### v1

# Little By Little: Continual Learning via Self-Activated Sparse Mixture-of-Rank Adaptive Learning

Abstract
Continual learning (CL) with large pre-trained models is challenged by catastrophic forgetting and task interference. Existing LoRA-based Mixture-of-Experts (MoE) approaches mitigate forgetting by assigning and freezing task-specific adapters, but suffer from interference, redundancy, and ambiguous routing due to coarse adapter-level selection. However, this design introduces three key challenges: 1) Interference: Activating full LoRA experts per input leads to subspace interference and prevents selective reuse of useful components across tasks. 2) Redundancy: Newly added experts often duplicate or contradict existing knowledge due to unnecessary activation of unrelated ranks and insufficient reuse of relevant ones. 3) Ambiguity: Overlapping features across tasks confuse the router, resulting in unstable expert assignments. As more experts accumulate, earlier task routing degrades, accelerating forgetting. We propose MoRA, a Mixture-of-Rank Adaptive learning approach with self-activated and sparse rank activation for CL. Unlike mixing multiple low-rank matrices, MoRA decomposes each rank-r update into r rank-1 components, each treated as an independent expert, enabling fine-grained mixture of rank-1 expert utilization while mitigating interference and redundancy. To avoid ambiguous routing, we propose that each rank-1 expert can infer its own relevance via intermediate activations. Coupled with our proposed rank pruning and activation budgets, MoRA adaptively selects a sparse mixture of ranks per input. We validate MoRA on continual learning tasks with CLIP and large language models (LLMs), analyzing both in-domain learning and out-of-domain forgetting/generalization during fine-tuning. MoRA shows significant effectiveness on enhancing CL with PTMs, and improving generalization while mitigating forgetting.
“Little by little, we gave you everything you ever dreamed of …” — “Little by little”, Oasis.
1 Introduction
Continual learning (CL) [29, 17, 73, 43, 21] aims to enable models to incrementally and efficiently acquire new knowledge from a stream of tasks, without catastrophic forgetting [55, 51] or the need for repeated fine-tuning on all previously seen data [77, 76, 85, 72]. The emergence of large pre-trained models (PTMs), including Vision Transformer (ViT) for vision [22], Large Language Model (LLM) [28, 61], and CLIP for vision–language embeddings [60, 33], has spurred a wave exploring CL with PTMs[94, 52, 72, 35, 81, 59, 62, 74, 58]. Despite the success of PTMs, CL on them across sequential tasks still faces key challenges, including preventing catastrophic forgetting and preserving—or even enhancing—the pre-trained capabilities to generalize to unseen or future cases [92, 85, 48, 58].
A common approach to integrating new knowledge into pre-trained models (PTMs) is continual full fine-tuning [92, 27] or parameter-efficient fine-tuning (PEFT) using a small set of learnable parameters, such as Low-Rank Adaptation (LoRA) [32] (Fig. 1(a)). However, continually updating the same set of parameters often causes interference and forgetting, even with a single new task [7]. Leveraging LoRA’s inherent regularization and flexibility, many CL methods for PTMs employ LoRA or its variants, along with specialized strategies to mitigate forgetting [72, 44, 74, 58, 44, 74].
Inspired by mixture-of-experts (MoE) architectures [65, 40, 24, 16] and the modular nature of LoRA adapters, MoE frameworks using LoRA as experts (Fig. 1(b)) have been widely adopted in CL [85, 72, 84, 13]. These methods incrementally train new LoRA experts while freezing previous ones, using a router to compose experts per sample. This isolation reduces cross-task interference and forgetting, while the composability of experts supports knowledge reuse and transfer [72, 58, 64]. In MoE-LoRA models, each expert comprises a full LoRA adapter with multiple ranks (i.e., subspaces), and expert selection occurs at this coarse-grained adapter level. Despite their success, this coarse granularity and the routing mechanism introduce three key challenges that limit effective utilization:
(1) Interference across experts and infeasible knowledge reuse. For any given sample, an entire LoRA expert must be either fully activated or not at all, leading to two major issues: interference among incompatible ranks from different experts, and limited flexibility in knowledge reuse.
(2) Learning redundancy and conflicts. Newly added experts may relearn redundant information or learn conflicting representation due to unnecessary activation of unrelated old ranks or insufficient reuse of relevant ones. This coarse selection exacerbates interference and forgetting during inference.
(3) Routing ambiguity and forgetting. Redundancy and conflicts among experts make it hard for a LoRA-level router to assign overlapping or common inputs correctly, leading to suboptimal mixtures [40, 24, 16]. As more new experts are accumulated in CL, the router’s assignments for earlier tasks become increasingly unreliable, makes it more prone to forgetting.
Figure 1: Conceptional illustration of CL with (a) LoRA, (b) MoE-LoRA, and (c) MoRA (Ours).
We propose MoRA, a Mixture-of-Rank Adaptive learning approaches with self-activated and sparsely selected ranks across the LoRAs added at different tasks. To address the limitations of coarse-grained MoE-LoRA methods, MoRA operates at a fine-grained level by treating each rank-one subspace within a LoRA adapter as an individual expert. Expert selection and routing are performed at the rank level, reducing interference from activating entire LoRA blocks. This decomposition promotes specialization on narrow regions of the data manifold, reduces redundancy, and enables effective reuse of shared knowledge across tasks.
We introduce a self-activation mechanism to select a sparse set of ranks without relying on a separately trained router, avoiding routing confusion and further mitigating forgetting (Fig. 1(c)). Each rank evaluates its own relevance to the input and determines its contribution, grounded in the view of low-rank updates as linear key–value memory. Sparsity is enforced via temperature-scaled softmax together with top-k masking and rank pruning under a rank activation budget. This design introduces no additional computation and eliminates the overhead associated with extra routing. Leveraging the connection between PTMs and low-dimensional manifolds [41, 1], MoRA enables precise, input-specific adaptation at test time through selective subspace updates. With fine-grained rank-level experts and sparse activation, MoRA facilitates continual learning by reusing relevant old ranks and selectively activating a few new ones per sample, enabling efficient, “little-by-little” adaptation that minimizes redundancy and supports essential updates. This selective, sample-specific activation at inference also enhances retention of pre-trained knowledge and improves generalization to unseen or future cases.
We apply and validate MoRA on CL with both vision-language CLIP and LLMs, as well as analyzing the effects on general pre-trained knowledge during low-rank fine-tuning. Across all settings, MoRA achieves strong downstream performance with significantly fewer active parameters, substantially reduces forgetting of both prior tasks and pre-trained knowledge, and even enhances generalization to unseen domains.
2 Related Work
Continual learning enables sequential knowledge acquisition without forgetting. Experience replay (ER) methods [49, 4, 10, 45, 11, 82, 83] interleave past examples with new data. Parameter regularization [36, 2, 86, 3, 34] penalizes updates to critical weights. Dynamic networks [70, 71, 93, 72, 52, 44, 77, 76, 66, 75, 72] allocate new capacity on the fly and preserve dedicated pathways for prior tasks.
Continual learning of PTMs. For CL on vision–language CLIP model [35, 89, 27], methods like ZSCL [92] retain zero-shot performance during adaptation, and follow-up work [85, 67, 81, 48] continually fine-tunes while leveraging frozen pre-trained predictions. The X-TAIL benchmark [81] further challenges models by mixing domain labels at test time. In large language models (LLMs) [18, 59, 62, 74, 58], continual learning uses capacity expansion or task-specific submodules to reduce interference. In this work, we also study how fine-tuning impacts pre-trained general knowledge.
Low-rank adaptation (LoRA) [32] is widely used for parameter-efficient fine-tuning of large pre-trained models. Building on this foundation, recent methods reformulate LoRA’s updates via SVD-based initialization and dynamic rank scheduling [87, 53, 20, 88, 46, 79], showing that task adaptation primarily lies in a compact subspace. In this work, we offer a complementary perspective by viewing both pre-trained weight matrix and its low-rank updates as a linear associative memory [41, 1, 37, 5]. Under this lens, a rank-$r$ update corresponds to writing $r$ new key–value entries into the pre-trained memory matrix, with each rank-1 component acting as an independent memory slot.
Mixture-of-experts (MoE) in LoRA fine-tuning. MoE increases model capacity by replacing the Transformer’s feed-forward layer with expert subnetworks, routing each input to a small subset and using load-balancing losses to even out usage [65, 40, 24, 16]. This paradigm has been adapted for standard fine-tuning [14, 42, 23] and CL [85, 72, 84, 13] with LoRA, treating each low-rank adapter as an expert: new adapters are trained and frozen per task to prevent forgetting. In contrast, we decompose each rank-$r$ update into $r$ rank-$1$ components and compute an input-dependent mixture over these fine-grained experts, greatly enhancing both expert specialization and mixture diversity.
3 Methods
3.1 Preliminaries
Continual learning. In continual learning, a model must sequentially learn from $T$ distinct tasks. For task $t\in\{1,\dots,T\}$, we denote its dataset by $\mathcal{D}^{t}=\{(\mathbf{x}^{t}_{i},y^{t}_{i})\}_{i=1}^{N^{t}}$, where $\mathbf{x}^{t}_{i}\in\mathbb{R}^{H\times W\times C}$ is the $i$th input sample, $y^{t}_{i}\in\mathcal{C}^{t}$ is its desired prediction, and $N^{t}$ is the number of examples in task $t$. At each time step $t$, the model may only access $\mathcal{D}^{t}$ and is prohibited from storing or replaying data from any $\mathcal{D}^{u}$ with $u<t$. This constraint makes it challenging to avoid catastrophic forgetting, which updates on new tasks overwrite knowledge acquired on earlier ones.
Low-rank adaptation. LoRA [32] parameterizes a low-rank update to a pre-trained weight matrix $\mathbf{W}_{0}\in\mathbb{R}^{d_{out}\times d_{in}}$ by introducing two factors $\mathbf{B}\in\mathbb{R}^{d_{out}\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times d_{in}}$, such that $\Delta\mathbf{W}=\mathbf{B}\mathbf{A}$, where $r\ll\min(d_{in},d_{out})$. The updated weight matrix is then defined as:
$\mathbf{W}=\mathbf{W}_{0}+\Delta\mathbf{W}=\mathbf{W}_{0}+\mathbf{B}\mathbf{A}.$ (1)
In this formulation, the original weights $\mathbf{W}_{0}$ remain fixed, and only $\mathbf{B}$ and $\mathbf{A}$ are trained, reducing the number of trainable parameters from $d_{in}d_{out}$ to $r(d_{in}+d_{out})$.
Mixture-of-Experts LoRA. Building on the Mixture-of-Experts (MoE) paradigm, a common framework of MoE-LoRA [85, 72] treats each LoRA as an independent expert and incrementally adds a new LoRA expert for each incoming task. After $T$ tasks, we have $T$ LoRA experts $\{(\mathbf{A}^{1},\mathbf{B}^{1}),\dots,(\mathbf{A}^{T},\mathbf{B}^{T})\}$. For an input token $x\in\mathbb{R}^{d_{\rm in}}$, the overall LoRA update in this framework is given by $\Delta\mathbf{W}=\sum\nolimits_{i=1}^{T}R(x)_{i}\mathbf{B}^{i}\,\mathbf{A}^{i},$ where the mixture weight $R(x)\in\mathbb{R}^{T}$ is produced by a learned router $R(\cdot)=\mathrm{softmax}(x\mathbf{W}_{r})$ and $\mathbf{W}_{r}\in\mathbb{R}^{d_{\rm in}\times T}$ contains the router’s trainable parameters. Each LoRA’s contribution is weighted by the learnable router, enabling the model to dynamically select and combine the most relevant low-rank updates for each token. In practice, a top-k masking is typically applied to mixture weights to enforce sparsity, activating only the $k$ most relevant experts and controlling computational overhead.
Figure 2: Overview of MoRA. For each new task, we freeze the ranks learned on previous tasks and introduce $r$ new ranks of updates. Our sparse self-activated mixture-of-ranks framework jointly considers all old and new ranks, adaptively inferring a sparse mixture weight for each rank. Panels (a,c) illustrate MoRA conceptually and (b,d) detail its computation for tasks $t$ and $t+1$, respectively.
3.2 Pre-trained Weight Matrices and Low-Rank Updates as Linear Associative Memory
Large PTMs have been shown to lie on a low “intrinsic dimension” manifold [41, 1]. From this perspective, the pre-trained weight matrix $\mathbf{W}_{0}\in\mathbb{R}^{d_{\mathrm{out}}\times d_{\mathrm{in}}}$ can itself be viewed as a linear associative memory [37, 5], having stored key–value pair of memories drawn from vast pre-training data. Specifically, any rank-$r$ update $\Delta\mathbf{W}\;=\;\mathbf{B}\,\mathbf{A}$ can be viewed as inserting or updating $r$ key–value entries into the pre-trained weight matrix. Concretely, let
$\mathbf{A}=\begin{bmatrix}\mathbf{A}_{1}^{\top}&\cdots&\mathbf{A}_{r}^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{r\times d_{\mathrm{in}}},\quad\mathbf{B}=\begin{bmatrix}\mathbf{B}_{1}&\cdots&\mathbf{B}_{r}\end{bmatrix}\in\mathbb{R}^{d_{\mathrm{out}}\times r},$ (2)
where each $\mathbf{A}_{i}\in\mathbb{R}^{1\times d_{\mathrm{in}}}$ serves as a key and each $\mathbf{B}_{i}\in\mathbb{R}^{d_{\mathrm{out}}\times 1}$ as its corresponding value. Then
$\Delta\mathbf{W}=\sum\nolimits_{i=1}^{r}\mathbf{B}_{i}\,\mathbf{A}_{i}$ (3)
writes the value $\mathbf{B}_{i}$ into the memory slot addressed by key $\mathbf{A}_{i}$. In this associative-memory view, each rank-$i$ component explicitly encodes “what to store” (the value $\mathbf{B}_{i}$) and “how to index it” (the key $\mathbf{A}_{i}$), thereby enriching the base model with newly acquired knowledge. More details are left in Appendix A.
3.3 Mixture of Ranks Adaptation
To overcome the coarse granularity of MoE-LoRA, we introduce mixture-of-ranks, applying MoE at the rank-level subspace within all LoRA adapters. By treating each rank as an independent expert, our fine-grained mixture framework enables: 1. Reduced redundancy and shared knowledge reuse: encoding common, cross-task features (e.g., generic semantics) can be reused, preventing redundant relearning. 2. Minimized interference and enhanced specialization: focusing exclusively on task-specific patterns, ensuring that new knowledge is learned without disrupting existing representations or interfered with future updates.
For any given input token $x$ from task $t$, we weight each rank’s contribution according to its relevance. We introduce a scalar mixture weight $w_{i}\in\mathbb{R}$ for each rank, yielding
$\Delta\mathbf{W}^{t}=\sum\nolimits_{i=1}^{r_{t}}w_{i}\mathbf{B}_{i}\mathbf{A}_{i},$ (4)
where each rank-1 update is analogous to an independent expert, $r_{t}$ is the total number of ranks learned up to task $t$, and $w_{i}$ modulates whether and how strongly the corresponding rank contributes to the final updated weight matrix. Unlike prior work that treats the entire rank-$r$ update as one expert [85, 72], this formulation enables fine-grained utilization of each rank’s update.
3.4 Self-Activated Adaptive Sparse Mixture of Ranks
Another key challenge in MoE and continual learning is the routing ambiguity and forgetting. To this end, we explore whether each rank-1 expert can autonomously infer its relevance to the current input, eliminating the need for a separate router and enabling self-activated expert selection.
Self-activated sparse mixture of ranks. Rather than relying on an external router, our framework derives mixture weights for each rank directly from the activations of each rank. Recall from Eq. (3) that each rank-$i$ update is parameterized by a key–value pair $(\mathbf{A}_{i},\mathbf{B}_{i})$. Given an input token $x\in\mathbb{R}^{d_{\mathrm{in}}}$ from task $t$, by treating $\mathbf{A}$ as the set of keys, we compute a raw score for all $r_{t}$ ranks accumulated up to task $t$:
$s_{i}\;=\;\frac{\mathbf{A}_{i}x}{\sqrt{\sum_{j=1}^{r_{t}}\|\mathbf{A}_{j}x\|_{2}^{2}}}\quad\in\mathbb{R},$ (5)
where the $\ell_{2}$–normalization over all $r_{t}$ ranks ensures numerical stability and compensates for variations in activation scales between layers and pre-trained weights. As shown in Table 5, using these raw scores alone matches or even exceeds the performance of a learned router for each rank, showing the feasibility that each rank can infer its own relevance using its activations.
To encourage rank specialization and focus the update on the most relevant ranks, we temperature-scale and normalize the scores:
$w_{i}\;=\;\mathrm{softmax}\!\bigl{(}\tfrac{s}{\tau_{\text{MoRA}}}\bigr{)}_{i},$ (6)
where the softmax is taken over the vector $s=[s_{1},\dots,s_{r_{t}}]$ and $\tau_{\text{MoRA}}$ is a temperature hyperparameter that controls the sharpness of the distribution. To ensure only a sparse, specialized subset of ranks contributes, we introduce an activation budget that caps the number of ranks that can be activated.
Adaptive sparse mixture of ranks. To ensure a stable and sparse activation pattern, we impose a fixed sparse rank-activation budget via top-k masking on the $\ell_{2}$–normalized scores $s_{i}:=\mathrm{TopK}(s,\,k),$ where
$[\mathrm{TopK}(s,k)]_{i}=\begin{cases}s_{i},&\text{if }s_{i}\text{ is among the top }k\text{ entries of }s,\\ -\infty,&\text{otherwise,}\end{cases}$ (7)
ensuring a sparse subset of at most $k$ out of the total $r_{t}$ ranks to be activated for task $t$ during CL.
We further refine the mixture weights by applying a fixed threshold $\delta$ to the $\ell_{2}$–normalized scores $s_{i}$. Concretely, after computing the softmax weights $w_{i}$, we define a binary mask $m_{i}=\mathbbm{1}\{s_{i}\geq\delta\}$, and update each gate via $w_{i}:=m_{i}\,\odot\,w_{i}$, where $\mathbbm{1}\{\cdot\}$ is an indicator function and $\odot$ denotes element-wise multiplication. This mask zeroes out any rank whose normalized activation falls below the threshold, ensuring that only the most relevant ranks contribute, resulting in a sparse mixture of an adaptive number of ranks based on the input.
Incremental mixture of ranks. Unlike methods that add a fixed number of LoRA experts per task [85], MoRA incrementally adds $r$ new ranks for each task, freezes all prior ranks, and adaptively selects a sparse subset of the most relevant ranks during both training and inference. This design seamlessly incorporates new task-specific knowledge into freshly initialized ranks while preserving earlier ranks unchanged to prevent forgetting. Moreover, it facilitates common knowledge reuse by allowing ranks that encode common patterns to be activated in later tasks.
Training Objectives. In our experiments, we only use the model’s standard training objective, without any extra regularization or load-balancing constraints. In practice, we empirically observe that the conventional load-balancing loss widely employed in the Mixture-of-Experts literature [65, 40, 24, 16] is not necessary, thanks to our fine-grained self-activated gating design. Imposing a strict load-balancing constraint might hinder the learning of the key vectors (i.e., matrix $\mathbf{A}$), since each rank should remain free to capture its own characteristic input distribution. Fig. 3 illustrates how each rank learns distinct characteristics.
Table 1: Comparisons on X-TAIL for each domain in terms of “Transfer”, “Average”, and “Last” scores (%). The best and the second best results are highlighted in red and blue, respectively.
    Method Aircraft [50] Caltech101 [25] DTD [15] EuroSAT [30] Flowers [56] Food [8] MNIST [19] OxfordPet [57] Cars [38] SUN397 [80] Average
CLIP
   Zero-shot 23.5 76.8 37.3 36.7 63.6 84.0 46.7 86.7 66.1 63.7 58.5
Transfer
   Zero-shot [60] 76.8 37.3 36.7 63.6 84.0 46.7 86.7 66.1 63.7 62.4
   LwF [43] 66.6 26.9 19.5 51.0 78.4 26.6 68.9 35.5 56.1 47.7
   WiSE-FT [78] 70.1 31.9 25.3 56.3 79.8 29.9 74.9 45.6 56.8 52.3
   iCaRL [63] 71.7 35.0 43.0 63.4 86.9 43.9 87.8 63.7 60.0 61.7
   ZSCL [92] 73.3 32.6 36.8 62.1 83.8 42.1 83.6 56.5 60.2 59.0
   MoE-Adapter [85] 71.0 34.9 19.2 63.0 86.6 20.0 87.2 63.7 58.6 56.0
   RAIL-Primal [81] 76.8 37.3 36.7 63.6 84.0 46.7 86.7 66.1 63.7 62.4
   CoDyRA [48] 74.3 36.8 44.2 69.9 83.5 42.8 88.9 64.6 63.4 63.2
   MoRA 74.5 38.1 46.9 65.3 82.9 45.8 88.2 65.1 62.9 63.3
Average
   LwF [43] 24.7 79.7 38.3 36.9 63.9 81.0 36.5 71.9 42.7 56.7 53.2
   WiSE-FT [78] 27.1 76.5 40.9 31.3 68.7 81.6 31.4 74.7 51.7 58.4 54.2
   iCaRL [63] 25.4 72.1 37.5 51.6 65.1 87.1 59.1 88.0 63.7 60.1 61.0
   ZSCL [92] 36.0 75.0 40.7 40.5 71.0 85.3 46.3 83.3 60.7 61.5 60.0
   MoE-Adapter [85] 43.6 77.9 52.1 34.7 75.9 86.3 45.2 87.4 66.6 60.2 63.0
   RAIL-Primal [81] 42.4 89.8 55.7 68.5 84.0 83.3 65.3 85.8 67.9 64.5 70.7
   CoDyRA [48] 41.4 81.0 58.7 77.8 83.4 84.6 64.5 90.4 67.2 64.4 71.3
   MoRA 44.1 81.6 64.6 79.6 83.9 84.4 66.5 89.7 68.4 64.1 72.7
Last
   LwF [43] 25.5 72.1 38.9 55.4 65.5 87.3 81.9 88.6 63.6 61.5 64.0
   WiSE-FT [78] 21.8 76.8 42.9 20.8 77.5 84.9 30.7 76.6 75.8 72.5 58.0
   iCaRL [63] 25.5 72.1 38.9 55.4 65.5 87.3 81.9 88.6 63.6 61.5 64.0
   ZSCL [92] 33.1 75.3 43.5 35.2 74.6 87.4 50.4 84.2 77.3 73.4 63.4
   MoE-Adapter [85] 43.2 78.7 57.6 32.8 79.4 86.0 86.7 87.8 78.2 74.2 70.5
   RAIL-Primal [81] 41.7 94.0 66.0 86.4 97.2 82.4 93.1 83.6 75.0 71.3 79.1
   CoDyRA [48] 37.7 81.5 65.1 89.9 91.4 85.5 96.8 93.3 77.3 73.5 79.2
   MoRA 37.7 81.5 70.7 92.4 95.0 86.0 97.6 92.6 81.0 74.7 80.9
4 Experiments
In this paper, we conduct experiments across a diverse set of tasks, including continual learning for both vision-language CLIP models and LLMs, and analyze catastrophic forgetting during fine-tuning. Detailed implementation settings and more experiment results are provided in the Appendix.
Table 2: Comparisons on 5-shot MTIL setting. Following the same protocol as in [85, 81, 48].
    Method Aircraft [50] Caltech101 [25] CIFAR100 [39] DTD [15] EuroSAT [30] Flowers [56] Food [8] MNIST [19] OxfordPet [57] Cars [38] SUN397 [80] Average
CLIP
   Zero-shot [60] 24.3 88.4 68.2 44.6 54.9 71.0 88.5 59.4 89.0 64.7 65.2 65.3
Transfer
   Zero-shot [60] 88.4 68.2 44.6 54.9 71.0 88.5 59.6 89.0 64.7 65.2 69.4
   LwF [43] 72.1 49.2 35.9 44.5 41.1 66.6 50.5 69.0 19.0 51.7 50.0
   LwF-VR [21] 82.2 62.5 40.1 40.1 56.3 80.0 60.9 77.6 40.5 60.8 60.1
   WiSE-FT [78] 77.6 60.0 41.3 39.4 53.0 76.6 58.1 75.5 37.3 58.2 57.7
   ZSCL [92] 84.0 68.1 44.8 46.8 63.6 84.9 61.4 81.4 55.5 62.2 65.3
   MoE-Adapter [85] 87.9 68.2 44.1 48.1 64.7 88.8 69.0 89.1 64.5 65.1 68.9
   RAIL-Primal [81] 88.4 68.2 44.6 54.9 71.0 88.5 59.6 89.0 64.7 65.2 69.4
   CoDyRA [48] 92.4 68.4 45.8 54.5 69.6 87.4 65.2 88.5 64.2 64.5 69.9
   MoRA 92.0 68.8 45.6 53.1 68.6 84.4 64.3 89.8 65.4 64.8 69.7
Average
   LwF [43] 23.5 77.4 43.5 41.7 43.5 52.2 54.6 63.4 68.0 21.3 52.6 49.2
   LwF-VR [21] 24.9 89.1 64.2 53.4 54.3 70.8 79.2 66.5 79.2 44.1 61.6 62.5
   WiSE-FT [78] 32.0 87.7 61.0 55.8 68.1 69.3 76.8 71.5 77.6 42.0 59.3 63.7
   ZSCL [92] 28.2 88.6 66.5 53.5 56.3 73.4 83.1 56.4 82.4 57.5 62.9 64.4
   MoE-Adapter [85] 30.0 89.6 73.9 58.7 69.3 79.3 88.1 76.5 89.1 65.3 65.8 71.4
   RAIL-Primal [81] 32.9 94.5 69.9 58.1 71.8 84.4 88.5 70.4 89.0 66.1 65.7 71.9
   CoDyRA [48] 34.6 95.8 73.9 60.0 77.1 81.3 86.6 75.9 89.9 66.1 65.3 73.3
   MoRA 36.7 95.4 74.9 61.9 77.1 82.6 85.3 76.0 90.5 67.0 65.6 73.9
Last
   LwF [43] 22.1 58.2 17.9 32.1 28.1 66.7 46.0 84.3 64.1 31.5 60.1 46.5
   LwF-VR [21] 22.9 89.8 59.3 57.1 57.6 79.2 78.3 77.7 83.6 60.1 69.8 66.9
   WiSE-FT [78] 30.8 88.9 59.6 60.3 80.9 81.7 77.1 94.9 83.2 62.8 70.0 71.9
   ZSCL [92] 26.8 88.5 63.7 55.7 60.2 82.1 82.6 58.6 85.9 66.7 70.4 67.4
   MoE-Adapter [85] 30.1 89.3 74.9 64.0 82.3 89.4 87.1 89.0 89.1 69.5 72.5 76.1
   RAIL-Primal [81] 32.9 95.1 70.3 63.2 81.5 95.6 88.5 89.7 89.0 72.5 71.0 77.2
   CoDyRA [48] 31.6 95.5 72.8 63.5 85.0 89.7 85.0 94.7 93.2 73.6 73.0 78.0
   MoRA 32.5 95.3 75.3 66.6 87.8 92.6 86.3 96.3 92.6 73.5 73.8 79.3
CL of CLIP We evaluate on two benchmarks using CLIP-ViT/B-16 [60, 33]: MTIL [92, 85], with 11 datasets learned sequentially under a 5-shot split, and X-TAIL [81, 48], with a 16-shot split. Following prior work, we report Transfer (zero-shot on unseen tasks), Last (retention on earlier tasks), and Average (mean of Transfer and Last) to assess continual learning performance.
CL of LLM. We follow prior work [59, 62, 74, 58] to continually fine-tune T5-large [61] on five text-classification benchmarks—AG News, Amazon Reviews, Yelp Reviews, DBpedia, and Yahoo Answers—under three different task orderings. We report the average final accuracy across all tasks after completing the last task, measuring the model’s ability to retain knowledge during CL.
Generalization and forgetting on unseen tasks. To assess effects on pre-trained general knowledge, we fine-tune Llama3 [28] on the CodeAlpaca code-generation dataset [9] and evaluate zero-shot in-domain performance on HumanEval [12], as well as out-of-domain accuracy on a broad selection of MMLU [31] subjects—Formal Logic, Philosophy, World Religions, Economics, Public Relations, STEM, Physics, and Machine Learning. These results reveal how fine-tuning impacts both task-specific adaptation and catastrophic forgetting or improved generalization to unseen domains.
4.1 Continual Learning of CLIP
Cross-domain task-agnostic incremental learning. Table 1 presents our results on the X-TAIL benchmark. Prior methods [85, 81] rely on zero-shot predictions to preserve pre-trained capabilities and mitigate forgetting, and as a result are capped by the base model’s zero-shot performance. Similarly as done in previous works [92, 48], MoRA aims to continuously refines the pre-trained model, enabling it to surpass this zero-shot ceiling and improve the generalization of the pre-trained model on unseen domains. In practice, continual parameter updates can slightly perturb some zero-shot predictions, our method ultimately improves overall performance on unseen domains by adapting the pre-trained model to new tasks while enhancing its generalization capabilities.
Table 3: Summary of the results on standard CL benchmarks with T5-large model. The averaged accuracy after training on the last task is reported. Experiments on long task sequences in Table 9.
Standard CL Benchmark
Method Order-1 Order-2 Order-3 Avg.
MTL 80.0
SeqFT 18.9 24.9 41.7 28.5
SeqLoRA 44.6 32.7 53.7 43.7
IncLoRA 66.0 64.9 68.3 66.4
Replay 55.2 56.9 61.3 57.8
EWC 48.7 47.7 54.5 50.3
LwF 54.4 53.1 49.6 52.3
L2P 60.3 61.7 61.1 60.7
LFPT5 67.6 72.6 77.9 72.7
O-LoRA 75.4 75.7 76.3 75.8
LB-CL 76.9 76.5 76.8 76.7
MoRA 77.43 77.47 77.93 77.61
On the Last and Average metrics, MoRA significantly outperforms prior methods. Unlike approaches that require domain ID prediction [85], high dimensional projection layers, or a memory bank of feature embeddings [81], MoRA enhances the pre-trained model’s own representation capability without relying on these external mechanisms. In contrast to the previous state of the art [48], which learns a fixed weight for each rank-$r$ update from each task, MoRA’s novel mixture-of-ranks framework selectively assigns each rank an input-dependent gate. Notably, only a small subset of ranks most relevant to each token are activated, resulting in far more fine-grained utilization and specialization of each ranks.
Multi-domain task incremental learning. We evaluate MoRA in the few-shot MTIL setting (Table 2) under the same protocols as [85, 81, 48]. Consistent with the results observed in the X-TAIL setting, our method demonstrates clear superiority in this scenario.
4.2 Continual Learning of LLM
Table 3 reports results across three task orderings, where MoRA consistently outperforms prior methods and nearly matches the multi-task learning (MTL) upper bound. In contrast to state-of-the-art techniques such as O-LoRA [74] and LB-CL [58], which rely on orthogonality constraints or gradient projections between per-task LoRA adapters—potentially limiting each adapter’s capacity—MoRA requires no additional regularization. By decomposing each rank-$r$ update into its rank-1 components and applying self-activated, sparse gating, MoRA allows every component to specialize on its own input distribution. This fine-grained specialization both prevents interference across tasks and more effectively captures diverse patterns.
4.3 Forgetting and Generalization during Fine-Tuning
Table 4 examines how standard fine-tuning affects both in-domain performance and out-of-domain generalization. We observe two key phenomena: 1. Forgetting of unrelated knowledge. Fine-tuning with a conventional LoRA leads to performance drops on topics completely unrelated to the fine-tuning data—for example, Philosophy, World Religions, and Public Relations—indicating that the model loses pre-trained general knowledge when adapting to a new domain. 2. Improved generalization. Fine-tuning on code generation data also boosts accuracy on tasks requiring logical understanding (Formal Logic, Economics, STEM, Physics, and Machine Learning), suggesting that representations learned during code fine-tuning transfer to related symbolic and quantitative domains.
By contrast, MoRA’s self-routed, sparse mixture-of-ranks framework both preserves unrelated general knowledge and enhances in-domain and out-of-domain generalization performance. By decomposing each low-rank update into specialized rank-1 components and activating only the most relevant subset, MoRA prevents catastrophic overwriting of prior knowledge. At the same time, it achieves strong code-generation performance using only one-third of the activated parameters required by a rank-32 LoRA, and significantly improves the out-of-domain generalization.
Table 4: Fine-tuning Llama3.1-8B on the code generation domain (CodeAlpaca). We report zero-shot in-domain performance on HumanEval (Pass@1) and zero-shot out-of-domain accuracy for selected MMLU subjects—Formal Logic, Philosophy, World Religions, Economics, Public Relations, STEM, Physics, and Machine Learning. The last two columns list trainable parameters (for MoRA: added / activated). The best and second best scores are highlighted in red and blue, respectively.
[…]
Logic Phil. Reli. Econ. Pub. Rel. STEM Phys. ML MMLU
Llama-3.1-8B 38.40 42.06 71.06 83.63 70.17 68.18 54.84 39.22 40.18 63.45
LoRA ($r=4$) 41.46 39.68 70.09 81.87 71.43 65.45 54.77 45.10 40.17 63.28 10.5 0.13%
LoRA ($r=8$) 44.51 39.68 70.74 81.87 71.85 64.54 54.17 42.16 39.29 63.03 21.0 0.26%
LoRA ($r=16$) 45.73 41.27 68.49 80.70 72.69 66.36 54.96 44.11 38.39 63.35 41.9 0.52%
LoRA ($r=32$) 47.56 42.85 69.45 81.87 72.27 66.36 55.44 45.10 39.29 63.59 83.9 1.03%
MoRA 47.56 48.41 70.09 82.46 73.53 68.18 55.53 46.08 41.96 63.70 41.9/26.2 0.52%/0.32%
[…]
(a) Rank Activations of MoRA on data from Task 1 after learning Task 1.
[…]
(b) Rank Activations of MoRA on data from Task 1 after learning Task 2.
[…]
(c) Rank Activations of MoRA on data from Task 2 after learning Task 2.
Figure 3: Visualization of MoRA rank activations during Task 1 and Task 2 training. Activations are extracted from the K projection in the attention module (layer 8) of the image encoder. Corresponding image patches are shown below each activation map, with regions relevant to each class marked by orange bounding boxes. Zoom in for details. More visualizations are in Figs. 5 and 6 of Appendix.
4.4 Visualizations on Fine-grained Rank Activations
We visualize rank activations during continual learning in Fig. 3. These visualizations demonstrate that MoRA (1) encourages each rank to specialize in its own input distribution, and (2) reduces cross-task interference to mitigate forgetting.
Each rank specializes in distinct input patterns. In Fig. 6(a), we observed that patches corresponding to the airplane object strongly activate rank 0, whereas background patches (blue sky) predominantly activate rank 11. Similarly as in Fig. 6(c), ranks at index 19, 20, and 29 (highlighted with orange bounding boxes) are activated for image patches of jaguar. Moreover, the complex background elements in Fig. 6(c)—such as tree leaves and shadows—elicit a more diverse activation pattern than the simple blue-sky backgrounds in Fig. 6(a), illustrating MoRA’s capacity to capture varying contextual complexity. Interestingly, we also observe that certain ranks learned on earlier tasks remain active when processing later-task inputs, indicating effective knowledge reuse.
Reduced cross-task interference to mitigate forgetting. By comparing the activation patterns in Fig. 6(a) (after learning Task 1) and Fig. 6(b) (after learning Task 2) on the same input image, we observe that the ranks activations are nearly identical: rank 0 continues to respond to airplane semantics, and rank 11 continues to respond to background blue-sky patches. This consistency demonstrates that our self-activated sparse mixture-of-ranks framework effectively prevents later-task updates from overwriting or intruding upon the knowledge learned for earlier tasks. As a result, inter-task interference is greatly reduced, and catastrophic forgetting is mitigated.
Table 5: Comparisons on different per LoRA/rank mixture calculation strategy
Mixture Calculation Transfer Average Last
Router@LoRA 62.56 69.45 74.53
Router@rank 60.09 65.97 69.76
Self-activated@rank 60.26 65.94 69.85
     + Sparse activation 62.59 72.28 80.22
     + Adaptive sparse (MoRA) 63.30 72.70 80.90
(a) Acti. Budget
(b) Temp. $\tau_{\text{MoRA}}$
(c) Thre. $\delta$
Figure 4: Ablation on (a) rank activation budget, (b) temperature $\tau_{\text{MoRA}}$, and (c) threshold $\delta$.
4.5 Ablation Studies
Per-LoRA/rank mixture weights. Table 5 compares different strategies for computing mixture weights, under a controlled setup with one LoRA per task, standard training loss only, and no additional regularization: (1) Router@LoRA trains a standard router to assign weights to each LoRA adapter. (2) Router@rank trains a fine-grained router for every rank-1 component, which increases training complexity and underperforms Router@LoRA. (3) Self-activated@rank directly infer mixture weights using softmax-normalized activations (Eq. 5), matching or exceeding Router@rank and showing that each rank can infer its own relevance without extra parameters. (4) Sparse activation adds temperature scaling (Eq. 6) and caps the number of activated ranks, yielding substantial gains by sharpening rank specialization. (5) Adaptive sparse activation further prunes low-activation ranks at test time, selecting the most relevant number of ranks to be activated adaptively for each input token, reducing noise from less relevant ranks and achieving the best overall performance.
Rank activation budget. Under our sparse rank-activation framework, we explore how many ranks are necessary by varying the activation budget (Fig. 4(a)). To focus exclusively on this budget parameter, we disable other sparsity enhancements, e.g. rank pruning. After training on 10 tasks (160 ranks total), raising the budget from 2 (1.25%) to 16 active ranks (10%) yields a dramatic boost in final accuracy, and gradually plateaus. This demonstrates that each rank specializes in a distinct input distribution, and a small, sparse subset of ranks suffices to handle diverse input distributions. Moreover, even when the budget is set very high, our self-activating gating remains robust: it still activates only the most relevant ranks for each token while suppressing contributions from less relevant ones.
Temperature setting of $\tau_{\text{MoRA}}$. We use temperature scaling (Eq. (6)) to control mixture sharpness: lower $\tau_{\text{MoRA}}$ yields sharper activations, while higher $\tau_{\text{MoRA}}$ flattens the distribution and engages more ranks. Figure4(b) shows that increasing $\tau_{\text{MoRA}}$ improves transfer to unseen tasks, even surpassing the reported state-of-the-art Transfer in Table 1 when $\tau_{\text{MoRA}}$ is high. A moderate $\tau_{\text{MoRA}}$ around 0.01 balances specificity and generalization, delivering strong Last and Average scores.
Adaptive mixture with rank pruning. In Fig. 4(c), we demonstrate the effects of adaptive sparsity constraints by varying the threshold $\delta$ for rank pruning. Pruning ranks with a relatively small threshold reduces noise and sharpens the mixture, improving both downstream adaptation and out-of-domain generalization, whereas raising $\delta$ above 0.2 begins to remove useful ranks and degrades performance.
5 Conclusion
We present MoRA, a self-activated sparse mixture-of-ranks framework that decomposes each LoRA update into rank-1 experts and self-infer their sparse mixture weights. By reducing knowledge redundancy and resolving router–expert mismatch during CL, MoRA delivers superior continual-learning performance, significantly reduces forgetting, and enhances generalization.
Limitations and future works. MoRA use a pre-defined temperature to control the sharpness of the mixture weights. While effective, the sparse activation mechanism can be further improved by incorporating awareness of the global data distribution. Learning this temperature or adapting the mixture’s sharpness based on data offers a promising avenue for future research. We also aim to extend the method’s applicability to a broader range of PTMs and applications.
Acknowledgments and Disclosure of Funding
This work was partially supported by the ARC DECRA Fellowship (DE230101591) awarded to D. Gong. H. Lu is affiliated with CSIRO Data61 through a PhD scholarship and acknowledges the support of the Google PhD Fellowship.
Appendix A Additional Details of MoRA
We provide more details of developing the proposed MoRA in this section.
A.1 Model Weights and Updates as Linear Associative Memory Model
Weight matrix as linear associative memory. Large pre-trained models (PTMs) have been shown to inhabit a low “intrinsic-dimension” manifold compared to their full parameter count [41, 1]. Any weight matrix of a pre-trained model $\mathbf{W}_{0}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$ can be treated as a linear associative memory model with a key-value structure [37, 5, 6, 54]. Concretely, for any matrix $\mathbf{W}\in\mathbb{R}^{d_{\mathrm{out}}\times d_{\mathrm{in}}}$ and input $\mathbf{x}\in\mathbb{R}^{d_{\mathrm{in}}}$, the product $\mathbf{W}\,\mathbf{x}$ implements a content-addressable memory read over $m$ stored key–value pairs $\{(\mathbf{K}_{i},\mathbf{V}_{i})\}_{i=1}^{m}$, where $m$ denotes the number of key-value memories of a specific weight matrix. Here, ${\mathbf{K}}=\begin{bmatrix}{\mathbf{K}}_{1}^{\top}&\cdots&{\mathbf{K}}_{m}^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{m\times d_{\text{in}}}$ and ${\mathbf{V}}=$$\begin{bmatrix}{\mathbf{V}}_{1}&\cdots&{\mathbf{V}}_{m}\end{bmatrix}\in\mathbb{R}^{d_{\text{out}}\times m}$ with $\mathbf{K}_{i}\in\mathbb{R}^{1\times d_{\text{in}}}$ is the $i$-th row of ${\mathbf{K}}$ and ${\mathbf{V}}_{i}\in\mathbb{R}^{d_{\text{out}}\times 1}$ is the $i$-th column of ${\mathbf{V}}$. Given any input $\mathbf{x}$, the output $\mathbf{y}\in\mathbb{R}^{d_{\text{out}}}$ is computed as
$\mathbf{y}=\mathbf{W}\mathbf{x}\approx\sum\nolimits_{i=1}^{m}\mathbf{V}_{i}(\mathbf{K}_{i}\mathbf{x}).$ (8)
In this formulation, the forward pass through the weight matrix with the input decomposes into two steps: (1) computing response scores or assignment weights via the key vectors $\mathbf{K}_{i}$, and (2) aggregating the corresponding value vectors $\mathbf{V}_{i}$ according to those weights.
Low-rank update as modifications to the linear associative memory. Building on the associative-memory view of weights, we can likewise interpret weight updates during training as key–value insertions into the existing memory. During training, for an arbitrary update $\Delta\mathbf{W}$ on a given $\mathbf{W}_{0}$ as the starting point, the weight update process can be represent as
$\mathbf{W}_{\text{new}}=\mathbf{W}_{0}+\Delta\mathbf{W},$ (9)
where $\mathbf{W}_{0}$ can be the pre-trained model weights. We can also formulate the weight updates $\Delta\mathbf{W}$ with the key-value memory model $\{\Delta\mathbf{K},\Delta\mathbf{V}\}$ as an update to the original memory model. With the operation on an input $\mathbf{x}$ via
$\Delta\mathbf{W}\mathbf{x}=\Delta\mathbf{V}\Delta\mathbf{K}\mathbf{x},$ (10)
the updates on weights can be seen as a retrievable key-value system for the updating of $\mathbf{W}\mathbf{x}$.
Low-Rank Adaptation (LoRA) [32] performs parameter-efficient fine-tuning by adding a rank-$r$ update $\Delta\mathbf{W}=\mathbf{B}\mathbf{A}$ to the pre-trained weight matrix $\mathbf{W}_{0}$, where $\mathbf{B}\in\mathbb{R}^{d_{\text{out}}\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times d_{in}}$. With the model update based on LoRA, given an input $\mathbf{x}$, the computation of updated weight becomes
$\mathbf{y}_{\text{new}}=\mathbf{W}_{\text{new}}\mathbf{x}=\mathbf{W}_{0}\mathbf{x}+\Delta\mathbf{W}\mathbf{x}=\mathbf{W}_{0}\mathbf{x}+\mathbf{B}\mathbf{A}\mathbf{x}.$ (11)
Decomposing the low-rank matrices into a collection of its $r$ rank-1 components gives
$\Delta\mathbf{W}\mathbf{x}=\mathbf{B}\mathbf{A}\mathbf{x}=\sum\nolimits_{i=1}^{r}\mathbf{B}_{i}(\mathbf{A}_{i}\mathbf{x}),$ (12)
with $\mathbf{A}=\begin{bmatrix}\mathbf{A}_{1}^{\top}&\cdots&\mathbf{A}_{r}^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{r\times d_{\text{in}}}$ where $\mathbf{A}_{i}$ is the $i$-th row of $\mathbf{A}$, and $\mathbf{B}=\begin{bmatrix}\mathbf{B}_{1}&\cdots&\mathbf{B}_{r}\end{bmatrix}\in\mathbb{R}^{d_{\text{out}}\times r}$ where $\mathbf{B}_{i}$ is the $i$-th column of $\mathbf{B}$.
Considering the connection between the formulation of LoRA Eq. (1) and associative-memory view of weight update Eq. (10), we can naturally interpret LoRA as maintaining a small key–value store in a low-rank subspace, where the rows of $\mathbf{A}$ serve as keys and the columns of $\mathbf{B}$ serve as their corresponding values. After training the LoRA module on a base weight matrix $\mathbf{W}_{0}$, the modified forward pass in Eq. (11) with Eq. (12) can be seen as: first, the original pre-trained computation $\mathbf{W}_{0}\,\mathbf{x}$ (Eq. (8)) is performed; then, each learned key $\mathbf{A}_{i}$ measures its relevance to the input $\mathbf{x}$ via the product $\mathbf{A}_{i}\mathbf{x}$, and each learned value $\mathbf{B}_{i}$ is scaled by this relevance score and aggregated, in a same way as shown in Eq. (8). In this way, LoRA effectively inserts or updates a small set of key–value memories that augment the pre-trained model’s capabilities.
A.2 Sparse Mixture-of-Rank Adaptation
In the key–value associative memory view of LoRA (Eq. (11) and Eq. (12)), a standard LoRA or its variants [53, 87] apply a dense, fixed update: every rank’s subspace contributes for any kinds of input $\mathbf{x}$. Even in MoE-LoRA, where each LoRA module is treated as a separate expert, all key-value pairs within a LoRA module are used densely as a whole, either fully activated or not activated at all. This leads to redundancy across experts and overlooked the strong potential of the key vectors $\mathbf{A}$ on expressing its relevance to the input.
Rather than activating all key-value memories of each LoRA for all kinds of inputs, where many of which may encode patterns irrelevant to the current input, we introduce a gated, sparse mixture-of-ranks update that selectively activates only the most relevant key–value pairs for each input. Concretely, we associate each rank-1 key–value pair with a scalar weight $w_{i}\in\mathbb{R}$, so that for an input token $\mathbf{x}\in\mathbb{R}^{d_{\mathrm{in}}}$ from task $t$, with $r_{t}$ total ranks, the update becomes:
$\Delta\mathbf{W}^{t}=\sum\nolimits_{i=1}^{r_{t}}w_{i}\mathbf{B}_{i}\mathbf{A}_{i},~{}~{}\text{and}~{}~{}\Delta\mathbf{W}^{t}\mathbf{x}=\sum\nolimits_{i=1}^{r_{t}}w_{i}\mathbf{B}_{i}\mathbf{A}_{i}\mathbf{x},$ (4)
where $w_{i}$ modulates whether and how strongly the corresponding rank contributes to the final updated weight matrix.
As discussed in Sec. 3.4, rather than obtaining mixture weights $w_{i}$’s through learning an additional router [85, 72], considering that $\mathbf{A}_{i}$’s in $\mathbf{A}$ works as keys in the memory model that reflect the significance on the correlation between input and the learned knowledge in each rank, we leverage the key vectors $\mathbf{A}_{i}$themselves to infer each rank’s relevance. Specifically, $w_{i}$’s are obtained through sparsity-inducing scaling with top-k masking and rank pruning of each rank’s activation $\mathbf{A}_{i}\mathbf{x}$, i.e., $[w_{i}]=\texttt{SparseOperator}([\mathbf{A}_{i}\mathbf{x}])$, where $\texttt{SparseOperator}()$ denotes the sparse-inducing operations.
Our fine-grained mixture framework offers two key advantages: (1) It reduces redundancy and enables shared-knowledge reuse: ranks that capture common, cross-task features (such as generic semantics) can be reactivated across multiple tasks, preventing those patterns from being relearned redundantly or interfered; (2) It minimizes routing ambiguity and prevents forgetting: each rank can specialize on more nuanced patterns, yielding an optimized, rank-wise mixture that integrates new knowledge without disrupting existing representations or hindering future updates.
Appendix B Additional Details and Results of Experiments
Table 6: Details of the 15 datasets used in our continual-learning experiments using LLM. NLI denotes natural language inference, and QA denotes question-answering tasks. The first five tasks comprise the standard CL benchmark; the remaining ten tasks are used for the extended long-sequence evaluations.
Dataset name Category Task Domain Metric
1. Yelp CL Benchmark sentiment analysis Yelp reviews accuracy
2. Amazon CL Benchmark sentiment analysis Amazon reviews accuracy
3. DBpedia CL Benchmark topic classification Wikipedia accuracy
4. Yahoo CL Benchmark topic classification Yahoo Q&A accuracy
5. AG News CL Benchmark topic classification news accuracy
6. MNLI GLUE NLI various accuracy
7. QQP GLUE paragraph detection Quora accuracy
8. RTE GLUE NLI news, Wikipedia accuracy
9. SST-2 GLUE sentiment analysis movie reviews accuracy
10. WiC SuperGLUE word sense disambiguation lexical databases accuracy
11. CB SuperGLUE NLI various accuracy
12. COPA SuperGLUE QA blogs, encyclopedia accuracy
13. BoolQA SuperGLUE boolean QA Wikipedia accuracy
14. MultiRC SuperGLUE QA various accuracy
15. IMDB SuperGLUE sentiment analysis movie reviews accuracy
Table 7: Instructions for different tasks.
Task Prompts
NLI
What is the logical relationship between the "sentence 1" and the "sentence 2"?
Choose one from the option.
QQP
Whether the "first sentence" and the "second sentence" have the same meaning?
Choose one from the option.
SC
What is the sentiment of the following paragraph? Choose one from the option.
TC
What is the topic of the following paragraph? Choose one from the option.
BoolQA
According to the following passage, is the question true or false? Choose one
from the option.
MultiRC
According to the following passage and question, is the candidate answer true
or false? Choose one from the option.
WiC
Given a word and two sentences, whether the word is used with the same sense
in both sentence? Choose one from the option.
Table 8: The six task-sequence orders used in our continual learning experiments. Sequences 1–3 follow the standard CL benchmarks employed in prior work. Sequences 4–6 extend to longer 15-task streams, as introduced in [62].
Order Task Sequence
1 dbpedia → amazon → yahoo → ag
2 dbpedia → amazon → ag → yahoo
3 yahoo → amazon → ag → dbpedia
4
mnli → cb → wic → copa → qqp → boolqa → rte → imdb →
yelp → amazon → sst-2 → dbpedia → ag → multirc → yahoo
5
multirc → boolqa → wic → mnli → cb → copa → qqp → rte
→ imdb → sst-2 → dbpedia → ag → yelp → amazon → yahoo
6
yelp → amazon → mnli → cb → copa → qqp → rte → imdb →
sst-2 → dbpedia → ag → yahoo → multirc → boolqa → wic
B.1 Detailed Experiment Settings
CL on CLIP. We follow the experimental setups in [92, 85, 81] and use the CLIP model with a ViT-B/16 backbone [60] for all experiments. By default, MoRA is applied to every pre-trained weight matrix in both the vision and text encoders, with an initial rank of 16 per update. Each task is trained for 500 iterations using AdamW [47] with learning rate of $45e-$ 4. During continual learning, we freeze the ranks learned from previous tasks and initialize new $r=16$ ranks for each incoming task. The rank activation budget is set to 16 throughout all tasks. We set the temperature $\tau_{\text{MoRA}}=0.1$ and the threshold $\delta=0.2$.
CL on LLMs. We follow the protocol of previous work in continually fine-tuning a T5-large model on a suite of text-classification tasks. We train on five standard benchmarks—AG News, Amazon Reviews, Yelp Reviews, DBpedia, and Yahoo Answers—using three distinct task orderings drawn from [59, 62, 74, 58]. To probe longer sequences, we extend this to a 15-dataset stream (Table 6), incorporating tasks from the original CL benchmark [90], GLUE [68], SuperGLUE [69], and the IMDB movie reviews corpus. Natural language prompts for each task are presented in Table 7, with NLI tasks (MNLI, RTE, CB), sentiment classification (Amazon, Yelp, SST-2, IMDB), and topic classification (AG News, DBpedia, Yahoo).
We evaluate three distinct task sequences for both the standard CL and 15-task benchmarks (Table 8). After completing the final task in each stream, we report the average accuracy across all tasks. All experiments use one epoch per task with DeepSpeed, a fixed learning rate of $31e-$ 3, batch size 64, and dropout of 0.1. MoRA is applied to both the query and key projection matrices within attention layers, initializing $r=8$ new ranks for each incoming task similarly as in [74, 58]. We maintain a constant activation budget of 4 ranks throughout continual learning, set the temperature $\tau_{\text{MoRA}}=0.1$, and the threshold $\delta=0.2$.
Generalization and forgetting on unseen tasks. To assess effects on pre-trained general knowledge, we fine-tune Llama3.1-8B [28] on the CodeAlpaca code-generation dataset [9] using llama-Factory [91] and evaluate using lm-eval-harness [26] on zero-shot in-domain performance on HumanEval [12], as well as out-of-domain accuracy on a broad selection of MMLU [31] subjects—Formal Logic, Philosophy, World Religions, Economics, Public Relations, STEM, Physics, and Machine Learning.
In this experiment, MoRA is applied to all linear weight matrices of the pre-trained model. We fine-tune on CodeAlpaca with a batch size of 32 over 3 epochs and a cosine learning-rate schedule starting at $45e-$ 4. We train with $r=16$ ranks and enforce a constant activation budget of 4 ranks. The self-routed gating uses a temperature $\tau_{\text{MoRA}}=0.5$ and a threshold $\delta=0.2$. We observe that, due to variations in hidden representations across architectures, the optimal temperature setting can differ across each pre-trained models.
Reproducibility Statement. Our source code will be publicly released upon paper acceptance. Results are averaged over three independent runs to ensure robustness. Experiments with the CLIP model were carried out on an NVIDIA GeForce RTX 3090 GPU, while experiments with LLMs utilized 4 NVIDIA H100 GPUs.
Table 9: Average accuracy on T5-large continual-learning benchmarks after the final task, evaluated over extended 15-task sequences. Results for prior methods are taken from [58].
Large Number of Tasks
Method Order-4 Order-5 Order-6 Avg.
MTL 76.5
SeqFT 7.4 7.4 7.5 7.4
SeqLoRA 2.3 0.6 1.9 1.6
IncLoRA 63.3 58.5 61.7 61.2
Replay 55 54.6 53.1 54.2
EWC 45.3 44.5 45.6 45.1
LwF 50.1 43.1 47.4 46.9
L2P 57.5 53.8 56.9 56.1
LFPT5 69.8 67.2 69.2 68.7
O-LoRA 70.5 65.5 70.5 68.8
LB-CL 68.4 67.3 71.8 69.2
MoRA 68.91 68.32 71.95 69.72
B.2 Continual Learning of LLM on Long Task Sequences
In Sec. 4.2 Table 3, we evaluate MoRA on standard CL benchmarks [62]. In Table 9, we extend the evaluation to challenged long task sequences using 15 datasets with 3 different orderings as in [74, 58]. Consistent with the findings in Table 3, MoRA outperforms previous methods in terms of averaged performance across three task orders, and largely close the gap to multi-task learning. Two key design choices drive this robust performance in the long-sequence regime. First, by decomposing each LoRA update into fine-grained rank-1 components and enforcing a small, fixed activation budget, MoRA encourages each rank to specialize on a narrow subspace of the data manifold. At inference time, only the most relevant subspaces are activated for a given input, which preserves earlier task representations and prevents catastrophic interference. Second, our self-routed gating mechanism enables each rank to assess its own relevance on a per-token basis, yielding stable mixture patterns as the expert pool grows. Coupling with our proposed rank pruning, these mechanisms ensure that MoRA continually incorporates new knowledge only when needed while robustly maintaining prior capabilities.
Table 10: Comparisons of trainable parameters for each pre-trained weight matrix during continual learning of each task. MoE-LoRA and MoE-Adapter trains additional router module with LoRA experts. In MoRA, $k$ denotes the rank activation budget (with $k\leq r$).
Method Trainable parameters per task
LoRA [32] $r\,(d_{\text{in}}+d_{\text{out}})$
MoE-LoRA (1 expert/task) $r\,(d_{\text{in}}+d_{\text{out}})\;+\;d_{\text{in}}$
MoE-Adapter (2 experts/task) [85] $2(r\,(d_{\text{in}}+d_{\text{out}})+d_{\text{in}})$
CoDyRA [48] $r\,(d_{\text{in}}+d_{\text{out}})$
O-LoRA [74] $r\,(d_{\text{in}}+d_{\text{out}})$
LB-CL [58] $r\,(d_{\text{in}}+d_{\text{out}})+r$
MoRA $rd_{\text{in}}+kd_{\text{out}}$
B.3 Computation Cost
Table 10 summarizes the per-task trainable parameters of various continual-learning methods. Standard LoRA [32], CoDyRA [48], and O-LoRA [74] each introduce $r(d_{\text{in}}+d_{\text{out}})$ new parameters per weight matrix. Mixture-of-Experts variants such as MoE-LoRA and MoE-Adapter [85] additionally train a router module to controll the usage of each LoRA experts, inducing $d_{\text{in}}$ additional paramters for each experts. LB-CL [58] introduce $r$ additional parameters, mimic the singular values of SVD.
By contrast, MoRA requires only $rd_{\text{in}}+kd_{\text{out}}$ activated trainable parameters per task, where $k\leq r$ is the activation budget, and the trainable parameter is at most the same as a standard LoRA. Despite the small number of parameter activated and trained, MoRA achieves superior continual learning performance, and reaches comparable performance in general fine-tuning with only one-third of the activated parameters of a standard LoRA.
[…]
(a) Rank Activations of MoRA on data from Task 1 after learning Task 1.
[…]
(b) Rank Activations of MoRA on data from Task 1 after learning Task 2.
[…]
(c) Rank Activations of MoRA on data from Task 1 after learning Task 10.
Figure 5: Extended view of Fig. 3 illustrating forgetting mitigation. Regions corresponding to object semantics are highlighted with orange bounding boxes. Zoom in for details.
[…]
(a) Rank Activations of MoRA on data from Task 1 after learning Task 1.
[…]
(b) Rank Activations of MoRA on data from Task 9 after learning Task 1.
[…]
(c) Rank Activations of MoRA on data from Task 9 after learning Task 9.
Figure 6: Extended view of Fig. 3 illustrating knowledge reuse. Regions corresponding to generic input tokens (e.g. blue sky) are highlighted with blue bounding boxes. Zoom in for details.
B.4 Extended Visualizations of Rank Activations
In Sec.4.4 (Fig.3), we illustrated rank activations during the learning of Task 1 and Task 2. Here, we extend these visualizations to additional tasks and scenarios in Fig.5 and Fig.6.
MoRA retains task-specific semantics without forgetting. Fig. 5 shows the activation maps for the same Task 1 image after training on Task 1 (a), Task 2 (b), and the last Task 10 (c). Patches corresponding to the airplane object are outlined in orange. In all three snapshots, Rank at index 0 remains consistently and exclusively activated for those airplane patches, demonstrating that MoRA has effectively memorized the airplane-specific knowledge into Rank 0. Even after 10 subsequent tasks, this pattern remains unchanged, indicating that later updates do not overwrite or interfere with the learned airplane representations. In other words, MoRA effectively memorizes and preserves task-relevant semantics, thereby mitigating catastrophic forgetting.
MoRA encodes generic semantics that are reused across tasks. Fig. 6 examines an input image from Task 9 before and after learning Task 9. Panel (a) shows the activation map of data from Task 1 after learning Task 1: Rank 11 (outlined in blue) already responds strongly to sky-background patches, demonstrating that MoRA has stored a generic “blue sky” concept in this rank. In panel (b), when we infer on the Task 9 image before training on Task 9, Rank 11 is again activated for the sky regions, confirming that MoRA reuses this shared knowledge for unseen data. Finally, panel (c) shows the activation map after learning Task 9: Rank 11 remains dedicated to the sky background, while newly initialized ranks specialize in the “car” object semantics. This persistent reuse of Rank 11 across tasks illustrates MoRA’s ability to capture and retain common features as reusable memory slots, reducing redundancy and facilitating knowledge reuse.
Figure 7: Statistical analyses on the number of ranks required to capture 99% of cumulative sum (indicated in red dashed line) of all rank activations. Activations were gathered from the model after training on all tasks, and results are shown for a representative selection of layers and positions within the pre-trained model.
Figure 8: Statistical analyses on the number of ranks required to capture 99% of cumulative sum (indicated in red dashed line) of all rank activations. Activations were gathered from the model after training on all tasks, and results are shown for a representative selection of layers and positions within the pre-trained model.
(a) Vision Encoder
(b) Text Encoder
Figure 9: Required ranks to capture 99 % of cumulative activations, shown across different pre-trained model layers and projection locations.
B.5 Statistical Analyses of Contributing Rank Activations
Fig. 7 and Fig. 8 plot the cumulative sum of averaged rank activations after training on all tasks, sorted in descending order, for several representative layers and locations within pre-trained models. The red dashed line marks the point at which 99% of the total activation mass is reached, allowing us to quantify how many ranks are truly contributing to the model’s adaptation. Two key observations emerge:
1. Sparse Mixture: only a small subset of ranks is needed. Across all layers and positions, we find that fewer than 10% of the total ranks suffice to capture 99% of the activations. This highlights the extreme sparsity of MoRA’s self-activated mixture: most ranks remain dormant for any given input, while a compact set of highly relevant ranks drives the adaptation.
2. Adaptive Activation: The number of ranks required varies by layers and modules. The number of ranks needed to capture 99% of the cumulative activation mass varies across both layer depth and module type. For example, in Fig. 7, the MLP’s output projection (c-proj) in Layer 1 of the vision encoder requires 16 ranks, whereas the same module in Layer 1 of the text encoder needs only 3 ranks.
To provide a broader view, Fig. 9 shows the required rank counts for every module in the pre-trained model. We observe that most attention modules requires around 6–12 ranks, while the second MLP projection generally demands more ranks in early layers, peaking in the first few blocks, and then steadily declines in deeper layers.
Coupling the rank activation budget with rank pruning, MoRA adapts the number of ranks needed to activate at each layer and module. This adaptive sparsity maximizes the efficient use of newly acquired knowledge during continual learning.
Broader Impacts
MoRA’s self-activated, sparse mixture-of-ranks method makes continual adaptation of large vision–language and language models far more efficient and reliable. By updating only the most relevant low-rank subspaces, it reduce compute and memory requirements, enabling on-device personalization in domains like education, healthcare, and finance. This democratizes access to powerful, continually evolved models for smaller teams and resource-constrained settings, driving innovation and improving quality of life. Preserving pre-trained knowledge also ensures stability in safety-critical systems, such as medical diagnostics and autonomous vehicles, where forgetting of pre-trained capabilities can be dangerous.
The focus of this paper is fundamental research, and is broadly applicable to model fine-tuning techniques. Thus, it is possible that MoRA’s lightweight updates could be misused to insert stealthy backdoors or reinforce hidden biases that are hard to detect. Because it leaves most pre-trained parameters untouched, existing biases may persist or become entrenched. To mitigate these risks, we recommend strict audit logging of rank-level updates, anomaly detection on sparse changes, robust access controls, and routine bias and fairness assessments alongside any deployment.

### v2

@@
 As more experts accumulate, earlier task routing degrades, accelerating forgetting.
-We propose MoRA, a Mixture-of-Rank Adaptive learning approach with self-activated and sparse rank activation for CL.
-Unlike mixing multiple low-rank matrices, MoRA decomposes each rank-r update into r rank-1 components, each treated as an independent expert, enabling fine-grained mixture of rank-1 expert utilization while mitigating interference and redundancy.
-To avoid ambiguous routing, we propose that each rank-1 expert can infer its own relevance via intermediate activations.
+We propose MoRA, a Mixture-of-Rank Adaptive learning approaches with self-activated and sparse rank activation for CL.
+Unlike mixing multiple low-rank matrices, MoRA decomposes each rank-r update into r rank-one components, each treated as an independent expert, enabling fine-grained rank-one expert utilization while mitigating interference and redundancy.
+To avoid ambiguous routing, we propose that each rank-one expert can infer its own relevance via intermediate activations.
 Coupled with our proposed rank pruning and activation budgets, MoRA adaptively selects a sparse mixture of ranks per input.
-We validate MoRA on continual learning tasks with CLIP and large language models (LLMs), analyzing both in-domain learning and out-of-domain forgetting/generalization during fine-tuning.
-MoRA shows significant effectiveness on enhancing CL with PTMs, and improving generalization while mitigating forgetting.
+We validate MoRA on continual learning benchmarks using CLIP and language models, analyzing both in-domain learning and out-of-domain forgetting/generalization during fine-tuning.
+MoRA shows significant effectiveness in enhancing CL with PTMs, and improving generalization while mitigating forgetting.
 “Little by little, we gave you everything you ever dreamed of …” — “Little by little”, Oasis.
 1 Introduction
-Continual learning (CL) [29, 17, 73, 43, 21] aims to enable models to incrementally and efficiently acquire new knowledge from a stream of tasks, without catastrophic forgetting [55, 51] or the need for repeated fine-tuning on all previously seen data [77, 76, 85, 72].
-The emergence of large pre-trained models (PTMs), including Vision Transformer (ViT) for vision [22], Large Language Model (LLM) [28, 61], and CLIP for vision–language embeddings [60, 33], has spurred a wave exploring CL with PTMs[94, 52, 72, 35, 81, 59, 62, 74, 58].
-Despite the success of PTMs, CL on them across sequential tasks still faces key challenges, including preventing catastrophic forgetting and preserving—or even enhancing—the pre-trained capabilities to generalize to unseen or future cases [92, 85, 48, 58].
-A common approach to integrating new knowledge into pre-trained models (PTMs) is continual full fine-tuning [92, 27] or parameter-efficient fine-tuning (PEFT) using a small set of learnable parameters, such as Low-Rank Adaptation (LoRA) [32] (Fig.
+Continual learning (CL) (Hadsell et al., 2020; De Lange et al., 2021; Ding et al., 2022) aims to enable models to incrementally and efficiently acquire new knowledge from a stream of tasks, without catastrophic forgetting (Nguyen et al., 2019; McCloskey & Cohen, 1989) or the need for repeated fine-tuning on all previously seen data (Wang et al., 2022e; 2024).
+The emergence of large pre-trained models (PTMs), including Vision Transformer (ViT) for vision (Dosovitskiy et al., 2020), Language Model (LM) (Grattafiori et al., 2024; Raffel et al., 2020), and CLIP for vision–language embeddings (Radford et al., 2021), has spurred a wave exploring CL with PTMs (Xu et al., 2024; Wang et al., 2023).
+Despite the success of PTMs, CL on them across sequential tasks still faces key challenges, including preventing catastrophic forgetting and preserving, or even enhancing, the pre-trained capabilities to generalize to unseen or future cases (Qiao & Mahdavi, 2024).
+A common way to integrate new knowledge into PTMs is full fine-tuning (Zheng et al., 2023; Garg et al., 2023) or PEFT with a small set of learnable parameters (Yang et al., 2024b), such as LoRA (Hu et al., 2021) (Fig.
 1(a)).
-However, continually updating the same set of parameters often causes interference and forgetting, even with a single new task [7].
-Leveraging LoRA’s inherent regularization and flexibility, many CL methods for PTMs employ LoRA or its variants, along with specialized strategies to mitigate forgetting [72, 44, 74, 58, 44, 74].
-Inspired by mixture-of-experts (MoE) architectures [65, 40, 24, 16] and the modular nature of LoRA adapters, MoE frameworks using LoRA as experts (Fig.
-1(b)) have been widely adopted in CL [85, 72, 84, 13].
-These methods incrementally train new LoRA experts while freezing previous ones, using a router to compose experts per sample.
-This isolation reduces cross-task interference and forgetting, while the composability of experts supports knowledge reuse and transfer [72, 58, 64].
-In MoE-LoRA models, each expert comprises a full LoRA adapter with multiple ranks (i.e., subspaces), and expert selection occurs at this coarse-grained adapter level.
-Despite their success, this coarse granularity and the routing mechanism introduce three key challenges that limit effective utilization:
-(1) Interference across experts and infeasible knowledge reuse.
-For any given sample, an entire LoRA expert must be either fully activated or not at all, leading to two major issues: interference among incompatible ranks from different experts, and limited flexibility in knowledge reuse.
+However, continually updating the same parameters often leads to interference and forgetting, even with a single new task (Biderman et al., 2024).
+Leveraging LoRA’s regularization and flexibility, many CL methods adopt LoRA or its variants with additional strategies to mitigate forgetting (Yu et al., 2024; Wang et al., 2023; Qiao & Mahdavi, 2024).
+Inspired by mixture-of-experts (MoE) architectures (Shazeer et al., 2017; Lepikhin et al., 2020; Fedus et al., 2022; Dai et al., 2024) and the modular nature of LoRA adapters, MoE frameworks using LoRA adapters as experts have been studied (Dou et al., 2023; Wu et al., 2024b) and applied in CL (Fig.
+1(b)) have been widely adopted in CL (Yu et al., 2024; Wang et al., 2024; Yang et al., 2024a; Chen et al., 2024; Li et al., 2025).
+These works either predefine an MoE with LoRA for CL (Yang et al., 2024a), or incrementally add experts (Wang et al., 2024) or task-specific routers (Yu et al., 2024), assuming MoE benefits CL by isolating task interference.
+Such methods (Wang et al., 2024; Qiao & Mahdavi, 2024; Rusu et al., 2016; Yu et al., 2024) freeze old components and add new ones (e.g., experts or routers) to reduce forgetting.
+Despite design differences, we collectively refer a plain and general design with LoRA-based MoE as MoE-LoRA.
+In MoE-LoRA models, each expert is a LoRA adapter with multiple ranks (subspaces), and the router selects among experts with each LoRA adapter as a unit.
+Despite their success, we argue that the coarse granularity of experts and routing introduces three key challenges that limit effective utilization:
+(1) Interference across experts and limited knowledge reuse.
+A coarse-grained multi-rank LoRA expert is packed with a rich set of knowledge that must be fully activated or not at all.
+For each sample (i.e., token), any two experts may hold both complementary and conflicting knowledge, creating a dilemma: selecting both causes interference, while only selecting either one prevents reuse.
 (2) Learning redundancy and conflicts.
-Newly added experts may relearn redundant information or learn conflicting representation due to unnecessary activation of unrelated old ranks or insufficient reuse of relevant ones.
+As a result of (1), (newly added) experts often have to carry conflicting or redundant information, either from activating other experts with both useful and unnecessary content or from insufficient reuse of existing ones.
 This coarse selection exacerbates interference and forgetting during inference.
 (3) Routing ambiguity and forgetting.
-Redundancy and conflicts among experts make it hard for a LoRA-level router to assign overlapping or common inputs correctly, leading to suboptimal mixtures [40, 24, 16].
-As more new experts are accumulated in CL, the router’s assignments for earlier tasks become increasingly unreliable, makes it more prone to forgetting.
+With coarse-grained experts and routing, redundancy and conflicts hinder precise input–module assignment, leading to suboptimal mixtures (Lepikhin et al., 2020; Fedus et al., 2022; Dai et al., 2024).
+As more experts accumulate in CL, router assignments for earlier tasks grow unreliable, increasing forgetting.
+The three challenges are intertwined causes and effects of coarse-grained experts and routing, limiting the potential of promising MoE-LoRA design.
 Figure 1: Conceptional illustration of CL with (a) LoRA, (b) MoE-LoRA, and (c) MoRA (Ours).
 We propose MoRA, a Mixture-of-Rank Adaptive learning approaches with self-activated and sparsely selected ranks across the LoRAs added at different tasks.
-To address the limitations of coarse-grained MoE-LoRA methods, MoRA operates at a fine-grained level by treating each rank-one subspace within a LoRA adapter as an individual expert.
-Expert selection and routing are performed at the rank level, reducing interference from activating entire LoRA blocks.
-This decomposition promotes specialization on narrow regions of the data manifold, reduces redundancy, and enables effective reuse of shared knowledge across tasks.
-We introduce a self-activation mechanism to select a sparse set of ranks without relying on a separately trained router, avoiding routing confusion and further mitigating forgetting (Fig.
+To address the limitations of coarse-grained MoE-LoRA methods, MoRA operates at a fine-grained level by treating each rank-one subspace within a LoRA adapter as an individual expert (or treating each rank-one LoRA as an expert).
+Expert selection and routing occur at the rank level, avoiding interference from whole LoRA blocks.
+This decomposition promotes specialization on narrower data regions, reduces redundancy, and enables effective knowledge reuse across tasks.
+We introduce a self-activation mechanism to select a sparse set of ranks without relying on a separately trained router.
+By defining each rank-one component as an expert, we avoid a separate router for expert assignment: the rank-one $\mathbf{A}$ in LoRA can act as a weighting router for its $\mathbf{B}$.
+Each rank evaluates its relevance to the input and determines its contribution, grounded in the view of low-rank updates as linear key–value memory.
+This design unifies rank-one LoRA with fine-grained MoE while preserving the original characteristics, saving parameters (for additional router), reducing routing confusion, and further mitigating forgetting (Fig.
 1(c)).
-Each rank evaluates its own relevance to the input and determines its contribution, grounded in the view of low-rank updates as linear key–value memory.
 Sparsity is enforced via temperature-scaled softmax together with top-k masking and rank pruning under a rank activation budget.
-This design introduces no additional computation and eliminates the overhead associated with extra routing.
-Leveraging the connection between PTMs and low-dimensional manifolds [41, 1], MoRA enables precise, input-specific adaptation at test time through selective subspace updates.
-With fine-grained rank-level experts and sparse activation, MoRA facilitates continual learning by reusing relevant old ranks and selectively activating a few new ones per sample, enabling efficient, “little-by-little” adaptation that minimizes redundancy and supports essential updates.
-This selective, sample-specific activation at inference also enhances retention of pre-trained knowledge and improves generalization to unseen or future cases.
-We apply and validate MoRA on CL with both vision-language CLIP and LLMs, as well as analyzing the effects on general pre-trained knowledge during low-rank fine-tuning.
-Across all settings, MoRA achieves strong downstream performance with significantly fewer active parameters, substantially reduces forgetting of both prior tasks and pre-trained knowledge, and even enhances generalization to unseen domains.
+As a complement, test-time thresholding further improves adaptability across cases.
+Leveraging the connection between PTMs and low-dimensional manifolds (Li et al., 2018; Aghajanyan et al., 2020), MoRA enables precise, input-specific adaptation through selective subspace updates.
+With fine-grained rank-level experts and sparse activation, MoRA enables continual learning by reusing old ranks and selectively activating a few new ones per sample, allowing efficient “little-by-little” adaptation with minimal redundancy.
+This sample-specific activation at inference also preserves pre-trained knowledge and improves generalization to unseen tasks.
+We validate MoRA on CL with both vision–language CLIP and LMs, analyzing its impact on pre-trained knowledge during low-rank fine-tuning.
+MoRA delivers strong downstream performance with far fewer active parameters, greatly reduces forgetting of prior tasks and pre-trained knowledge, and even improves generalization to unseen domains.
 2 Related Work
 Continual learning enables sequential knowledge acquisition without forgetting.
-Experience replay (ER) methods [49, 4, 10, 45, 11, 82, 83] interleave past examples with new data.
-Parameter regularization [36, 2, 86, 3, 34] penalizes updates to critical weights.
-Dynamic networks [70, 71, 93, 72, 52, 44, 77, 76, 66, 75, 72] allocate new capacity on the fly and preserve dedicated pathways for prior tasks.
+Experience replay (ER) methods (Luo et al., 2023; Aljundi et al., 2019b; Chaudhry et al., 2018a; Liu et al., 2020; Chaudhry et al., 2018b; Yan et al., 2022; 2021) interleave past examples with new data.
+Parameter regularization (Kirkpatrick et al., 2017; Aljundi et al., 2018; Zenke et al., 2017; Aljundi et al., 2019a; Jha et al., 2023) penalizes updates to critical weights.
+Dynamic networks (Zhou et al., 2024; Wang et al., 2022a; b; Zhou et al., 2022; Wang et al., 2024; McDonnell et al., 2024; Liang & Li, 2024; Wang et al., 2022e; d; Smith et al., 2023; Wang et al., 2022c; 2024) allocate new capacity on the fly and preserve dedicated pathways for prior tasks.
 Continual learning of PTMs.
-For CL on vision–language CLIP model [35, 89, 27], methods like ZSCL [92] retain zero-shot performance during adaptation, and follow-up work [85, 67, 81, 48] continually fine-tunes while leveraging frozen pre-trained predictions.
-The X-TAIL benchmark [81] further challenges models by mixing domain labels at test time.
-In large language models (LLMs) [18, 59, 62, 74, 58], continual learning uses capacity expansion or task-specific submodules to reduce interference.
-In this work, we also study how fine-tuning impacts pre-trained general knowledge.
-Low-rank adaptation (LoRA) [32] is widely used for parameter-efficient fine-tuning of large pre-trained models.
-Building on this foundation, recent methods reformulate LoRA’s updates via SVD-based initialization and dynamic rank scheduling [87, 53, 20, 88, 46, 79], showing that task adaptation primarily lies in a compact subspace.
-In this work, we offer a complementary perspective by viewing both pre-trained weight matrix and its low-rank updates as a linear associative memory [41, 1, 37, 5].
-Under this lens, a rank-$r$ update corresponds to writing $r$ new key–value entries into the pre-trained memory matrix, with each rank-1 component acting as an independent memory slot.
-Mixture-of-experts (MoE) in LoRA fine-tuning.
-MoE increases model capacity by replacing the Transformer’s feed-forward layer with expert subnetworks, routing each input to a small subset and using load-balancing losses to even out usage [65, 40, 24, 16].
-This paradigm has been adapted for standard fine-tuning [14, 42, 23] and CL [85, 72, 84, 13] with LoRA, treating each low-rank adapter as an expert: new adapters are trained and frozen per task to prevent forgetting.
-In contrast, we decompose each rank-$r$ update into $r$ rank-$1$ components and compute an input-dependent mixture over these fine-grained experts, greatly enhancing both expert specialization and mixture diversity.
+For CL on vision–language CLIP model (Jha et al., 2024; Zhang et al., 2024b; Garg et al., 2023), methods like ZSCL (Zheng et al., 2023) retain zero-shot performance during adaptation, and follow-up work (Yu et al., 2024; Tang et al., 2025; Xu et al., 2024; Lu et al., 2024) continually fine-tunes while leveraging frozen pre-trained predictions.
+The X-TAIL benchmark (Xu et al., 2024) further challenges models by mixing domain labels at test time.
+In language models (LMs) (de Masson D’Autume et al., 2019; Qin & Joty, 2021; Razdaibiedina et al., 2023; Wang et al., 2023; Qiao & Mahdavi, 2024), continual learning uses capacity expansion or task-specific submodules to reduce interference.
+Low‐rank adaptation (LoRA) (Hu et al., 2021) is widely used for parameter‐efficient fine‐tuning of large pre‐trained models.
+Building on this foundation, recent methods reformulate LoRA’s updates via SVD‐based initialization and dynamic rank scheduling (Zhang et al., 2024a; Meng et al., 2024; Ding et al., 2023; Zhang et al., 2023; Liu et al., 2024; Wu et al., 2024a), showing that task adaptation primarily lies in a compact subspace.
+In this work, we offer a complementary perspective by viewing both pre‐trained weight matrix and its low‐rank updates as a linear associative memory (Li et al., 2018; Aghajanyan et al., 2020; Kohonen, 1972; Anderson, 1972).
+Under this lens, a rank-$r$ update corresponds to writing $r$ new key–value entries into the pre-trained memory matrix, with each rank-one component acting as an independent memory slot.
+Mixture‐of‐experts (MoE) in LoRA fine‐tuning.
+MoE increases model capacity by replacing the Transformer’s feed-forward layer with expert subnetworks, routing each input to a small subset and using load‐balancing losses to even out usage (Shazeer et al., 2017; Lepikhin et al., 2020; Fedus et al., 2022; Dai et al., 2024).
+This paradigm has been adapted for standard fine-tuning (Chen et al., 2023; Li et al., 2024; Dou et al., 2023) and CL (Yu et al., 2024; Wang et al., 2024; Yang et al., 2024a; Chen et al., 2024) with LoRA, treating each low‐rank adapter as an expert: new adapters are trained and frozen per task to prevent forgetting.
+In contrast, we decompose each rank‐$r$ update into $r$ rank‐11 components and compute an input‐dependent mixture over these fine‐grained experts, greatly enhancing both expert specialization and mixture diversity.
 3 Methods
@@
 Continual learning.
-In continual learning, a model must sequentially learn from $T$ distinct tasks.
-For task $t\in\{1,\dots,T\}$, we denote its dataset by $\mathcal{D}^{t}=\{(\mathbf{x}^{t}_{i},y^{t}_{i})\}_{i=1}^{N^{t}}$, where $\mathbf{x}^{t}_{i}\in\mathbb{R}^{H\times W\times C}$ is the $i$th input sample, $y^{t}_{i}\in\mathcal{C}^{t}$ is its desired prediction, and $N^{t}$ is the number of examples in task $t$.
-At each time step $t$, the model may only access $\mathcal{D}^{t}$ and is prohibited from storing or replaying data from any $\mathcal{D}^{u}$ with $u<t$.
-This constraint makes it challenging to avoid catastrophic forgetting, which updates on new tasks overwrite knowledge acquired on earlier ones.
+In CL, a model sequentially learns $T$ tasks.
+For task $t\in{1,\dots,T}$, let $\mathcal{D}^{t}={(\mathbf{x}^{t}_{i},y^{t}_{i})}_{i=1}^{N^{t}}$, where $\mathbf{x}^{t}_{i}\in\mathbb{R}^{n\times d}$, $y^{t}_{i}\in\mathcal{C}^{t}$, and $N^{t}$ is the number of examples.
+In the memory-free setting, the model may access only $\mathcal{D}^{t}$ and cannot access data from any $\mathcal{D}^{u}$ with $u<t$.
 Low-rank adaptation.
-LoRA [32] parameterizes a low-rank update to a pre-trained weight matrix $\mathbf{W}_{0}\in\mathbb{R}^{d_{out}\times d_{in}}$ by introducing two factors $\mathbf{B}\in\mathbb{R}^{d_{out}\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times d_{in}}$, such that $\Delta\mathbf{W}=\mathbf{B}\mathbf{A}$, where $r\ll\min(d_{in},d_{out})$.
+LoRA (Hu et al., 2021) parameterizes a low‐rank update to a pre‐trained weight matrix $\mathbf{W}_{0}\in\mathbb{R}^{d_{out}\times d_{in}}$ by introducing two factors $\mathbf{B}\in\mathbb{R}^{d_{out}\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times d_{in}}$, such that $\Delta\mathbf{W}=\mathbf{B}\mathbf{A}$, where $r\ll\min(d_{in},d_{out})$.
 The updated weight matrix is then defined as:
@@
 In this formulation, the original weights $\mathbf{W}_{0}$ remain fixed, and only $\mathbf{B}$ and $\mathbf{A}$ are trained, reducing the number of trainable parameters from $d_{in}d_{out}$ to $r(d_{in}+d_{out})$.
-Mixture-of-Experts LoRA.
-Building on the Mixture-of-Experts (MoE) paradigm, a common framework of MoE-LoRA [85, 72] treats each LoRA as an independent expert and incrementally adds a new LoRA expert for each incoming task.
-After $T$ tasks, we have $T$ LoRA experts $\{(\mathbf{A}^{1},\mathbf{B}^{1}),\dots,(\mathbf{A}^{T},\mathbf{B}^{T})\}$.
+Mixture‐of‐Experts LoRA.
+Building on the Mixture‐of‐Experts (MoE) paradigm, a common generic framework of MoE‐LoRA (Yu et al., 2024; Wang et al., 2024) treats each LoRA as an independent expert.
+Suppose after $T$ tasks, we have $T$ LoRA experts $\{(\mathbf{A}^{1},\mathbf{B}^{1}),\dots,(\mathbf{A}^{T},\mathbf{B}^{T})\}$.
 For an input token $x\in\mathbb{R}^{d_{\rm in}}$, the overall LoRA update in this framework is given by $\Delta\mathbf{W}=\sum\nolimits_{i=1}^{T}R(x)_{i}\mathbf{B}^{i}\,\mathbf{A}^{i},$ where the mixture weight $R(x)\in\mathbb{R}^{T}$ is produced by a learned router $R(\cdot)=\mathrm{softmax}(x\mathbf{W}_{r})$ and $\mathbf{W}_{r}\in\mathbb{R}^{d_{\rm in}\times T}$ contains the router’s trainable parameters.
-Each LoRA’s contribution is weighted by the learnable router, enabling the model to dynamically select and combine the most relevant low-rank updates for each token.
+Each LoRA’s contribution is weighted by the learnable router, enabling the model to dynamically select and combine the most relevant low‐rank updates for each token.
 In practice, a Top-$k$ masking is typically applied to mixture weights to enforce sparsity, activating only the $k$ most relevant experts and controlling computational overhead.
+3.2 Model Weights and Updates as Linear Associative Memory Model
+Weight matrix as linear associative memory.
+Large pre‐trained models (PTMs) have been shown to inhabit a low “intrinsic‐dimension” manifold compared to their full parameter count (Li et al., 2018; Aghajanyan et al., 2020).
+Any weight matrix of a pre-trained model $\mathbf{W}_{0}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$ can be treated as a linear associative memory model with a key-value structure (Kohonen, 1972; Anderson, 1972; Bau et al., 2020; Meng et al., 2022).
+Concretely, for any matrix $\mathbf{W}\in\mathbb{R}^{d_{\mathrm{out}}\times d_{\mathrm{in}}}$ and input $\mathbf{x}\in\mathbb{R}^{d_{\mathrm{in}}}$, the product $\mathbf{W}\,\mathbf{x}$ implements a content‐addressable memory read over $m$ stored key–value pairs $\{(\mathbf{K}_{i,:},\mathbf{V}_{:,i})\}_{i=1}^{m}$, with $m$ potentially large and varied across each weight matrix.
+Here, ${\mathbf{K}}=\begin{bmatrix}{\mathbf{K}}_{1,:}^{\top}&\cdots&{\mathbf{K}}_{m,:}^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{m\times d_{\text{in}}}$ and $\mathbf{V}$$\begin{bmatrix}{\mathbf{V}}_{:,1}&\cdots&{\mathbf{V}}_{:,m}\end{bmatrix}\in\mathbb{R}^{d_{\text{out}}\times m}$ with $\mathbf{K}_{i,:}\in\mathbb{R}^{1\times d_{\text{in}}}$ is the $i$-th row of $\mathbf{K}$} and ${\mathbf{V}}_{:,i}\in\mathbb{R}^{d_{\text{out}}\times 1}$ is the $i$-th column of $\mathbf{V}$}.
+Given any input $\mathbf{x}$, the output $\mathbf{y}\in\mathbb{R}^{d_{\text{out}}}$ is computed as
+$\mathbf{y}=\mathbf{W}\mathbf{x}\approx\sum\nolimits_{i=1}^{m}\mathbf{V}_{:,i}(\mathbf{K}_{i,:}\mathbf{x}).$ (2)
+In this formulation, the forward pass through the weight matrix with the input decomposes into two steps: (1) computing response scores or assignment weights via the key vectors $\mathbf{K}_{i,:}$, and (2) aggregating the corresponding value vectors $\mathbf{V}_{:,i}$ according to those weights.
 Figure 2: Overview of MoRA.
 For each new task, we freeze the ranks learned on previous tasks and introduce $r$ new ranks of updates.
-Our sparse self-activated mixture-of-ranks framework jointly considers all old and new ranks, adaptively inferring a sparse mixture weight for each rank.
+Our sparse self-activated mixture‐of‐ranks framework jointly considers all old and new ranks, adaptively inferring a sparse mixture weight for each rank.
 Panels (a,c) illustrate MoRA conceptually and (b,d) detail its computation for tasks $t$ and $t+1$, respectively.
-3.2 Pre-trained Weight Matrices and Low-Rank Updates as Linear Associative Memory
-Large PTMs have been shown to lie on a low “intrinsic dimension” manifold [41, 1].
-From this perspective, the pre-trained weight matrix $\mathbf{W}_{0}\in\mathbb{R}^{d_{\mathrm{out}}\times d_{\mathrm{in}}}$ can itself be viewed as a linear associative memory [37, 5], having stored key–value pair of memories drawn from vast pre-training data.
-Specifically, any rank-$r$ update $\Delta\mathbf{W}\;=\;\mathbf{B}\,\mathbf{A}$ can be viewed as inserting or updating $r$ key–value entries into the pre-trained weight matrix.
-Concretely, let
-$\mathbf{A}=\begin{bmatrix}\mathbf{A}_{1}^{\top}&\cdots&\mathbf{A}_{r}^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{r\times d_{\mathrm{in}}},\quad\mathbf{B}=\begin{bmatrix}\mathbf{B}_{1}&\cdots&\mathbf{B}_{r}\end{bmatrix}\in\mathbb{R}^{d_{\mathrm{out}}\times r},$ (2)
-where each $\mathbf{A}_{i}\in\mathbb{R}^{1\times d_{\mathrm{in}}}$ serves as a key and each $\mathbf{B}_{i}\in\mathbb{R}^{d_{\mathrm{out}}\times 1}$ as its corresponding value. Then
-$\Delta\mathbf{W}=\sum\nolimits_{i=1}^{r}\mathbf{B}_{i}\,\mathbf{A}_{i}$ (3)
-writes the value $\mathbf{B}_{i}$ into the memory slot addressed by key $\mathbf{A}_{i}$.
-In this associative-memory view, each rank-$i$ component explicitly encodes “what to store” (the value $\mathbf{B}_{i}$) and “how to index it” (the key $\mathbf{A}_{i}$), thereby enriching the base model with newly acquired knowledge.
-More details are left in Appendix A.
+Parameter update as modifications to the linear associative memory.
+Building on the associative-memory view of weights, we can likewise interpret weight updates during training as key–value insertions and modifications into the existing memory.
+During training, for an arbitrary update $\Delta\mathbf{W}$ on a given $\mathbf{W}_{0}$ as the starting point, the weight update process can be represented as
+$\mathbf{W}_{\text{new}}=\mathbf{W}_{0}+\Delta\mathbf{W},$ (3)
+where $\mathbf{W}_{0}$ can be the pre-trained model weights.
+We can also formulate the weight updates $\Delta\mathbf{W}$ with the key-value memory model $\{\Delta\mathbf{K},\Delta\mathbf{V}\}$ as an update to the original memory model.
+With the operation on an input $\mathbf{x}$ via
+$\Delta\mathbf{W}\mathbf{x}=\Delta\mathbf{V}\Delta\mathbf{K}\mathbf{x},$ (4)
+the updates on weights can be seen as a retrievable key-value system for the updating of $\mathbf{W}\mathbf{x}$.
+LoRA as low-rank key–value memory updates.
+LoRA (Hu et al., 2021) performs parameter‐efficient fine‐tuning by adding a rank‐$r$ update $\Delta\mathbf{W}=\mathbf{B}\mathbf{A}$ to the pre‐trained weight matrix $\mathbf{W}_{0}$, where $\mathbf{B}\in\mathbb{R}^{d_{\text{out}}\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times d_{in}}$.
+With the model update based on LoRA, given an input $\mathbf{x}$, the computation of updated weight becomes
+$\mathbf{y}_{\text{new}}=\mathbf{W}_{\text{new}}\mathbf{x}=\mathbf{W}_{0}\mathbf{x}+\Delta\mathbf{W}\mathbf{x}=\mathbf{W}_{0}\mathbf{x}+\mathbf{B}\mathbf{A}\mathbf{x}.$ (5)
+Decomposing the low-rank matrices into a collection of its $r$ rank-one components gives
+$\Delta\mathbf{W}\mathbf{x}=\mathbf{B}\mathbf{A}\mathbf{x}=\sum\nolimits_{i=1}^{r}\mathbf{B}_{:,i}(\mathbf{A}_{i,:}\mathbf{x}),$ (6)
+with $\mathbf{A}=\begin{bmatrix}\mathbf{A}_{1,:}^{\top}&\cdots&\mathbf{A}_{r,:}^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{r\times d_{\text{in}}}$ where $\mathbf{A}_{i,:}$ is the $i$-th row of $\mathbf{A}$, and $\mathbf{B}=\begin{bmatrix}\mathbf{B}_{:,1}&\cdots&\mathbf{B}_{:,r}\end{bmatrix}\in\mathbb{R}^{d_{\text{out}}\times r}$ where $\mathbf{B}_{:,i}$ is the $i$-th column of $\mathbf{B}$.
+Considering the connection between the formulation of LoRA Eq.
+(1) and associative‐memory view (Anderson, 1972; Meng et al., 2022) of weight update Eq.
+(4), we can intuitively interpret LoRA as obtaining a small key–value store in a low‐rank subspace, where the rows of $\mathbf{A}$ serve as keys and the columns of $\mathbf{B}$ serve as their corresponding values.
+After training the LoRA module on a base weight matrix $\mathbf{W}_{0}$, the modified forward pass in Eq.
+(5) with Eq.
+(6) can be seen as: first, the original pre‐trained computation $\mathbf{W}_{0}\,\mathbf{x}$ (Eq.
+(2)) is performed; then, each learned key $\mathbf{A}_{i,:}$ measures its relevance to the input $\mathbf{x}$ via the product $\mathbf{A}_{i,:}\mathbf{x}$, and each learned value $\mathbf{B}_{:,i}$ is scaled by this relevance score and aggregated, in a same way as shown in Eq.
+(2).
+In this way, LoRA effectively inserts or updates a small set of key–value memories that augment the pre‐trained model’s capabilities.
 3.3 Mixture of Ranks Adaptation
-To overcome the coarse granularity of MoE-LoRA, we introduce mixture-of-ranks, applying MoE at the rank-level subspace within all LoRA adapters.
-By treating each rank as an independent expert, our fine-grained mixture framework enables: 1.
-Reduced redundancy and shared knowledge reuse: encoding common, cross-task features (e.g., generic semantics) can be reused, preventing redundant relearning.
-2.
-Minimized interference and enhanced specialization: focusing exclusively on task-specific patterns, ensuring that new knowledge is learned without disrupting existing representations or interfered with future updates.
-For any given input token $x$ from task $t$, we weight each rank’s contribution according to its relevance. We introduce a scalar mixture weight $w_{i}\in\mathbb{R}$ for each rank, yielding
-$\Delta\mathbf{W}^{t}=\sum\nolimits_{i=1}^{r_{t}}w_{i}\mathbf{B}_{i}\mathbf{A}_{i},$ (4)
-where each rank-1 update is analogous to an independent expert, $r_{t}$ is the total number of ranks learned up to task $t$, and $w_{i}$ modulates whether and how strongly the corresponding rank contributes to the final updated weight matrix.
-Unlike prior work that treats the entire rank-$r$ update as one expert [85, 72], this formulation enables fine-grained utilization of each rank’s update.
+In the key–value associative‐memory view of LoRA (Eq.
+(5) and Eq.
+(6)), a standard LoRA or its variants (Hu et al., 2021; Meng et al., 2024; Zhang et al., 2024a) apply a dense update: every rank’s subspace contributes to $\Delta\mathbf{W}\,\mathbf{x}$ for any input $\mathbf{x}$.
+Even in MoE-LoRA, where each LoRA module is treated as a separate expert, all key–value pairs within a LoRA module are typically used densely as a whole (either fully activated or not), which creates redundancy across experts and underutilizes the expressive potential of the key vectors $\mathbf{A}$ to indicate input relevance.
+Rather than by default activating all key–value memories for every input, many of which may encode patterns irrelevant to the current token, we introduce a gated, sparse mixture-of-ranks update that selectively activates only the most relevant rank-one key–value pairs.
+Concretely, we associate each rank-one key–value pair with a scalar weight ${\textnormal{w}}_{i}\in\mathbb{R}$.
+For an input token $\mathbf{x}\in\mathbb{R}^{d_{\mathrm{in}}}$ from task $t$, with $r_{t}$ total ranks, the update becomes
+$\Delta\mathbf{W}^{t}=\sum\nolimits_{i=1}^{r_{t}}{\textnormal{w}}_{i}\mathbf{B}_{:,i}\mathbf{A}_{i,:},{}$ (7)
+where ${\textnormal{w}}_{i}$ modulates whether and how strongly each rank $i$ contributes to the updated weight matrix.
 3.4 Self-Activated Adaptive Sparse Mixture of Ranks
-Another key challenge in MoE and continual learning is the routing ambiguity and forgetting.
-To this end, we explore whether each rank-1 expert can autonomously infer its relevance to the current input, eliminating the need for a separate router and enabling self-activated expert selection.
-Self-activated sparse mixture of ranks.
-Rather than relying on an external router, our framework derives mixture weights for each rank directly from the activations of each rank.
+A key challenge for MoE-style CL is routing ambiguity, leading to catastrophic forgetting.
+Conventional routers assign examples to coarse-grained experts, but adding new experts often makes routing suboptimal, causing over-activation or underuse, i.e., the three challenges discussed in Sec.
+1.
+To avoid these pitfalls, we treat each rank-one LoRA (or rank) as an expert based on the rank-level mixture model (Eq.
+(7)), enabling sparse activation of fine-grained experts for specific tasks.
+Thus, CL with incrementally added rank-one experts becomes a process of accumulating atomic key–value knowledge that can be reused in future tasks.
+Unlike conventional MoE-LoRA (Yang et al., 2024a; Wang et al., 2024; Yu et al., 2024), we use the rank-one $\mathbf{A}_{i}$’s (keys) as routers to their corresponding $\mathbf{B}_{i}$’s (values) through self-activation, with each $\mathbf{A}_{i}$$\mathbf{B}_{i}$ pair functioning as a router-expert pair.
+Mixture of self-activated rank-one experts.
+Rather than relying on an external router, our framework derives mixture weights for each rank directly from that rank’s activation.
 Recall from Eq.
-(3) that each rank-$i$ update is parameterized by a key–value pair $(\mathbf{A}_{i},\mathbf{B}_{i})$.
-Given an input token $x\in\mathbb{R}^{d_{\mathrm{in}}}$ from task $t$, by treating $\mathbf{A}$ as the set of keys, we compute a raw score for all $r_{t}$ ranks accumulated up to task $t$:
-$s_{i}\;=\;\frac{\mathbf{A}_{i}x}{\sqrt{\sum_{j=1}^{r_{t}}\|\mathbf{A}_{j}x\|_{2}^{2}}}\quad\in\mathbb{R},$ (5)
-where the $\ell_{2}$–normalization over all $r_{t}$ ranks ensures numerical stability and compensates for variations in activation scales between layers and pre-trained weights.
-As shown in Table 5, using these raw scores alone matches or even exceeds the performance of a learned router for each rank, showing the feasibility that each rank can infer its own relevance using its activations.
-To encourage rank specialization and focus the update on the most relevant ranks, we temperature-scale and normalize the scores:
-$w_{i}\;=\;\mathrm{softmax}\!\bigl{(}\tfrac{s}{\tau_{\text{MoRA}}}\bigr{)}_{i},$ (6)
-where the softmax is taken over the vector $s=[s_{1},\dots,s_{r_{t}}]$ and $\tau_{\text{MoRA}}$ is a temperature hyperparameter that controls the sharpness of the distribution.
-To ensure only a sparse, specialized subset of ranks contributes, we introduce an activation budget that caps the number of ranks that can be activated.
-Adaptive sparse mixture of ranks.
-To ensure a stable and sparse activation pattern, we impose a fixed sparse rank-activation budget via top-k masking on the $\ell_{2}$–normalized scores $s_{i}:=\mathrm{TopK}(s,\,k),$ where
-$[\mathrm{TopK}(s,k)]_{i}=\begin{cases}s_{i},&\text{if }s_{i}\text{ is among the top }k\text{ entries of }s,\\ -\infty,&\text{otherwise,}\end{cases}$ (7)
-ensuring a sparse subset of at most $k$ out of the total $r_{t}$ ranks to be activated for task $t$ during CL.
-We further refine the mixture weights by applying a fixed threshold $\delta$ to the $\ell_{2}$–normalized scores $s_{i}$.
-Concretely, after computing the softmax weights $w_{i}$, we define a binary mask $m_{i}=\mathbbm{1}\{s_{i}\geq\delta\}$, and update each gate via $w_{i}:=m_{i}\,\odot\,w_{i}$, where $\mathbbm{1}\{\cdot\}$ is an indicator function and $\odot$ denotes element-wise multiplication.
-This mask zeroes out any rank whose normalized activation falls below the threshold, ensuring that only the most relevant ranks contribute, resulting in a sparse mixture of an adaptive number of ranks based on the input.
+(6) that each rank-one update is parameterized by a key–value memory pair $(\mathbf{A}_{i,:},\mathbf{B}_{:,i})_{i=1}^{r_{t}}$; the key $\mathbf{A}_{i,:}$ therefore provides a natural, input-dependent relevance signal.
+Specifically, given an input token $x\in\mathbb{R}^{d_{\mathrm{in}}}$ from task $t$, and treating $\mathbf{A}$ as the set of keys, we compute a raw score for all $r_{t}$ ranks accumulated up to task $t$:
+${\textnormal{s}}_{i}\;=\;\frac{\mathbf{A}_{i}x}{\sqrt{\sum_{j=1}^{r_{t}}\|\mathbf{A}_{j}x\|_{2}^{2}}}\quad\in\mathbb{R},$ (8)
+where the $\ell_{2}$-normalization across all $r_{t}$ ranks ensures numerical stability and compensates for activation-scale variations between layers and pre-trained weights.
+Using raw scores alone matches or surpasses a separately learned router (with 1.5 times parameters) without self-activation (Table 4), showing that each rank can infer its own relevance from activation.
+Self-activated rank selection.
+To ensure a stable and gated, sparse activation pattern, we impose a fixed sparse rank‐activation budget via top-k masking on the $\ell_{2}$–normalized scores ${\textnormal{s}}_{i}:=\mathrm{TopK}({\textnormal{s}},\,k),$ where
+[…]
+ensuring that at most $k$ out of the total $r_{t}$ ranks are eligible to be activated for task $t$.
+This encourages rank specialization: only a small set of the most relevant ranks are emphasized and trained to capture each kind of input-specific dynamics, and it prevents tiny, noisy activations from affecting gradients and learning dynamics.
+Sharpness control on rank activations. To further encourage rank specialization and concentrate the update on the most relevant ranks, we temperature-scale and normalize the scores:
+${\textnormal{w}}_{i}\;=\;\mathrm{softmax}\!\bigl(\tfrac{s}{\tau_{\text{MoRA}}}\bigr)_{i},$ (10)
+where the softmax is taken over the vector ${\textnormal{s}}=[{\textnormal{s}}_{1},\dots,{\textnormal{s}}_{r_{t}}]$ and $\tau_{\text{MoRA}}$ controls the distribution’s sharpness.
+A lower $\tau_{\text{MoRA}}$ concentrates mass on a few ranks (specialists), sharpens the distribution and routes larger learning signals to top ranks, which accelerates specialization and reduces noisy updates to low-relevance ranks, while a higher $\tau_{\text{MoRA}}$ spreads mass (shared representations).
+Threshold-based test-time rank selection.
+To further suppress low-signal ranks at inference, we optionally apply a threshold $\delta$ to the $\ell_{2}$–normalized raw scores and zero out gates below it.
+We do not hard-threshold during training such that gradients can still flow to near-threshold ranks and allow exploration and adaptation; thresholding only at test time reduces noise and runtime cost:
+${\textnormal{w}}_{i}:=\mathbbm{1}\{{\textnormal{s}}_{i}\geq\delta\}\odot{\textnormal{w}}_{i}.$ (11)
+This yields a sparse, input-dependent mixture of only the most significant rank-one experts.
+Our fine-grained mixture framework offers two key advantages: (1) It reduces redundancy and enables shared-knowledge reuse: ranks that capture common, cross-task features (such as generic semantics) can be reactivated across multiple tasks, preventing those patterns from being relearned redundantly or interfered; (2) It minimizes routing ambiguity and prevents forgetting: each rank can specialize on more nuanced patterns, yielding an optimized, rank‐wise mixture that integrates new knowledge without disrupting existing representations or hindering future updates.
 Incremental mixture of ranks.
-Unlike methods that add a fixed number of LoRA experts per task [85], MoRA incrementally adds $r$ new ranks for each task, freezes all prior ranks, and adaptively selects a sparse subset of the most relevant ranks during both training and inference.
+Unlike methods that add a fixed number of LoRA experts per task (Yu et al., 2024), MoRA incrementally adds $r$ new ranks for each task, freezes all prior ranks, and adaptively selects a sparse subset of the most relevant ranks during both training and inference.
 This design seamlessly incorporates new task-specific knowledge into freshly initialized ranks while preserving earlier ranks unchanged to prevent forgetting.
@@
 Training Objectives.
-In our experiments, we only use the model’s standard training objective, without any extra regularization or load-balancing constraints.
-In practice, we empirically observe that the conventional load-balancing loss widely employed in the Mixture-of-Experts literature [65, 40, 24, 16] is not necessary, thanks to our fine-grained self-activated gating design.
+In our experiments, we only use the model’s standard training objective, without any extra regularization or load‐balancing constraints.
+In practice, we empirically observe that the conventional load-balancing loss widely employed in the Mixture-of-Experts literature (Shazeer et al., 2017; Lepikhin et al., 2020; Fedus et al., 2022; Dai et al., 2024) is not necessary, thanks to our fine-grained self-activated gating design.
 Imposing a strict load-balancing constraint might hinder the learning of the key vectors (i.e., matrix $\mathbf{A}$), since each rank should remain free to capture its own characteristic input distribution.
@@
 Table 1: Comparisons on X-TAIL for each domain in terms of “Transfer”, “Average”, and “Last” scores (%). The best and the second best results are highlighted in red and blue, respectively.
-    Method Aircraft [50] Caltech101 [25] DTD [15] EuroSAT [30] Flowers [56] Food [8] MNIST [19] OxfordPet [57] Cars [38] SUN397 [80] Average
+Method
+Aircraft Caltech DTD EuroSAT Flowers Food MNIST OxPet Cars SUN397 Average
 CLIP
    Zero-shot 23.5 76.8 37.3 36.7 63.6 84.0 46.7 86.7 66.1 63.7 58.5
+   Fine-tune 39.6 84.7 70.0 94.7 97.0 85.8 97.6 93.4 81.0 74.7 81.9
 Transfer
-   Zero-shot [60] 76.8 37.3 36.7 63.6 84.0 46.7 86.7 66.1 63.7 62.4
-   LwF [43] 66.6 26.9 19.5 51.0 78.4 26.6 68.9 35.5 56.1 47.7
-   WiSE-FT [78] 70.1 31.9 25.3 56.3 79.8 29.9 74.9 45.6 56.8 52.3
-   iCaRL [63] 71.7 35.0 43.0 63.4 86.9 43.9 87.8 63.7 60.0 61.7
-   ZSCL [92] 73.3 32.6 36.8 62.1 83.8 42.1 83.6 56.5 60.2 59.0
-   MoE-Adapter [85] 71.0 34.9 19.2 63.0 86.6 20.0 87.2 63.7 58.6 56.0
-   RAIL-Primal [81] 76.8 37.3 36.7 63.6 84.0 46.7 86.7 66.1 63.7 62.4
-   CoDyRA [48] 74.3 36.8 44.2 69.9 83.5 42.8 88.9 64.6 63.4 63.2
+   Zero-shot (Radford et al., 2021) 76.8 37.3 36.7 63.6 84.0 46.7 86.7 66.1 63.7 62.4
+   LwF (Li & Hoiem, 2017) 66.6 26.9 19.5 51.0 78.4 26.6 68.9 35.5 56.1 47.7
+   WiSE-FT (Wortsman et al., 2022) 70.1 31.9 25.3 56.3 79.8 29.9 74.9 45.6 56.8 52.3
+   iCaRL (Rebuffi et al., 2017) 71.7 35.0 43.0 63.4 86.9 43.9 87.8 63.7 60.0 61.7
+   ZSCL (Zheng et al., 2023) 73.3 32.6 36.8 62.1 83.8 42.1 83.6 56.5 60.2 59.0
+   MoE-Adapter (Yu et al., 2024) 71.0 34.9 19.2 63.0 86.6 20.0 87.2 63.7 58.6 56.0
+   RAIL-Primal (Xu et al., 2024) 76.8 37.3 36.7 63.6 84.0 46.7 86.7 66.1 63.7 62.4
+   CoDyRA (Lu et al., 2024) 74.3 36.8 44.2 69.9 83.5 42.8 88.9 64.6 63.4 63.2
    MoRA 74.5 38.1 46.9 65.3 82.9 45.8 88.2 65.1 62.9 63.3
 Average
-   LwF [43] 24.7 79.7 38.3 36.9 63.9 81.0 36.5 71.9 42.7 56.7 53.2
-   WiSE-FT [78] 27.1 76.5 40.9 31.3 68.7 81.6 31.4 74.7 51.7 58.4 54.2
-   iCaRL [63] 25.4 72.1 37.5 51.6 65.1 87.1 59.1 88.0 63.7 60.1 61.0
-   ZSCL [92] 36.0 75.0 40.7 40.5 71.0 85.3 46.3 83.3 60.7 61.5 60.0
-   MoE-Adapter [85] 43.6 77.9 52.1 34.7 75.9 86.3 45.2 87.4 66.6 60.2 63.0
-   RAIL-Primal [81] 42.4 89.8 55.7 68.5 84.0 83.3 65.3 85.8 67.9 64.5 70.7
-   CoDyRA [48] 41.4 81.0 58.7 77.8 83.4 84.6 64.5 90.4 67.2 64.4 71.3
+   LwF (Li & Hoiem, 2017) 24.7 79.7 38.3 36.9 63.9 81.0 36.5 71.9 42.7 56.7 53.2
+   WiSE-FT (Wortsman et al., 2022) 27.1 76.5 40.9 31.3 68.7 81.6 31.4 74.7 51.7 58.4 54.2
+   iCaRL (Rebuffi et al., 2017) 25.4 72.1 37.5 51.6 65.1 87.1 59.1 88.0 63.7 60.1 61.0
+   ZSCL (Zheng et al., 2023) 36.0 75.0 40.7 40.5 71.0 85.3 46.3 83.3 60.7 61.5 60.0
+   MoE-Adapter (Yu et al., 2024) 43.6 77.9 52.1 34.7 75.9 86.3 45.2 87.4 66.6 60.2 63.0
+   RAIL-Primal (Xu et al., 2024) 42.4 89.8 55.7 68.5 84.0 83.3 65.3 85.8 67.9 64.5 70.7
+   CoDyRA (Lu et al., 2024) 41.4 81.0 58.7 77.8 83.4 84.6 64.5 90.4 67.2 64.4 71.3
    MoRA 44.1 81.6 64.6 79.6 83.9 84.4 66.5 89.7 68.4 64.1 72.7
 Last
-   LwF [43] 25.5 72.1 38.9 55.4 65.5 87.3 81.9 88.6 63.6 61.5 64.0
-   WiSE-FT [78] 21.8 76.8 42.9 20.8 77.5 84.9 30.7 76.6 75.8 72.5 58.0
-   iCaRL [63] 25.5 72.1 38.9 55.4 65.5 87.3 81.9 88.6 63.6 61.5 64.0
-   ZSCL [92] 33.1 75.3 43.5 35.2 74.6 87.4 50.4 84.2 77.3 73.4 63.4
-   MoE-Adapter [85] 43.2 78.7 57.6 32.8 79.4 86.0 86.7 87.8 78.2 74.2 70.5
-   RAIL-Primal [81] 41.7 94.0 66.0 86.4 97.2 82.4 93.1 83.6 75.0 71.3 79.1
-   CoDyRA [48] 37.7 81.5 65.1 89.9 91.4 85.5 96.8 93.3 77.3 73.5 79.2
+   LwF (Li & Hoiem, 2017) 25.5 72.1 38.9 55.4 65.5 87.3 81.9 88.6 63.6 61.5 64.0
+   WiSE-FT (Wortsman et al., 2022) 21.8 76.8 42.9 20.8 77.5 84.9 30.7 76.6 75.8 72.5 58.0
+   iCaRL (Rebuffi et al., 2017) 25.5 72.1 38.9 55.4 65.5 87.3 81.9 88.6 63.6 61.5 64.0
+   ZSCL (Zheng et al., 2023) 33.1 75.3 43.5 35.2 74.6 87.4 50.4 84.2 77.3 73.4 63.4
+   MoE-Adapter (Yu et al., 2024) 43.2 78.7 57.6 32.8 79.4 86.0 86.7 87.8 78.2 74.2 70.5
+   RAIL-Primal (Xu et al., 2024) 41.7 94.0 66.0 86.4 97.2 82.4 93.1 83.6 75.0 71.3 79.1
+   CoDyRA (Lu et al., 2024) 37.7 81.5 65.1 89.9 91.4 85.5 96.8 93.3 77.3 73.5 79.2
    MoRA 37.7 81.5 70.7 92.4 95.0 86.0 97.6 92.6 81.0 74.7 80.9
 4 Experiments
-In this paper, we conduct experiments across a diverse set of tasks, including continual learning for both vision-language CLIP models and LLMs, and analyze catastrophic forgetting during fine-tuning.
-Detailed implementation settings and more experiment results are provided in the Appendix.
-Table 2: Comparisons on 5-shot MTIL setting. Following the same protocol as in [85, 81, 48].
-    Method Aircraft [50] Caltech101 [25] CIFAR100 [39] DTD [15] EuroSAT [30] Flowers [56] Food [8] MNIST [19] OxfordPet [57] Cars [38] SUN397 [80] Average
-CLIP
-   Zero-shot [60] 24.3 88.4 68.2 44.6 54.9 71.0 88.5 59.4 89.0 64.7 65.2 65.3
-Transfer
-   Zero-shot [60] 88.4 68.2 44.6 54.9 71.0 88.5 59.6 89.0 64.7 65.2 69.4
-   LwF [43] 72.1 49.2 35.9 44.5 41.1 66.6 50.5 69.0 19.0 51.7 50.0
-   LwF-VR [21] 82.2 62.5 40.1 40.1 56.3 80.0 60.9 77.6 40.5 60.8 60.1
-   WiSE-FT [78] 77.6 60.0 41.3 39.4 53.0 76.6 58.1 75.5 37.3 58.2 57.7
-   ZSCL [92] 84.0 68.1 44.8 46.8 63.6 84.9 61.4 81.4 55.5 62.2 65.3
-   MoE-Adapter [85] 87.9 68.2 44.1 48.1 64.7 88.8 69.0 89.1 64.5 65.1 68.9
-   RAIL-Primal [81] 88.4 68.2 44.6 54.9 71.0 88.5 59.6 89.0 64.7 65.2 69.4
-   CoDyRA [48] 92.4 68.4 45.8 54.5 69.6 87.4 65.2 88.5 64.2 64.5 69.9
-   MoRA 92.0 68.8 45.6 53.1 68.6 84.4 64.3 89.8 65.4 64.8 69.7
-Average
-   LwF [43] 23.5 77.4 43.5 41.7 43.5 52.2 54.6 63.4 68.0 21.3 52.6 49.2
-   LwF-VR [21] 24.9 89.1 64.2 53.4 54.3 70.8 79.2 66.5 79.2 44.1 61.6 62.5
-   WiSE-FT [78] 32.0 87.7 61.0 55.8 68.1 69.3 76.8 71.5 77.6 42.0 59.3 63.7
-   ZSCL [92] 28.2 88.6 66.5 53.5 56.3 73.4 83.1 56.4 82.4 57.5 62.9 64.4
-   MoE-Adapter [85] 30.0 89.6 73.9 58.7 69.3 79.3 88.1 76.5 89.1 65.3 65.8 71.4
-   RAIL-Primal [81] 32.9 94.5 69.9 58.1 71.8 84.4 88.5 70.4 89.0 66.1 65.7 71.9
-   CoDyRA [48] 34.6 95.8 73.9 60.0 77.1 81.3 86.6 75.9 89.9 66.1 65.3 73.3
-   MoRA 36.7 95.4 74.9 61.9 77.1 82.6 85.3 76.0 90.5 67.0 65.6 73.9
-Last
-   LwF [43] 22.1 58.2 17.9 32.1 28.1 66.7 46.0 84.3 64.1 31.5 60.1 46.5
-   LwF-VR [21] 22.9 89.8 59.3 57.1 57.6 79.2 78.3 77.7 83.6 60.1 69.8 66.9
-   WiSE-FT [78] 30.8 88.9 59.6 60.3 80.9 81.7 77.1 94.9 83.2 62.8 70.0 71.9
-   ZSCL [92] 26.8 88.5 63.7 55.7 60.2 82.1 82.6 58.6 85.9 66.7 70.4 67.4
-   MoE-Adapter [85] 30.1 89.3 74.9 64.0 82.3 89.4 87.1 89.0 89.1 69.5 72.5 76.1
-   RAIL-Primal [81] 32.9 95.1 70.3 63.2 81.5 95.6 88.5 89.7 89.0 72.5 71.0 77.2
-   CoDyRA [48] 31.6 95.5 72.8 63.5 85.0 89.7 85.0 94.7 93.2 73.6 73.0 78.0
-   MoRA 32.5 95.3 75.3 66.6 87.8 92.6 86.3 96.3 92.6 73.5 73.8 79.3
-CL of CLIP We evaluate on two benchmarks using CLIP-ViT/B-16 [60, 33]: MTIL [92, 85], with 11 datasets learned sequentially under a 5-shot split, and X-TAIL [81, 48], with a 16-shot split.
-Following prior work, we report Transfer (zero-shot on unseen tasks), Last (retention on earlier tasks), and Average (mean of Transfer and Last) to assess continual learning performance.
-CL of LLM.
-We follow prior work [59, 62, 74, 58] to continually fine-tune T5-large [61] on five text-classification benchmarks—AG News, Amazon Reviews, Yelp Reviews, DBpedia, and Yahoo Answers—under three different task orderings.
-We report the average final accuracy across all tasks after completing the last task, measuring the model’s ability to retain knowledge during CL.
+In this paper, we conduct experiments across a diverse set of tasks, including continual learning for both vision-language CLIP models and LMs, and analyze catastrophic forgetting during fine‐tuning.
+Detailed implementation settings and more experiment results are provided in Appendix A.1.
+CL of CLIP.
+We evaluate on two benchmarks using CLIP-ViT/B-16 (Radford et al., 2021; Ilharco et al., 2021) on MTIL (Zheng et al., 2023; Yu et al., 2024) and X-TAIL (Xu et al., 2024; Lu et al., 2024), with a range of image classification datasets.
+Following prior work, we report Transfer (zero-shot on unseen tasks), Last (retention on earlier tasks), and Average (mean accuracy across all datasets of all learning tasks) to assess continual learning performance.
+CL of LM.
+We follow prior works (Qin & Joty, 2021; Razdaibiedina et al., 2023; Wang et al., 2023; Qiao & Mahdavi, 2024) to continually fine-tune T5-large (Raffel et al., 2020) and LLaMA2-7B (Touvron et al., 2023) on five text-classification benchmarks, under three different task orderings.
+We report the average final accuracy across all tasks after completing the last task.
 Generalization and forgetting on unseen tasks.
-To assess effects on pre-trained general knowledge, we fine-tune Llama3 [28] on the CodeAlpaca code-generation dataset [9] and evaluate zero-shot in-domain performance on HumanEval [12], as well as out-of-domain accuracy on a broad selection of MMLU [31] subjects—Formal Logic, Philosophy, World Religions, Economics, Public Relations, STEM, Physics, and Machine Learning.
-These results reveal how fine-tuning impacts both task-specific adaptation and catastrophic forgetting or improved generalization to unseen domains.
+To assess effects on pre-trained general knowledge, we fine-tune Llama3.1-8B (Grattafiori et al., 2024) on the CodeAlpaca code-generation dataset (Chaudhary, 2023) and evaluate zero-shot in-domain performance on HumanEval (Chen et al., 2021), as well as out-of-domain accuracy on a broad selection of MMLU (Hendrycks et al., 2021) subjects.
+Table 2: Zero-shot in-domain performance on HumanEval (Pass@1) for the code generation domain and out-of-domain accuracy on selected MMLU subjects (formal logic, philosophy, world religions, economics, public relations, STEM, physics, machine learning) after fine-tuning Llama-3.1-8B on CodeAlpaca.
+The last two columns report trainable parameters (for MoRA: added / activated).
+Method
+HumanEval
+(Pass@1)
+Out-of-Domain (Acc.)
+Params
+[…]
+Logic Phil. Reli. Econ. Pub. Rel. STEM Phys. ML MMLU
+Llama-3.1-8B 38.40 42.06 71.06 83.63 70.17 68.18 54.84 39.22 40.18 63.45
+LoRA ($r=4$) 41.46 39.68 70.09 81.87 71.43 65.45 54.77 45.10 40.17 63.28 10.5 0.13%
+LoRA ($r=8$) 44.51 39.68 70.74 81.87 71.85 64.54 54.17 42.16 39.29 63.03 21.0 0.26%
+LoRA ($r=16$) 45.73 41.27 68.49 80.70 72.69 66.36 54.96 44.11 38.39 63.35 41.9 0.52%
+LoRA ($r=32$) 47.56 42.85 69.45 81.87 72.27 66.36 55.44 45.10 39.29 63.59 83.9 1.03%
+MoRA 47.56 48.41 70.09 82.46 73.53 68.18 55.53 46.08 41.96 63.70 41.9/26.2 0.52%/0.32%
 4.1 Continual Learning of CLIP
-Cross-domain task-agnostic incremental learning.
-Table 1 presents our results on the X-TAIL benchmark.
-Prior methods [85, 81] rely on zero-shot predictions to preserve pre-trained capabilities and mitigate forgetting, and as a result are capped by the base model’s zero-shot performance.
-Similarly as done in previous works [92, 48], MoRA aims to continuously refines the pre-trained model, enabling it to surpass this zero-shot ceiling and improve the generalization of the pre-trained model on unseen domains.
-In practice, continual parameter updates can slightly perturb some zero-shot predictions, our method ultimately improves overall performance on unseen domains by adapting the pre-trained model to new tasks while enhancing its generalization capabilities.
-Table 3: Summary of the results on standard CL benchmarks with T5-large model. The averaged accuracy after training on the last task is reported. Experiments on long task sequences in Table 9.
+Table 1 reports X-TAIL results (MTIL results in Appendix Table 11).
+Prior methods (Yu et al., 2024; Xu et al., 2024) preserve pre-trained capabilities via zero-shot predictions and are therefore capped by the base model’s zero-shot performance.
+Following recent practice (Zheng et al., 2023; Lu et al., 2024), MoRA continuously trains the pre-trained model so it can surpass that zero-shot ceiling and better generalize to unseen domains, while still preserving existing capabilities.
+Across both Last and Average metrics, MoRA substantially outperforms prior works.
+Unlike methods that rely on domain-ID prediction (Yu et al., 2024), high-dimensional projection layers or feature-banks (Xu et al., 2024), MoRA improves the pre-trained representation without external mechanisms; compared with (Lu et al., 2024), which learns fixed weights per rank-$r$ update, MoRA uses input-dependent gates on individual rank-one updates for finer expert specialization and dynamic utilization.
+Table 3: Summary of results on standard CL benchmarks with T5-large. We report averaged accuracy after training on the last task across three task orderings.
 Standard CL Benchmark
@@
 LFPT5 67.6 72.6 77.9 72.7
+InfLoRA 75.2 75.4 75.8 75.5
 O-LoRA 75.4 75.7 76.3 75.8
 LB-CL 76.9 76.5 76.8 76.7
-MoRA 77.43 77.47 77.93 77.61
-On the Last and Average metrics, MoRA significantly outperforms prior methods.
-Unlike approaches that require domain ID prediction [85], high dimensional projection layers, or a memory bank of feature embeddings [81], MoRA enhances the pre-trained model’s own representation capability without relying on these external mechanisms.
-In contrast to the previous state of the art [48], which learns a fixed weight for each rank-$r$ update from each task, MoRA’s novel mixture-of-ranks framework selectively assigns each rank an input-dependent gate.
-Notably, only a small subset of ranks most relevant to each token are activated, resulting in far more fine-grained utilization and specialization of each ranks.
-Multi-domain task incremental learning.
-We evaluate MoRA in the few-shot MTIL setting (Table 2) under the same protocols as [85, 81, 48].
-Consistent with the results observed in the X-TAIL setting, our method demonstrates clear superiority in this scenario.
-4.2 Continual Learning of LLM
-Table 3 reports results across three task orderings, where MoRA consistently outperforms prior methods and nearly matches the multi-task learning (MTL) upper bound.
-In contrast to state-of-the-art techniques such as O-LoRA [74] and LB-CL [58], which rely on orthogonality constraints or gradient projections between per-task LoRA adapters—potentially limiting each adapter’s capacity—MoRA requires no additional regularization.
-By decomposing each rank-$r$ update into its rank-1 components and applying self-activated, sparse gating, MoRA allows every component to specialize on its own input distribution.
-This fine-grained specialization both prevents interference across tasks and more effectively captures diverse patterns.
-4.3 Forgetting and Generalization during Fine-Tuning
+MoRA 77.4 77.5 77.9 77.6
+4.2 Continual Learning of LM
+Table 3 reports results across three task orderings: MoRA consistently outperforms prior methods and closely approaches the multi-task learning (MTL) upper bound.
+Unlike O-LoRA (Wang et al., 2023) and LB-CL (Qiao & Mahdavi, 2024), which rely on orthogonality constraints or gradient projections between per-task LoRA adapters (potentially limiting adapter capacity), MoRA needs no extra regularization.
+By decomposing each rank-r update into rank-one components and applying self-activated, sparse gating, MoRA lets each component specialize on its own input distribution, reducing interference and more effectively capturing diverse patterns.
+We further evaluate MoRA on LLaMA2-7B (Touvron et al., 2023) under the same continual-learning setup (Table 9 in the Appendix), MoRA also outperforms O-LoRA by 2.3% averaged over 3 task orders.
+4.3 Forgetting and Generalization during Fine‐Tuning
 Table 2 examines how standard fine-tuning affects both in-domain performance and out-of-domain generalization.
-We observe two key phenomena: 1.
-Forgetting of unrelated knowledge.
-Fine-tuning with a conventional LoRA leads to performance drops on topics completely unrelated to the fine-tuning data—for example, Philosophy, World Religions, and Public Relations—indicating that the model loses pre-trained general knowledge when adapting to a new domain.
-2.
-Improved generalization.
-Fine-tuning on code generation data also boosts accuracy on tasks requiring logical understanding (Formal Logic, Economics, STEM, Physics, and Machine Learning), suggesting that representations learned during code fine-tuning transfer to related symbolic and quantitative domains.
-By contrast, MoRA’s self-routed, sparse mixture-of-ranks framework both preserves unrelated general knowledge and enhances in-domain and out-of-domain generalization performance.
-By decomposing each low-rank update into specialized rank-1 components and activating only the most relevant subset, MoRA prevents catastrophic overwriting of prior knowledge.
-At the same time, it achieves strong code-generation performance using only one-third of the activated parameters required by a rank-32 LoRA, and significantly improves the out-of-domain generalization.
-Table 4: Fine-tuning Llama3.1-8B on the code generation domain (CodeAlpaca).
-We report zero-shot in-domain performance on HumanEval (Pass@1) and zero-shot out-of-domain accuracy for selected MMLU subjects—Formal Logic, Philosophy, World Religions, Economics, Public Relations, STEM, Physics, and Machine Learning.
-The last two columns list trainable parameters (for MoRA: added / activated).
-The best and second best scores are highlighted in red and blue, respectively.
-[…]
-Logic Phil. Reli. Econ. Pub. Rel. STEM Phys. ML MMLU
-Llama-3.1-8B 38.40 42.06 71.06 83.63 70.17 68.18 54.84 39.22 40.18 63.45
-LoRA ($r=4$) 41.46 39.68 70.09 81.87 71.43 65.45 54.77 45.10 40.17 63.28 10.5 0.13%
-LoRA ($r=8$) 44.51 39.68 70.74 81.87 71.85 64.54 54.17 42.16 39.29 63.03 21.0 0.26%
-LoRA ($r=16$) 45.73 41.27 68.49 80.70 72.69 66.36 54.96 44.11 38.39 63.35 41.9 0.52%
-LoRA ($r=32$) 47.56 42.85 69.45 81.87 72.27 66.36 55.44 45.10 39.29 63.59 83.9 1.03%
-MoRA 47.56 48.41 70.09 82.46 73.53 68.18 55.53 46.08 41.96 63.70 41.9/26.2 0.52%/0.32%
+We observe two key phenomena.
+(1) Forgetting of unrelated knowledge: conventional LoRA fine-tuning causes drops on topics that are unrelated to the fine-tuning data (e.g., Philosophy, World Religions, Public Relations), indicating that pre-trained general knowledge is overwritten during adaptation.
+(2) Improved generalization: fine-tuning on code data increases accuracy on logically and quantitatively oriented tasks (Formal Logic, Economics, STEM, Physics, Machine Learning), showing that representations learned from code transfer to related symbolic domains.
+In contrast, MoRA both preserves unrelated pre-trained knowledge and improves in- and out-of-domain generalization.
+Decomposing each low-rank update into specialized rank-one components and activating only the most relevant subset prevents catastrophic overwriting of prior knowledge while allowing targeted adaptation.
+At the same time, MoRA achieves strong code-generation performance using roughly one-third of the activated parameters required by a rank-32 LoRA and substantially improves out-of-domain accuracy.
+4.4 Visualizations of Fine-grained Rank Activations
+Figure 3 shows rank activations recorded during continual learning.
+These visualizations illustrate two properties of MoRA: (1) individual ranks specialize on distinct input patterns, and (2) the model substantially reduces cross-task interference, thereby mitigating forgetting.
+Extended visualizations across more tasks and scenarios appear in Fig.
+4 and Fig.
+5 in the Appendix.
+Each rank specializes in distinct input patterns.
+In Fig.
+3(a), airplane patches strongly activate Rank 0, while blue-sky backgrounds predominantly activate Rank 11.
+In Fig.
+3(c), Ranks 19, 20, and 29 (orange boxes) respond to jaguar patches.
+The more complex backgrounds in Fig.
+3(c) (e.g., leaves, shadows) yield a richer, more distributed pattern than the simple blue sky in Fig.
+3(a), highlighting MoRA’s capacity to model contextual complexity.
+Some ranks learned earlier are also reused on later tasks (Fig.
+5, Appendix), indicating transfer of shared semantics.
+Reduced cross-task interference and mitigated forgetting.
+Comparing the same input after Task 1 (Fig.
+3(a)) and after Task 2 (Fig.
+3(b)) shows almost identical activations (more in Fig.
+4, Appendix): Rank 0 still responds to airplane semantics and Rank 11 to blue-sky patches.
+This stability indicates that our self-activated, sparse mixture-of-ranks prevents later updates from overwriting earlier task representations, reducing inter-task interference and mitigating catastrophic forgetting.
 […]
@@
 More visualizations are in Figs.
-5 and 6 of Appendix.
-4.4 Visualizations on Fine-grained Rank Activations
-We visualize rank activations during continual learning in Fig.
-3.
-These visualizations demonstrate that MoRA (1) encourages each rank to specialize in its own input distribution, and (2) reduces cross-task interference to mitigate forgetting.
-Each rank specializes in distinct input patterns.
-In Fig.
-6(a), we observed that patches corresponding to the airplane object strongly activate rank 0, whereas background patches (blue sky) predominantly activate rank 11.
-Similarly as in Fig.
-6(c), ranks at index 19, 20, and 29 (highlighted with orange bounding boxes) are activated for image patches of jaguar.
-Moreover, the complex background elements in Fig.
-6(c)—such as tree leaves and shadows—elicit a more diverse activation pattern than the simple blue-sky backgrounds in Fig.
-6(a), illustrating MoRA’s capacity to capture varying contextual complexity.
-Interestingly, we also observe that certain ranks learned on earlier tasks remain active when processing later-task inputs, indicating effective knowledge reuse.
-Reduced cross-task interference to mitigate forgetting.
-By comparing the activation patterns in Fig.
-6(a) (after learning Task 1) and Fig.
-6(b) (after learning Task 2) on the same input image, we observe that the ranks activations are nearly identical: rank 0 continues to respond to airplane semantics, and rank 11 continues to respond to background blue-sky patches.
-This consistency demonstrates that our self-activated sparse mixture-of-ranks framework effectively prevents later-task updates from overwriting or intruding upon the knowledge learned for earlier tasks.
-As a result, inter-task interference is greatly reduced, and catastrophic forgetting is mitigated.
-Table 5: Comparisons on different per LoRA/rank mixture calculation strategy
+4 and 5 of the Appendix, demonstrating forgetting mitigation and knowledge reuse.
+4.5 Ablation Studies
+Routing/activation strategy.
+Table 4 compares different strategies for computing mixture weights, under a controlled setup with one LoRA per task, standard training loss, and no additional regularization: (1) MoE-LoRA trains a router to assign weights to each LoRA adapter; its coarse-grained design and lack of mechanisms that encourage expert specialization lead to greater interference and increased forgetting compared to MoRA.
+(2) Non-self-activation trains a finer-grained router for mixing each rank-one component.
+However, with a much larger number of experts (ranks), the router becomes increasingly prone to confusion, resulting in higher training complexity and degraded performance compared with MoE-LoRA.
+(3) Self-activation infers weights directly from softmax-normalized activations (Eq.
+8), matching or exceeding Non-self-activation and demonstrating that each rank can infer its own relevance without extra trainable router modules.
+(4) Top-k rank selection restricts updates to the k most relevant ranks, reducing interference and improving accuracy.
+(5) Sharpness control concentrates activation mass and substantially boosts specialization.
+(6) Test-time rank thresholding removes low-importance ranks at inference, yielding a slight gain by eliminating noisy contributions.
+Table 4: Routing/activation strategy
 Mixture Calculation Transfer Average Last
-Router@LoRA 62.56 69.45 74.53
-Router@rank 60.09 65.97 69.76
-Self-activated@rank 60.26 65.94 69.85
-     + Sparse activation 62.59 72.28 80.22
-     + Adaptive sparse (MoRA) 63.30 72.70 80.90
+MoE-LoRA 62.56 69.45 74.53
+Non-self-activation 60.09 65.97 69.76
+Self-activation 60.26 65.94 69.85
+    w/ top-k rank selection 60.69 66.52 70.62
+    w/ sharpness control with $\tau_{\text{MoRA}}$ 62.07 71.15 79.62
+    w/ test-time rank thresholding 60.78 66.83 71.08
+    MoRA 63.30 72.70 80.90
+[…]
 (a) Acti. Budget
@@
 Figure 3: Ablation on (a) rank activation budget, (b) temperature $\tau_{\text{MoRA}}$, and (c) threshold $\delta$.
-4.5 Ablation Studies
-Per-LoRA/rank mixture weights.
-Table 5 compares different strategies for computing mixture weights, under a controlled setup with one LoRA per task, standard training loss only, and no additional regularization: (1) Router@LoRA trains a standard router to assign weights to each LoRA adapter.
-(2) Router@rank trains a fine-grained router for every rank-1 component, which increases training complexity and underperforms Router@LoRA.
-(3) Self-activated@rank directly infer mixture weights using softmax-normalized activations (Eq.
-5), matching or exceeding Router@rank and showing that each rank can infer its own relevance without extra parameters.
-(4) Sparse activation adds temperature scaling (Eq.
-6) and caps the number of activated ranks, yielding substantial gains by sharpening rank specialization.
-(5) Adaptive sparse activation further prunes low-activation ranks at test time, selecting the most relevant number of ranks to be activated adaptively for each input token, reducing noise from less relevant ranks and achieving the best overall performance.
 Rank activation budget.
-Under our sparse rank-activation framework, we explore how many ranks are necessary by varying the activation budget (Fig.
+We probe how many ranks are necessary by varying the activation budget (Fig.
 4(a)).
-To focus exclusively on this budget parameter, we disable other sparsity enhancements, e.g.
-rank pruning.
-After training on 10 tasks (160 ranks total), raising the budget from 2 (1.25%) to 16 active ranks (10%) yields a dramatic boost in final accuracy, and gradually plateaus.
-This demonstrates that each rank specializes in a distinct input distribution, and a small, sparse subset of ranks suffices to handle diverse input distributions.
-Moreover, even when the budget is set very high, our self-activating gating remains robust: it still activates only the most relevant ranks for each token while suppressing contributions from less relevant ones.
-Temperature setting of $\tau_{\text{MoRA}}$.
+To isolate this effect, we disable other sparsity mechanisms (e.g., thresholding).
+After training on 10 tasks (160 ranks total), increasing the budget from 2 (1.25%) to 16 active ranks (10%) yields a large improvement in final accuracy and then gradually plateaus.
+This suggests that each rank specializes in distinct input patterns and that a small, sparse subset of ranks is sufficient to cover diverse inputs.
+Even with a high budget, the self-activating gating remains robust: it still privileges the most relevant ranks per token while suppressing lower-relevance contributions.
+Sharpness control of $\tau_{\text{MoRA}}$.
 We use temperature scaling (Eq.
 (10)) to control mixture sharpness: lower $\tau_{\text{MoRA}}$ yields sharper activations, while higher $\tau_{\text{MoRA}}$ flattens the distribution and engages more ranks.
-Figure4(b) shows that increasing $\tau_{\text{MoRA}}$ improves transfer to unseen tasks, even surpassing the reported state-of-the-art Transfer in Table 1 when $\tau_{\text{MoRA}}$ is high.
+Fig.
+4(b) shows that increasing $\tau_{\text{MoRA}}$ improves transfer to unseen tasks, even surpassing the reported state-of-the-art Transfer in Table 1 when $\tau_{\text{MoRA}}$ is high.
 A moderate $\tau_{\text{MoRA}}$ around 0.01 balances specificity and generalization, delivering strong Last and Average scores.
-Adaptive mixture with rank pruning.
-In Fig.
-4(c), we demonstrate the effects of adaptive sparsity constraints by varying the threshold $\delta$ for rank pruning.
-Pruning ranks with a relatively small threshold reduces noise and sharpens the mixture, improving both downstream adaptation and out-of-domain generalization, whereas raising $\delta$ above 0.2 begins to remove useful ranks and degrades performance.
+Threshold-based rank selection.
+Figure 4(c) shows the effect of the test-time rank selection threshold $\delta$.
+Applying a modest threshold removes low-activation ranks and reduces noisy contributions, which improves both downstream adaptation and out-of-domain generalization.
 5 Conclusion
-We present MoRA, a self-activated sparse mixture-of-ranks framework that decomposes each LoRA update into rank-1 experts and self-infer their sparse mixture weights.
-By reducing knowledge redundancy and resolving router–expert mismatch during CL, MoRA delivers superior continual-learning performance, significantly reduces forgetting, and enhances generalization.
-Limitations and future works.
-MoRA use a pre-defined temperature to control the sharpness of the mixture weights.
-While effective, the sparse activation mechanism can be further improved by incorporating awareness of the global data distribution.
+We present MoRA, a self-activated sparse mixture-of-ranks framework that decomposes each LoRA update into rank‐1 experts and self‐infer their sparse mixture weights.
+By reducing knowledge redundancy and resolving router–expert mismatch during CL, MoRA delivers superior continual‐learning performance, significantly reduces forgetting, and enhances generalization.
+Future works.
+MoRA controls mixture sharpness and sparsity via top-k selection, temperature scaling, and thresholding.
+While effective, the sparse activation mechanism can be further improved by incorporating awareness of the input data distribution.
 Learning this temperature or adapting the mixture’s sharpness based on data offers a promising avenue for future research.
 We also aim to extend the method’s applicability to a broader range of PTMs and applications.
-Acknowledgments and Disclosure of Funding
-This work was partially supported by the ARC DECRA Fellowship (DE230101591) awarded to D.
-Gong.
-H.
-Lu is affiliated with CSIRO Data61 through a PhD scholarship and acknowledges the support of the Google PhD Fellowship.
-Appendix A Additional Details of MoRA
-We provide more details of developing the proposed MoRA in this section.
-A.1 Model Weights and Updates as Linear Associative Memory Model
-Weight matrix as linear associative memory.
-Large pre-trained models (PTMs) have been shown to inhabit a low “intrinsic-dimension” manifold compared to their full parameter count [41, 1].
-Any weight matrix of a pre-trained model $\mathbf{W}_{0}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$ can be treated as a linear associative memory model with a key-value structure [37, 5, 6, 54].
-Concretely, for any matrix $\mathbf{W}\in\mathbb{R}^{d_{\mathrm{out}}\times d_{\mathrm{in}}}$ and input $\mathbf{x}\in\mathbb{R}^{d_{\mathrm{in}}}$, the product $\mathbf{W}\,\mathbf{x}$ implements a content-addressable memory read over $m$ stored key–value pairs $\{(\mathbf{K}_{i},\mathbf{V}_{i})\}_{i=1}^{m}$, where $m$ denotes the number of key-value memories of a specific weight matrix.
-Here, ${\mathbf{K}}=\begin{bmatrix}{\mathbf{K}}_{1}^{\top}&\cdots&{\mathbf{K}}_{m}^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{m\times d_{\text{in}}}$ and ${\mathbf{V}}=$$\begin{bmatrix}{\mathbf{V}}_{1}&\cdots&{\mathbf{V}}_{m}\end{bmatrix}\in\mathbb{R}^{d_{\text{out}}\times m}$ with $\mathbf{K}_{i}\in\mathbb{R}^{1\times d_{\text{in}}}$ is the $i$-th row of ${\mathbf{K}}$ and ${\mathbf{V}}_{i}\in\mathbb{R}^{d_{\text{out}}\times 1}$ is the $i$-th column of ${\mathbf{V}}$.
-Given any input $\mathbf{x}$, the output $\mathbf{y}\in\mathbb{R}^{d_{\text{out}}}$ is computed as
-$\mathbf{y}=\mathbf{W}\mathbf{x}\approx\sum\nolimits_{i=1}^{m}\mathbf{V}_{i}(\mathbf{K}_{i}\mathbf{x}).$ (8)
-In this formulation, the forward pass through the weight matrix with the input decomposes into two steps: (1) computing response scores or assignment weights via the key vectors $\mathbf{K}_{i}$, and (2) aggregating the corresponding value vectors $\mathbf{V}_{i}$ according to those weights.
-Low-rank update as modifications to the linear associative memory.
-Building on the associative-memory view of weights, we can likewise interpret weight updates during training as key–value insertions into the existing memory.
-During training, for an arbitrary update $\Delta\mathbf{W}$ on a given $\mathbf{W}_{0}$ as the starting point, the weight update process can be represent as
-$\mathbf{W}_{\text{new}}=\mathbf{W}_{0}+\Delta\mathbf{W},$ (9)
-where $\mathbf{W}_{0}$ can be the pre-trained model weights.
-We can also formulate the weight updates $\Delta\mathbf{W}$ with the key-value memory model $\{\Delta\mathbf{K},\Delta\mathbf{V}\}$ as an update to the original memory model.
-With the operation on an input $\mathbf{x}$ via
-$\Delta\mathbf{W}\mathbf{x}=\Delta\mathbf{V}\Delta\mathbf{K}\mathbf{x},$ (10)
-the updates on weights can be seen as a retrievable key-value system for the updating of $\mathbf{W}\mathbf{x}$.
-Low-Rank Adaptation (LoRA) [32] performs parameter-efficient fine-tuning by adding a rank-$r$ update $\Delta\mathbf{W}=\mathbf{B}\mathbf{A}$ to the pre-trained weight matrix $\mathbf{W}_{0}$, where $\mathbf{B}\in\mathbb{R}^{d_{\text{out}}\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times d_{in}}$.
-With the model update based on LoRA, given an input $\mathbf{x}$, the computation of updated weight becomes
-$\mathbf{y}_{\text{new}}=\mathbf{W}_{\text{new}}\mathbf{x}=\mathbf{W}_{0}\mathbf{x}+\Delta\mathbf{W}\mathbf{x}=\mathbf{W}_{0}\mathbf{x}+\mathbf{B}\mathbf{A}\mathbf{x}.$ (11)
-Decomposing the low-rank matrices into a collection of its $r$ rank-1 components gives
-$\Delta\mathbf{W}\mathbf{x}=\mathbf{B}\mathbf{A}\mathbf{x}=\sum\nolimits_{i=1}^{r}\mathbf{B}_{i}(\mathbf{A}_{i}\mathbf{x}),$ (12)
-with $\mathbf{A}=\begin{bmatrix}\mathbf{A}_{1}^{\top}&\cdots&\mathbf{A}_{r}^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{r\times d_{\text{in}}}$ where $\mathbf{A}_{i}$ is the $i$-th row of $\mathbf{A}$, and $\mathbf{B}=\begin{bmatrix}\mathbf{B}_{1}&\cdots&\mathbf{B}_{r}\end{bmatrix}\in\mathbb{R}^{d_{\text{out}}\times r}$ where $\mathbf{B}_{i}$ is the $i$-th column of $\mathbf{B}$.
-Considering the connection between the formulation of LoRA Eq.
-(1) and associative-memory view of weight update Eq.
-(10), we can naturally interpret LoRA as maintaining a small key–value store in a low-rank subspace, where the rows of $\mathbf{A}$ serve as keys and the columns of $\mathbf{B}$ serve as their corresponding values.
-After training the LoRA module on a base weight matrix $\mathbf{W}_{0}$, the modified forward pass in Eq.
-(11) with Eq.
-(12) can be seen as: first, the original pre-trained computation $\mathbf{W}_{0}\,\mathbf{x}$ (Eq.
-(8)) is performed; then, each learned key $\mathbf{A}_{i}$ measures its relevance to the input $\mathbf{x}$ via the product $\mathbf{A}_{i}\mathbf{x}$, and each learned value $\mathbf{B}_{i}$ is scaled by this relevance score and aggregated, in a same way as shown in Eq.
-(8).
-In this way, LoRA effectively inserts or updates a small set of key–value memories that augment the pre-trained model’s capabilities.
-A.2 Sparse Mixture-of-Rank Adaptation
-In the key–value associative memory view of LoRA (Eq.
-(11) and Eq.
-(12)), a standard LoRA or its variants [53, 87] apply a dense, fixed update: every rank’s subspace contributes for any kinds of input $\mathbf{x}$.
-Even in MoE-LoRA, where each LoRA module is treated as a separate expert, all key-value pairs within a LoRA module are used densely as a whole, either fully activated or not activated at all.
-This leads to redundancy across experts and overlooked the strong potential of the key vectors $\mathbf{A}$ on expressing its relevance to the input.
-Rather than activating all key-value memories of each LoRA for all kinds of inputs, where many of which may encode patterns irrelevant to the current input, we introduce a gated, sparse mixture-of-ranks update that selectively activates only the most relevant key–value pairs for each input.
-Concretely, we associate each rank-1 key–value pair with a scalar weight $w_{i}\in\mathbb{R}$, so that for an input token $\mathbf{x}\in\mathbb{R}^{d_{\mathrm{in}}}$ from task $t$, with $r_{t}$ total ranks, the update becomes:
-$\Delta\mathbf{W}^{t}=\sum\nolimits_{i=1}^{r_{t}}w_{i}\mathbf{B}_{i}\mathbf{A}_{i},~{}~{}\text{and}~{}~{}\Delta\mathbf{W}^{t}\mathbf{x}=\sum\nolimits_{i=1}^{r_{t}}w_{i}\mathbf{B}_{i}\mathbf{A}_{i}\mathbf{x},$ (4)
-where $w_{i}$ modulates whether and how strongly the corresponding rank contributes to the final updated weight matrix.
-As discussed in Sec.
-3.4, rather than obtaining mixture weights $w_{i}$’s through learning an additional router [85, 72], considering that $\mathbf{A}_{i}$’s in $\mathbf{A}$ works as keys in the memory model that reflect the significance on the correlation between input and the learned knowledge in each rank, we leverage the key vectors $\mathbf{A}_{i}$themselves to infer each rank’s relevance.
-Specifically, $w_{i}$’s are obtained through sparsity-inducing scaling with top-k masking and rank pruning of each rank’s activation $\mathbf{A}_{i}\mathbf{x}$, i.e., $[w_{i}]=\texttt{SparseOperator}([\mathbf{A}_{i}\mathbf{x}])$, where $\texttt{SparseOperator}()$ denotes the sparse-inducing operations.
-Our fine-grained mixture framework offers two key advantages: (1) It reduces redundancy and enables shared-knowledge reuse: ranks that capture common, cross-task features (such as generic semantics) can be reactivated across multiple tasks, preventing those patterns from being relearned redundantly or interfered; (2) It minimizes routing ambiguity and prevents forgetting: each rank can specialize on more nuanced patterns, yielding an optimized, rank-wise mixture that integrates new knowledge without disrupting existing representations or hindering future updates.
-Appendix B Additional Details and Results of Experiments
-Table 6: Details of the 15 datasets used in our continual-learning experiments using LLM.
-NLI denotes natural language inference, and QA denotes question-answering tasks.
+Reproducibility Statement
+Further implementation details and dataset descriptions are provided in Sec. 4 and Appendix A.1. Our source code will be publicly released upon paper acceptance.
+Appendix A Additional Details and Results of Experiments
+Table 5: Details of the 15 datasets used in our continual‐learning experiments using LMs.
+NLI denotes natural language inference, and QA denotes question‐answering tasks.
 The first five tasks comprise the standard CL benchmark; the remaining ten tasks are used for the extended long-sequence evaluations.
@@
 in both sentence? Choose one from the option.
-Table 8: The six task-sequence orders used in our continual learning experiments.
+Table 7: The six task‐sequence orders used in our continual learning experiments.
 Sequences 1–3 follow the standard CL benchmarks employed in prior work.
-Sequences 4–6 extend to longer 15-task streams, as introduced in [62].
+Sequences 4–6 extend to longer 15‐task streams, as introduced in (Razdaibiedina et al., 2023).
 Order Task Sequence
@@
 sst-2 → dbpedia → ag → yahoo → multirc → boolqa → wic
-B.1 Detailed Experiment Settings
+A.1 Detailed Experiment Settings
 CL on CLIP.
-We follow the experimental setups in [92, 85, 81] and use the CLIP model with a ViT-B/16 backbone [60] for all experiments.
+We follow the experimental setups in (Zheng et al., 2023; Yu et al., 2024; Xu et al., 2024) and use the CLIP model with a ViT-B/16 backbone (Radford et al., 2021) for all experiments.
 By default, MoRA is applied to every pre-trained weight matrix in both the vision and text encoders, with an initial rank of 16 per update.
-Each task is trained for 500 iterations using AdamW [47] with learning rate of $45e-$ 4.
+Each task is trained for 500 iterations using AdamW (Loshchilov & Hutter, 2017) with a learning rate of 5e45e-4.
 During continual learning, we freeze the ranks learned from previous tasks and initialize new $r=16$ ranks for each incoming task.
@@
 We set the temperature $\tau_{\text{MoRA}}=0.1$ and the threshold $\delta=0.2$.
-CL on LLMs.
-We follow the protocol of previous work in continually fine-tuning a T5-large model on a suite of text-classification tasks.
-We train on five standard benchmarks—AG News, Amazon Reviews, Yelp Reviews, DBpedia, and Yahoo Answers—using three distinct task orderings drawn from [59, 62, 74, 58].
-To probe longer sequences, we extend this to a 15-dataset stream (Table 6), incorporating tasks from the original CL benchmark [90], GLUE [68], SuperGLUE [69], and the IMDB movie reviews corpus.
+CL on LMs.
+We follow the protocol of previous work in continually fine-tuning the T5-large (Raffel et al., 2020) and LLaMA2-7B (Touvron et al., 2023) on a suite of text-classification tasks.
+We train on five standard benchmarks—AG News, Amazon Reviews, Yelp Reviews, DBpedia, and Yahoo Answers—using three distinct task orderings drawn from (Qin & Joty, 2021; Razdaibiedina et al., 2023; Wang et al., 2023; Qiao & Mahdavi, 2024).
+To probe longer sequences, we extend this to a 15-dataset stream (Table 5), incorporating tasks from the original CL benchmark (Zhang et al., 2015), GLUE (Wang et al., 2018), SuperGLUE (Wang et al., 2019), and the IMDB movie reviews corpus.
 Natural language prompts for each task are presented in Table 6, with NLI tasks (MNLI, RTE, CB), sentiment classification (Amazon, Yelp, SST-2, IMDB), and topic classification (AG News, DBpedia, Yahoo).
@@
 After completing the final task in each stream, we report the average accuracy across all tasks.
-All experiments use one epoch per task with DeepSpeed, a fixed learning rate of $31e-$ 3, batch size 64, and dropout of 0.1.
-MoRA is applied to both the query and key projection matrices within attention layers, initializing $r=8$ new ranks for each incoming task similarly as in [74, 58].
+All experiments use one epoch per task with DeepSpeed, a fixed learning rate of 1e31e-3, batch size 64, and dropout of 0.1.
+MoRA is applied to both the query and key projection matrices within attention layers, initializing $r=8$ new ranks for each incoming task similarly as in (Wang et al., 2023; Qiao & Mahdavi, 2024).
 We maintain a constant activation budget of 4 ranks throughout continual learning, set the temperature $\tau_{\text{MoRA}}=0.1$, and the threshold $\delta=0.2$.
 Generalization and forgetting on unseen tasks.
-To assess effects on pre-trained general knowledge, we fine-tune Llama3.1-8B [28] on the CodeAlpaca code-generation dataset [9] using llama-Factory [91] and evaluate using lm-eval-harness [26] on zero-shot in-domain performance on HumanEval [12], as well as out-of-domain accuracy on a broad selection of MMLU [31] subjects—Formal Logic, Philosophy, World Religions, Economics, Public Relations, STEM, Physics, and Machine Learning.
+To assess effects on pre-trained general knowledge, we fine-tune Llama3.1-8B (Grattafiori et al., 2024) on the CodeAlpaca code-generation dataset (Chaudhary, 2023) using llama-Factory (Zheng et al., 2024) and evaluate using lm-eval-harness (Gao et al., 2024) on zero-shot in-domain performance on HumanEval (Chen et al., 2021), as well as out-of-domain accuracy on a broad selection of MMLU (Hendrycks et al., 2021) subjects—Formal Logic, Philosophy, World Religions, Economics, Public Relations, STEM, Physics, and Machine Learning.
 In this experiment, MoRA is applied to all linear weight matrices of the pre-trained model.
-We fine-tune on CodeAlpaca with a batch size of 32 over 3 epochs and a cosine learning-rate schedule starting at $45e-$ 4.
+We fine-tune on CodeAlpaca with a batch size of 32 over 3 epochs and a cosine learning-rate schedule starting at 5e45e-4.
 We train with $r=16$ ranks and enforce a constant activation budget of 4 ranks.
@@
 We observe that, due to variations in hidden representations across architectures, the optimal temperature setting can differ across each pre-trained models.
-Reproducibility Statement.
-Our source code will be publicly released upon paper acceptance.
-Results are averaged over three independent runs to ensure robustness.
-Experiments with the CLIP model were carried out on an NVIDIA GeForce RTX 3090 GPU, while experiments with LLMs utilized 4 NVIDIA H100 GPUs.
-Table 9: Average accuracy on T5-large continual-learning benchmarks after the final task, evaluated over extended 15-task sequences. Results for prior methods are taken from [58].
+Further details of the used datasets in MTIL and X-TAIL.
+The MTIL setting consists of 1,201 classes drawn from 11 diverse datasets: Aircraft (Maji et al., 2013), Caltech101 (Fei-Fei et al., 2004), CIFAR100 (Krizhevsky et al., 2009), DTD (Cimpoi et al., 2014), EuroSAT (Helber et al., 2019), Flowers (Nilsback & Zisserman, 2008), Food (Bossard et al., 2014), MNIST (Deng, 2012), OxfordPet (Parkhi et al., 2012), Cars (Krause et al., 2013), and SUN397 (Xiao et al., 2010).
+In the X-TAIL setting, a total of 10 datasets are used, with CIFAR100 (Fei-Fei et al., 2004) excluded to prevent domain overlap, following the protocol in (Xu et al., 2024).
+In line with (Xu et al., 2024), we use a 5-shot split for MTIL and a 16-shot split for X-TAIL.
+Table 8: Average accuracy on T5-large continual‐learning benchmarks after the final task, evaluated over extended 15-task sequences. Results for prior methods are taken from (Qiao & Mahdavi, 2024).
 Large Number of Tasks
@@
 MoRA 68.91 68.32 71.95 69.72
-B.2 Continual Learning of LLM on Long Task Sequences
+Table 9: Continual learning results on standard CL benchmarks with the LLaMA2-7B model.
+Method Order-1 Order-2 Order-3 Avg.
+O-LoRA 76.8 75.7 75.7 76.1
+MoRA 77.8 78.0 79.3 78.4
+A.2 Continual Learning of LMs on Long Task Sequences
 In Sec.
-4.2 Table 3, we evaluate MoRA on standard CL benchmarks [62].
-In Table 9, we extend the evaluation to challenged long task sequences using 15 datasets with 3 different orderings as in [74, 58].
+4.2 Table 3, we evaluate MoRA on standard CL benchmarks (Razdaibiedina et al., 2023).
+In Table 8, we extend the evaluation to challenged long task sequences using 15 datasets with 3 different orderings as in (Wang et al., 2023; Qiao & Mahdavi, 2024).
 Consistent with the findings in Table 3, MoRA outperforms previous methods in terms of averaged performance across three task orders, and largely close the gap to multi-task learning.
-Two key design choices drive this robust performance in the long-sequence regime.
-First, by decomposing each LoRA update into fine-grained rank-1 components and enforcing a small, fixed activation budget, MoRA encourages each rank to specialize on a narrow subspace of the data manifold.
+Two key design choices drive this robust performance in the long‐sequence regime.
+First, by decomposing each LoRA update into fine‐grained rank-1 components and enforcing a small, fixed activation budget, MoRA encourages each rank to specialize on a narrow subspace of the data manifold.
 At inference time, only the most relevant subspaces are activated for a given input, which preserves earlier task representations and prevents catastrophic interference.
-Second, our self-routed gating mechanism enables each rank to assess its own relevance on a per-token basis, yielding stable mixture patterns as the expert pool grows.
+Second, our self‐routed gating mechanism enables each rank to assess its own relevance on a per‐token basis, yielding stable mixture patterns as the expert pool grows.
 Coupling with our proposed rank pruning, these mechanisms ensure that MoRA continually incorporates new knowledge only when needed while robustly maintaining prior capabilities.
@@
 Method Trainable parameters per task
-LoRA [32] $r\,(d_{\text{in}}+d_{\text{out}})$
+LoRA (Hu et al., 2021) $r\,(d_{\text{in}}+d_{\text{out}})$
 MoE-LoRA (1 expert/task) $r\,(d_{\text{in}}+d_{\text{out}})\;+\;d_{\text{in}}$
-MoE-Adapter (2 experts/task) [85] $2(r\,(d_{\text{in}}+d_{\text{out}})+d_{\text{in}})$
-CoDyRA [48] $r\,(d_{\text{in}}+d_{\text{out}})$
-O-LoRA [74] $r\,(d_{\text{in}}+d_{\text{out}})$
-LB-CL [58] $r\,(d_{\text{in}}+d_{\text{out}})+r$
+MoE-Adapter (2 experts/task) (Yu et al., 2024) $2(r\,(d_{\text{in}}+d_{\text{out}})+d_{\text{in}})$
+CoDyRA (Lu et al., 2024) $r\,(d_{\text{in}}+d_{\text{out}})+r$
+O-LoRA (Wang et al., 2023) $r\,(d_{\text{in}}+d_{\text{out}})$
+LB-CL (Qiao & Mahdavi, 2024) $r\,(d_{\text{in}}+d_{\text{out}})+r$
 MoRA $rd_{\text{in}}+kd_{\text{out}}$
-B.3 Computation Cost
-Table 10 summarizes the per-task trainable parameters of various continual-learning methods.
-Standard LoRA [32], CoDyRA [48], and O-LoRA [74] each introduce $r(d_{\text{in}}+d_{\text{out}})$ new parameters per weight matrix.
-Mixture-of-Experts variants such as MoE-LoRA and MoE-Adapter [85] additionally train a router module to controll the usage of each LoRA experts, inducing $d_{\text{in}}$ additional paramters for each experts.
-LB-CL [58] introduce $r$ additional parameters, mimic the singular values of SVD.
+A.3 Multi‐domain task incremental learning.
+We evaluate MoRA in the few‐shot MTIL setting (Table 11) under the same protocols as (Yu et al., 2024; Xu et al., 2024; Lu et al., 2024).
+Consistent with the results observed in the X-TAIL setting, our method demonstrates clear superiority in this scenario.
+In this challenging scenario, the model must learn 11 diverse tasks sequentially, with only five examples per class.
+These findings validate that our self-activated sparse mixture-of-ranks framework both facilitates continual acquisition of new knowledge and mitigates forgetting from the pre-trained model and earlier tasks.
+Table 11: Comparisons on 5-shot MTIL setting. Following the same protocol as in (Yu et al., 2024; Xu et al., 2024; Lu et al., 2024).
+Method
+Aircraft Caltech101 CIFAR100 DTD EuroSAT Flowers Food MNIST OxfordPet Cars SUN397 Average
+CLIP
+   Zero-shot (Radford et al., 2021) 24.3 88.4 68.2 44.6 54.9 71.0 88.5 59.4 89.0 64.7 65.2 65.3
+Transfer
+   Zero-shot (Radford et al., 2021) 88.4 68.2 44.6 54.9 71.0 88.5 59.6 89.0 64.7 65.2 69.4
+   LwF (Li & Hoiem, 2017) 72.1 49.2 35.9 44.5 41.1 66.6 50.5 69.0 19.0 51.7 50.0
+   LwF-VR (Ding et al., 2022) 82.2 62.5 40.1 40.1 56.3 80.0 60.9 77.6 40.5 60.8 60.1
+   WiSE-FT (Wortsman et al., 2022) 77.6 60.0 41.3 39.4 53.0 76.6 58.1 75.5 37.3 58.2 57.7
+   ZSCL (Zheng et al., 2023) 84.0 68.1 44.8 46.8 63.6 84.9 61.4 81.4 55.5 62.2 65.3
+   MoE-Adapter (Yu et al., 2024) 87.9 68.2 44.1 48.1 64.7 88.8 69.0 89.1 64.5 65.1 68.9
+   RAIL-Primal (Xu et al., 2024) 88.4 68.2 44.6 54.9 71.0 88.5 59.6 89.0 64.7 65.2 69.4
+   CoDyRA (Lu et al., 2024) 92.4 68.4 45.8 54.5 69.6 87.4 65.2 88.5 64.2 64.5 69.9
+   MoRA 92.0 68.8 45.6 53.1 68.6 84.4 64.3 89.8 65.4 64.8 69.7
+Average
+   LwF (Li & Hoiem, 2017) 23.5 77.4 43.5 41.7 43.5 52.2 54.6 63.4 68.0 21.3 52.6 49.2
+   LwF-VR (Ding et al., 2022) 24.9 89.1 64.2 53.4 54.3 70.8 79.2 66.5 79.2 44.1 61.6 62.5
+   WiSE-FT (Wortsman et al., 2022) 32.0 87.7 61.0 55.8 68.1 69.3 76.8 71.5 77.6 42.0 59.3 63.7
+   ZSCL (Zheng et al., 2023) 28.2 88.6 66.5 53.5 56.3 73.4 83.1 56.4 82.4 57.5 62.9 64.4
+   MoE-Adapter (Yu et al., 2024) 30.0 89.6 73.9 58.7 69.3 79.3 88.1 76.5 89.1 65.3 65.8 71.4
+   RAIL-Primal (Xu et al., 2024) 32.9 94.5 69.9 58.1 71.8 84.4 88.5 70.4 89.0 66.1 65.7 71.9
+   CoDyRA (Lu et al., 2024) 34.6 95.8 73.9 60.0 77.1 81.3 86.6 75.9 89.9 66.1 65.3 73.3
+   MoRA 36.7 95.4 74.9 61.9 77.1 82.6 85.3 76.0 90.5 67.0 65.6 73.9
+Last
+   LwF (Li & Hoiem, 2017) 22.1 58.2 17.9 32.1 28.1 66.7 46.0 84.3 64.1 31.5 60.1 46.5
+   LwF-VR (Ding et al., 2022) 22.9 89.8 59.3 57.1 57.6 79.2 78.3 77.7 83.6 60.1 69.8 66.9
+   WiSE-FT (Wortsman et al., 2022) 30.8 88.9 59.6 60.3 80.9 81.7 77.1 94.9 83.2 62.8 70.0 71.9
+   ZSCL (Zheng et al., 2023) 26.8 88.5 63.7 55.7 60.2 82.1 82.6 58.6 85.9 66.7 70.4 67.4
+   MoE-Adapter (Yu et al., 2024) 30.1 89.3 74.9 64.0 82.3 89.4 87.1 89.0 89.1 69.5 72.5 76.1
+   RAIL-Primal (Xu et al., 2024) 32.9 95.1 70.3 63.2 81.5 95.6 88.5 89.7 89.0 72.5 71.0 77.2
+   CoDyRA (Lu et al., 2024) 31.6 95.5 72.8 63.5 85.0 89.7 85.0 94.7 93.2 73.6 73.0 78.0
+   MoRA 32.5 95.3 75.3 66.6 87.8 92.6 86.3 96.3 92.6 73.5 73.8 79.3
+A.4 More Comparison on HumanEval for Code Generation
+To further evaluate code generation performance, we compare MoRA against LoRI-D and LoRI-S (Zhang et al., 2025) on the HumanEval benchmark (Table 12).
+MoRA consistently outperforms both LoRI variants across all metrics, achieving notable improvements in Pass@1, Pass@5, and especially Pass@10.
+Table 12: Performance comparison on the HumanEval benchmark, reported in terms of Pass@1, Pass@5, and Pass@10.
+HumanEval Pass@1 Pass@5 Pass@10
+LoRI-D 43.2 57.6 63.2
+LoRI-S 41.3 54.4 59.6
+MoRA 47.6 60.9 70.1
+A.5 performance robustness
+MoRA employs a sparse mixture of previously learned and newly introduced rank-1 experts to capture both shared and task-specific knowledge, resulting in substantially improved Last performance.
+To assess statistical significance and robustness, we report mean and standard deviation over three independent runs (Table 13).
+MoRA consistently outperforms competing methods across all metrics and exhibits lower variance, highlighting its effectiveness and stability in continual-learning scenarios.
+Table 13: Comparison to InfLoRA and performance robustness. We report mean and standard deviation across 3 independent runs. Best performances are marked in bold.
+Method
+Cars Aircraft OxfordPet Food SUN397 MNIST Flowers DTD Caltech101 EuroSAT Average
+Transfer
+   InfLoRA $72.26^{\pm 0.56}$ $36.19^{\pm 0.64}$ $38.46^{\pm 0.39}$ $55.22^{\pm 1.65}$ $73.19^{\pm 0.55}$ $39.32^{\pm 1.54}$ $80.29^{\pm 0.91}$ $1.19^{\pm 1.16} 55$$05^{\pm 0.51} 55.$$5.69^{\pm 0.24}$
+   CoDyRA $74.3^{\pm 0.52}$ $36.8^{\pm 0.23}$ $44.2^{\pm 0.56}$ $\mathbf{69.9^{\pm 0.56}}$ $\mathbf{83.5^{\pm 0.23}}$ $42.8^{\pm 0.18}$ $\mathbf{88.9^{\pm 0.42}}$ $64.6^{\pm 0.47}$ $\mathbf{63.4^{\pm 0.56}}$ $63.2^{\pm 0.28}$
+   MoRA $\mathbf{74.5^{\pm 0.51}}$ $\mathbf{38.1^{\pm 0.24}}$ $\mathbf{46.9^{\pm 0.56}}$ $65.3^{\pm 0.44}$ $82.9^{\pm 0.18}$ $\mathbf{45.8^{\pm 0.31}}$ $88.2^{\pm 0.15}$ $\mathbf{65.1^{\pm 0.35}}$ $62.9^{\pm 0.10}$ $\mathbf{63.3^{\pm 0.26}}$
+Average
+   InfLoRA $20.49^{\pm 0.98}$ $78.58^{\pm 1.02}$ $48.5^{\pm 1.18}$ $66.59^{\pm 1.51}$ $71.83^{\pm 0.80}$ $76.79^{\pm 0.34}$ $61.45^{\pm 1.36}$ $82.59^{\pm 0.86}$ $55.3^{\pm 1.34}$ $56.67^{\pm 0.59}$ $62.48^{\pm 0.31}$
+   CoDyRA $41.4^{\pm 0.28}$ $81^{\pm 0.38}$ $58.7^{\pm 0.26}$ $77.8^{\pm 0.47}$ $83.4^{\pm 0.39}$ $\mathbf{84.6^{\pm 0.28}}$ $64.5^{\pm 0.14}$ $\mathbf{90.4^{\pm 0.40}}$ $67.2^{\pm 0.23}$ $\mathbf{64.4^{\pm 0.47}}$ $71.3^{\pm 0.18}$
+   MoRA $\mathbf{44.1^{\pm 0.24}}$ $\mathbf{81.6^{\pm 0.34}}$ $\mathbf{64.6^{\pm 0.34}}$ $\mathbf{79.6^{\pm 0.37}}$ $\mathbf{83.9^{\pm 0.36}}$ $84.4^{\pm 0.15}$ $\mathbf{66.5^{\pm 0.24}}$ $89.7^{\pm 0.07}$ $\mathbf{68.4^{\pm 0.38}}$ $64.1^{\pm 0.09}$ $\mathbf{72.7^{\pm 0.17}}$
+Last
+   InfLoRA $18.26^{\pm 0.49}$ $\mathbf{82.36^{\pm 0.92}}$ $46.57^{\pm 0.89}$ $79.38^{\pm 2.22}$ $76.16^{\pm 1.61}$ $79.58^{\pm 0.60}$ $95.74^{\pm 0.44}$ $87.78^{\pm 0.85}$ $71.11^{\pm 0.73}$ $73.05^{\pm 0.19}$ $70.99^{\pm 0.24}$
+   CoDyRA $\mathbf{37.7^{\pm 0.42}}$ $81.5^{\pm 0.24}$ $65.1^{\pm 0.63}$ $89.9^{\pm 0.55}$ $91.4^{\pm 0.38}$ $85.5^{\pm 0.16}$ $96.8^{\pm 0.08}$ $\mathbf{93.3^{\pm 0.30}}$ $77.3^{\pm 0.66}$ $73.5^{\pm 0.21}$ $79.2^{\pm 0.18}$
+   MoRA $\mathbf{37.7^{\pm 0.28}}$ $81.5^{\pm 0.22}$ $\mathbf{70.7^{\pm 0.49}}$ $\mathbf{92.4^{\pm 0.20}}$ $\mathbf{95^{\pm 0.34}}$ $\mathbf{86^{\pm 0.13}}$ $\mathbf{97.6^{\pm 0.19}}$ $92.6^{\pm 0.10}$ $\mathbf{81^{\pm 0.35}}$ $\mathbf{74.7^{\pm 0.06}}$ $\mathbf{80.9^{\pm 0.12}}$
+A.6 More comparisons on X-TAIL
+To further validate MoRA’s robustness, we evaluated it under other continual-learning task orderings, i.e., X-TAIL (Order 2), as shown in Table 14.
+The results align with those in Table 1, confirming that MoRA consistently achieves state-of-the-art performance.
+Table 14: Comparisons on X-TAIL (Order 2) for each domain in terms of “Transfer”, “Average”, and “Last” scores (%).
+Method
+Cars Aircraft OxfordPet Food SUN397 MNIST Flowers DTD Caltech101 EuroSAT Average
+CLIP
+   Zero-shot 66.1 23.5 86.7 84 63.7 46.7 63.6 37.3 76.8 36.7 58.5
+Transfer
+   Zero-shot (Radford et al., 2021) 23.5 86.7 84 63.7 46.7 63.6 37.3 76.8 36.7 57.7
+   LwF (Li & Hoiem, 2017) 20.0 74.1 79.6 58.1 34.1 48.9 27.7 64.4 15.1 46.9
+   WiSE-FT (Wortsman et al., 2022) 21.3 79.5 83.3 61.0 39.9 56.5 29.6 68.0 20.8 51.1
+   ZSCL (Zheng et al., 2023) 23.0 84.3 87.2 63.0 42.1 65.2 34.6 71.4 40.9 56.9
+   MoE-Adapter (Yu et al., 2024) 17.1 87.2 87.5 58.4 12.6 65.5 35.9 70.0 17.9 50.2
+   RAIL-Primal (Xu et al., 2024) 23.5 86.7 84 63.7 46.7 63.6 37.3 76.8 36.7 57.7
+   CoDyRA (Lu et al., 2024) 23.6 89.2 83 62 51 71.4 38 77.4 39 59.4
+   MoRA 23.6 88.7 83.4 62.6 51.2 69.9 39.3 77.5 39 59.5
+Average
+   LwF (Li & Hoiem, 2017) 49.0 27.4 69.7 83.0 65.7 42.2 63.5 33.1 68.5 17.5 52.0
+   WiSE-FT (Wortsman et al., 2022) 57.9 29.6 77.8 85.4 68.0 51.6 69.3 35.5 71.0 23.0 56.9
+   ZSCL (Zheng et al., 2023) 74.4 36.4 86.7 88.7 68.9 50.0 75.1 40.1 72.5 43.7 63.6
+   MoE-Adapter (Yu et al., 2024) 74.4 38.6 87.7 87.3 67.9 50.6 76.5 43.7 72.3 18.8 61.8
+   RAIL-Primal (Xu et al., 2024) 77.9 40.4 85.6 83.3 68.3 62.2 76.6 45.8 80.4 41.7 66.2
+   CoDyRA (Lu et al., 2024) 80 39.2 92.5 85.2 69.2 73.7 79.6 46.2 78.6 44.1 68.8
+   MoRA 80.2 40.1 92.5 84.7 70.1 74 80.1 48.7 78.4 44.4 69.3
+Last
+   LwF (Li & Hoiem, 2017) 29.6 17.5 63.0 83.8 67.7 44.9 79.3 44.8 84.6 39.0 55.4
+   WiSE-FT (Wortsman et al., 2022) 46.1 23.5 71.3 85.7 70.2 59.1 85.5 47.9 82.4 42.8 61.5
+   ZSCL (Zheng et al., 2023) 71.7 35.3 86.5 89.2 71.8 52.3 89.8 52.0 77.1 68.4 69.4
+   MoE-Adapter (Yu et al., 2024) 75.1 41.1 87.9 87.1 74.1 89.7 92.6 61.2 81.0 27.4 71.7
+   RAIL-Primal (Xu et al., 2024) 77.7 41.9 86.1 83.3 71.8 91.6 97.3 66.4 94.8 86.9 79.8
+   CoDyRA (Lu et al., 2024) 79 38.6 92.6 86.4 74.7 95.2 93 64.7 81.9 92.2 79.8
+   MoRA 79.3 38.9 93.1 85.4 74.9 96.4 94.1 69.9 82 92.9 80.7
+Table 15: Trainable parameters and averaged training GPU memory per task.
+Method Trainable Params. (Million) GPU Mem. (MiB)
+LWF (Li & Hoiem, 2017) 129.6 32172
+ZSCL (Zheng et al., 2023) 129.6 26290
+MoE-Adapters (Yu et al., 2024) 59.8 22358
+CoDyRA (Lu et al., 2024) 4.4 21770
+MoRA 4.4 21090
+A.7 Computation Cost
+Table 10 summarizes the per‐task trainable parameters of various continual‐learning methods.
+Standard LoRA (Hu et al., 2021), CoDyRA (Lu et al., 2024), and O-LoRA (Wang et al., 2023) each introduce $r(d_{\text{in}}+d_{\text{out}})$ new parameters per weight matrix.
+Mixture‐of‐Experts variants such as MoE-LoRA and MoE-Adapter (Yu et al., 2024) additionally train a router module to controll the usage of each LoRA experts, inducing $d_{\text{in}}$ additional paramters for each experts.
+LB-CL (Qiao & Mahdavi, 2024) introduce $r$ additional parameters, mimic the singular values of SVD.
 By contrast, MoRA requires only $rd_{\text{in}}+kd_{\text{out}}$ activated trainable parameters per task, where $k\leq r$ is the activation budget, and the trainable parameter is at most the same as a standard LoRA.
-Despite the small number of parameter activated and trained, MoRA achieves superior continual learning performance, and reaches comparable performance in general fine-tuning with only one-third of the activated parameters of a standard LoRA.
+Despite the small number of parameter activated and trained, MoRA achieves superior continual learning performance, and reaches comparable performance in general fine-tuning with only one‐third of the activated parameters of a standard LoRA.
+Trainable parameters and training GPU memory.
+Beyond the estimated parameter counts in Table 10, in Table 15, we measured the actual trainable parameters for each continual‐learning task and GPU memory usage, under the same settings as Table 1 in the main paper.
+LWF (Li & Hoiem, 2017) and ZSCL (Zheng et al., 2023) perform full‐parameter fine‐tuning, consuming the most parameters and memory.
+MoE‐Adapters (Yu et al., 2024) maintains a router with 22 rank‐64 adapter experts (top-2 activated) and a DDAS domain predictor.
+CoDyRA (Lu et al., 2024) trains a single rank‐16 LoRA per task, reducing its footprint to 4.4 M parameters.
+MoRA introduces 16 rank-1 experts per task, with no additional router, for a total of 4.4 M trainable parameters and keeps a low gpu memory usage, thanks to our novel self‐activated sparse mixture of ranks design.
 […]
@@
 Figure 5: Extended view of Fig. 3 illustrating knowledge reuse. Regions corresponding to generic input tokens (e.g. blue sky) are highlighted with blue bounding boxes. Zoom in for details.
-B.4 Extended Visualizations of Rank Activations
+A.8 Extended Visualizations of Rank Activations
 In Sec.4.4 (Fig.3), we illustrated rank activations during the learning of Task 1 and Task 2. Here, we extend these visualizations to additional tasks and scenarios in Fig.4 and Fig.5.
-MoRA retains task-specific semantics without forgetting.
+MoRA retains task‐specific semantics without forgetting.
 Fig.
@@
 Patches corresponding to the airplane object are outlined in orange.
-In all three snapshots, Rank at index 0 remains consistently and exclusively activated for those airplane patches, demonstrating that MoRA has effectively memorized the airplane-specific knowledge into Rank 0.
+In all three snapshots, Rank at index 0 remains consistently and exclusively activated for those airplane patches, demonstrating that MoRA has effectively memorized the airplane‐specific knowledge into Rank 0.
 Even after 10 subsequent tasks, this pattern remains unchanged, indicating that later updates do not overwrite or interfere with the learned airplane representations.
-In other words, MoRA effectively memorizes and preserves task-relevant semantics, thereby mitigating catastrophic forgetting.
+In other words, MoRA effectively memorizes and preserves task‐relevant semantics, thereby mitigating catastrophic forgetting.
 MoRA encodes generic semantics that are reused across tasks.
@@
 5 examines an input image from Task 9 before and after learning Task 9.
-Panel (a) shows the activation map of data from Task 1 after learning Task 1: Rank 11 (outlined in blue) already responds strongly to sky-background patches, demonstrating that MoRA has stored a generic “blue sky” concept in this rank.
+Panel (a) shows the activation map of data from Task 1 after learning Task 1: Rank 11 (outlined in blue) already responds strongly to sky‐background patches, demonstrating that MoRA has stored a generic “blue sky” concept in this rank.
 In panel (b), when we infer on the Task 9 image before training on Task 9, Rank 11 is again activated for the sky regions, confirming that MoRA reuses this shared knowledge for unseen data.
@@
 Figure 8: Required ranks to capture 99 % of cumulative activations, shown across different pre-trained model layers and projection locations.
-B.5 Statistical Analyses of Contributing Rank Activations
+A.9 Statistical Analyses of Contributing Rank Activations
 Fig.
@@
 For example, in Fig.
-7, the MLP’s output projection (c-proj) in Layer 1 of the vision encoder requires 16 ranks, whereas the same module in Layer 1 of the text encoder needs only 3 ranks.
+6, the MLP’s output projection (c-proj) in Layer 1 of the vision encoder requires 16 ranks, whereas the same module in Layer 1 of the text encoder needs only 3 ranks.
 To provide a broader view, Fig.
@@
 This adaptive sparsity maximizes the efficient use of newly acquired knowledge during continual learning.
-Broader Impacts
-MoRA’s self-activated, sparse mixture-of-ranks method makes continual adaptation of large vision–language and language models far more efficient and reliable.
-By updating only the most relevant low-rank subspaces, it reduce compute and memory requirements, enabling on-device personalization in domains like education, healthcare, and finance.
-This democratizes access to powerful, continually evolved models for smaller teams and resource-constrained settings, driving innovation and improving quality of life.
-Preserving pre-trained knowledge also ensures stability in safety-critical systems, such as medical diagnostics and autonomous vehicles, where forgetting of pre-trained capabilities can be dangerous.
-The focus of this paper is fundamental research, and is broadly applicable to model fine-tuning techniques.
-Thus, it is possible that MoRA’s lightweight updates could be misused to insert stealthy backdoors or reinforce hidden biases that are hard to detect.
-Because it leaves most pre-trained parameters untouched, existing biases may persist or become entrenched.
-To mitigate these risks, we recommend strict audit logging of rank-level updates, anomaly detection on sparse changes, robust access controls, and routine bias and fairness assessments alongside any deployment.

### v3

@@
-Little By Little: Continual Learning via Self-Activated Sparse Mixture-of-Rank Adaptive Learning
+Little By Little: Continual Learning via Incremental Mixture of Rank-1 Associative Memory Experts
 Abstract
-Continual learning (CL) with large pre-trained models is challenged by catastrophic forgetting and task interference.
-Existing LoRA-based Mixture-of-Experts (MoE) approaches mitigate forgetting by assigning and freezing task-specific adapters, but suffer from interference, redundancy, and ambiguous routing due to coarse adapter-level selection.
-However, this design introduces three key challenges: 1) Interference: Activating full LoRA experts per input leads to subspace interference and prevents selective reuse of useful components across tasks.
-2) Redundancy: Newly added experts often duplicate or contradict existing knowledge due to unnecessary activation of unrelated ranks and insufficient reuse of relevant ones.
-3) Ambiguity: Overlapping features across tasks confuse the router, resulting in unstable expert assignments.
-As more experts accumulate, earlier task routing degrades, accelerating forgetting.
-We propose MoRA, a Mixture-of-Rank Adaptive learning approaches with self-activated and sparse rank activation for CL.
-Unlike mixing multiple low-rank matrices, MoRA decomposes each rank-r update into r rank-one components, each treated as an independent expert, enabling fine-grained rank-one expert utilization while mitigating interference and redundancy.
-To avoid ambiguous routing, we propose that each rank-one expert can infer its own relevance via intermediate activations.
-Coupled with our proposed rank pruning and activation budgets, MoRA adaptively selects a sparse mixture of ranks per input.
-We validate MoRA on continual learning benchmarks using CLIP and language models, analyzing both in-domain learning and out-of-domain forgetting/generalization during fine-tuning.
-MoRA shows significant effectiveness in enhancing CL with PTMs, and improving generalization while mitigating forgetting.
+Continual learning (CL) with large pre-trained models is challenged by task interference and catastrophic forgetting.
+Existing LoRA-based Mixture-of-Experts (MoE) methods mitigate forgetting by adding new task-specific adapters and freezing old ones, but often suffer from redundancy, interference, and ambiguous routing due to coarse-grained experts and routing.
+Coarse-grained experts (i.e., full LoRA adapters with large rank) encode low-specialty information.
+Newly added experts often duplicate or conflict with existing ones, causing redundancy and interference.
+Their low specialization further confuses the router, accelerating routing degradation and forgetting as experts accumulate.
+In this work, we propose MoRAM (Mixture of Rank-1 Associative Memory).
+Grounded in the view that weight matrices function as linear associative memories, MoRAM achieves CL as gradual incrementing of atomic rank-1 memory experts.
+Each rank-1 adapter acts as a fine-grained MoE expert or an associative memory unit.
+By viewing rank-1 adapters as key–value pairs, we eliminate explicit routers in MoE-LoRA, using a self-activation mechanism where each memory atom evaluates its own relevance via its intrinsic key.
+This transforms the adaptation process into robust, content-addressable retrieval.
+Extensive experiments on CLIP and LLMs demonstrate that MoRAM significantly outperforms state-of-the-art baselines, achieving superior plasticity-stability trade-offs, improving generalization while mitigating forgetting.
 “Little by little, we gave you everything you ever dreamed of …” — “Little by little”, Oasis.
 1 Introduction
-Continual learning (CL) (Hadsell et al., 2020; De Lange et al., 2021; Ding et al., 2022) aims to enable models to incrementally and efficiently acquire new knowledge from a stream of tasks, without catastrophic forgetting (Nguyen et al., 2019; McCloskey & Cohen, 1989) or the need for repeated fine-tuning on all previously seen data (Wang et al., 2022e; 2024).
-The emergence of large pre-trained models (PTMs), including Vision Transformer (ViT) for vision (Dosovitskiy et al., 2020), Language Model (LM) (Grattafiori et al., 2024; Raffel et al., 2020), and CLIP for vision–language embeddings (Radford et al., 2021), has spurred a wave exploring CL with PTMs (Xu et al., 2024; Wang et al., 2023).
-Despite the success of PTMs, CL on them across sequential tasks still faces key challenges, including preventing catastrophic forgetting and preserving, or even enhancing, the pre-trained capabilities to generalize to unseen or future cases (Qiao & Mahdavi, 2024).
-A common way to integrate new knowledge into PTMs is full fine-tuning (Zheng et al., 2023; Garg et al., 2023) or PEFT with a small set of learnable parameters (Yang et al., 2024b), such as LoRA (Hu et al., 2021) (Fig.
-1(a)).
-However, continually updating the same parameters often leads to interference and forgetting, even with a single new task (Biderman et al., 2024).
-Leveraging LoRA’s regularization and flexibility, many CL methods adopt LoRA or its variants with additional strategies to mitigate forgetting (Yu et al., 2024; Wang et al., 2023; Qiao & Mahdavi, 2024).
-Inspired by mixture-of-experts (MoE) architectures (Shazeer et al., 2017; Lepikhin et al., 2020; Fedus et al., 2022; Dai et al., 2024) and the modular nature of LoRA adapters, MoE frameworks using LoRA adapters as experts have been studied (Dou et al., 2023; Wu et al., 2024b) and applied in CL (Fig.
-1(b)) have been widely adopted in CL (Yu et al., 2024; Wang et al., 2024; Yang et al., 2024a; Chen et al., 2024; Li et al., 2025).
-These works either predefine an MoE with LoRA for CL (Yang et al., 2024a), or incrementally add experts (Wang et al., 2024) or task-specific routers (Yu et al., 2024), assuming MoE benefits CL by isolating task interference.
-Such methods (Wang et al., 2024; Qiao & Mahdavi, 2024; Rusu et al., 2016; Yu et al., 2024) freeze old components and add new ones (e.g., experts or routers) to reduce forgetting.
-Despite design differences, we collectively refer a plain and general design with LoRA-based MoE as MoE-LoRA.
-In MoE-LoRA models, each expert is a LoRA adapter with multiple ranks (subspaces), and the router selects among experts with each LoRA adapter as a unit.
-Despite their success, we argue that the coarse granularity of experts and routing introduces three key challenges that limit effective utilization:
-(1) Interference across experts and limited knowledge reuse.
-A coarse-grained multi-rank LoRA expert is packed with a rich set of knowledge that must be fully activated or not at all.
-For each sample (i.e., token), any two experts may hold both complementary and conflicting knowledge, creating a dilemma: selecting both causes interference, while only selecting either one prevents reuse.
-(2) Learning redundancy and conflicts.
-As a result of (1), (newly added) experts often have to carry conflicting or redundant information, either from activating other experts with both useful and unnecessary content or from insufficient reuse of existing ones.
-This coarse selection exacerbates interference and forgetting during inference.
-(3) Routing ambiguity and forgetting.
-With coarse-grained experts and routing, redundancy and conflicts hinder precise input–module assignment, leading to suboptimal mixtures (Lepikhin et al., 2020; Fedus et al., 2022; Dai et al., 2024).
-As more experts accumulate in CL, router assignments for earlier tasks grow unreliable, increasing forgetting.
-The three challenges are intertwined causes and effects of coarse-grained experts and routing, limiting the potential of promising MoE-LoRA design.
-Figure 1: Conceptional illustration of CL with (a) LoRA, (b) MoE-LoRA, and (c) MoRA (Ours).
-We propose MoRA, a Mixture-of-Rank Adaptive learning approaches with self-activated and sparsely selected ranks across the LoRAs added at different tasks.
-To address the limitations of coarse-grained MoE-LoRA methods, MoRA operates at a fine-grained level by treating each rank-one subspace within a LoRA adapter as an individual expert (or treating each rank-one LoRA as an expert).
-Expert selection and routing occur at the rank level, avoiding interference from whole LoRA blocks.
-This decomposition promotes specialization on narrower data regions, reduces redundancy, and enables effective knowledge reuse across tasks.
-We introduce a self-activation mechanism to select a sparse set of ranks without relying on a separately trained router.
-By defining each rank-one component as an expert, we avoid a separate router for expert assignment: the rank-one $\mathbf{A}$ in LoRA can act as a weighting router for its $\mathbf{B}$.
-Each rank evaluates its relevance to the input and determines its contribution, grounded in the view of low-rank updates as linear key–value memory.
-This design unifies rank-one LoRA with fine-grained MoE while preserving the original characteristics, saving parameters (for additional router), reducing routing confusion, and further mitigating forgetting (Fig.
+Continual learning (CL) [29, 16, 20] aims to enable models to incrementally and efficiently acquire new knowledge from a stream of tasks, without catastrophic forgetting [57, 54] or the need for repeated fine-tuning on all previously seen data [85, 78].
+With the ascendancy of Large Pre-trained Models (PTMs) in vision [21, 63] and language [64, 28], the challenge has shifted: how do we efficiently insert specific new capabilities into a massive, static memory structure without disrupting its existing generalized knowledge?
+Figure 1: Conceptual illustration of CL with (a) LoRA, (b) MoE-LoRA, and (c) MoRAM (Ours).
+A dominant approach for adapting PTMs is Parameter-Efficient Fine-Tuning (PEFT), particularly Low-Rank Adaptation (LoRA, Fig.
+1(a)) [33].
+While efficient, standard LoRA applies updates as dense modifications to the weight space.
+This creates a fundamental conflict: to learn a new task, the model must inevitably alter the representation space used by previous tasks, leading to interference and forgetting [6].
+To mitigate this, recent works have adopted Mixture-of-Experts MoE frameworks using LoRA adapters as experts (Fig.
+1(b)) have been studied [22, 88] and have been widely adopted in CL [96, 78, 94, 12, 43].
+These works either predefine an MoE with LoRA for CL [94], or incrementally add experts [78] or task-specific routers [96], assuming MoE benefits CL by isolating task interference.
+Such methods [78, 61, 67, 96] freeze old components and add new ones (e.g., experts or routers) to reduce forgetting.
+Despite design differences, we collectively refer to a plain and general design with LoRA-based MoE as MoE-LoRA.
+In MoE-LoRA models, each expert is a LoRA adapter with multiple ranks (in subspaces), and the router selects among experts with each LoRA adapter as a unit.
+However, current MoE-LoRA approaches operate at a coarse granularity that limits their effectiveness in CL.
+Treating entire rank-$r$ adapters as indivisible experts forces each coarse-grained expert to be activated and learned as a single unit.
+It enables each expert to learn a broad range of information, leading to low specialty and combinatorial expressivity [30, 51, 15].
+It specifically introduces three challenges:
+(1) Interference: A coarse expert contains a mix of low-specialty knowledge.
+Activating coarse-grained experts (i.e., LoRA adapters with large rank) for a specific input inevitably triggers irrelevant subspaces, causing interference.
+(2) Redundancy: New experts cannot selectively reuse specific atoms of knowledge from old experts; if the limited combination cannot cover the new task, they must relearn entire blocks, leading to inefficient capacity usage.
+(3) Routing/Retrieval Collapse: The low-specialization of coarse-grained experts further confuse the router.
+As the number of experts grows, routers increasingly struggle to reliably index the expanding expert pool (i.e., “routing ambiguity”), leading to routing collapse and drift of old experts, and ultimately accelerating forgetting.
+The coarse-grained experts and these resulting challenges limit the potential of the promising MoE-LoRA design.
+To address the limitations, we fundamentally rethink adaptation by synthesizing the intrinsic dimensionality of PTMs [1] through the lens of Linear Associative Memory [37].
+While PTMs reside in a high-dimensional parameter space, they operate on a low-rank manifold, effectively functioning as a composition of atomic key-value pairs of a memory system [37].
+From this perspective, the optimal CL update is not a “dense matrix” or “coarse adapters”, but an associative memory system with expandable and retrievable atomic memory units.
+Conventional methods that update entire matrices of PTM inevitably alter the global interference patterns of the memory, easily causing catastrophic forgetting.
+Coarse-grained MoE-LoRA methods suffer from brittle expert expansion and retrieval.
+From a memory-augmentation perspective, we treat CL as incremental external memory expansion with fine-grained atomic memory units for specialization and efficiency, while casting inference as input-specific memory retrieval and application.
+Based on this insight, we propose MoRAM (Mixture of Rank-1 Associative Memory), a framework that continually updates PTMs by incrementing an associative memory system with atomic rank-1 experts “little by little” (Fig.
 1(c)).
-Sparsity is enforced via temperature-scaled softmax together with top-k masking and rank pruning under a rank activation budget.
-As a complement, test-time thresholding further improves adaptability across cases.
-Leveraging the connection between PTMs and low-dimensional manifolds (Li et al., 2018; Aghajanyan et al., 2020), MoRA enables precise, input-specific adaptation through selective subspace updates.
-With fine-grained rank-level experts and sparse activation, MoRA enables continual learning by reusing old ranks and selectively activating a few new ones per sample, allowing efficient “little-by-little” adaptation with minimal redundancy.
-This sample-specific activation at inference also preserves pre-trained knowledge and improves generalization to unseen tasks.
-We validate MoRA on CL with both vision–language CLIP and LMs, analyzing its impact on pre-trained knowledge during low-rank fine-tuning.
-MoRA delivers strong downstream performance with far fewer active parameters, greatly reduces forgetting of prior tasks and pre-trained knowledge, and even improves generalization to unseen domains.
+The associative memory can also be seen as a mixture-of-expert model with rank-1 adapters as experts.
+Each rank-1 adapter acts as a fine-grained MoE expert or an atomic memory unit.
+By viewing rank-1 experts as key–value memory pairs, we can naturally eliminate the additional explicit routers in MoE-LoRA.
+With the proposed self-activation mechanism, each expert can evaluate its own relevance to the input via its intrinsic key.
+This shifts the paradigm from address-based routing (learning where to send data) to content-addressable retrieval (inputs automatically triggering the correct memory).
+This operation eliminates the source of routing collapse and forgetting from the additional router, increasing the reliability of the memory retrieval, even as the memory bank expands, while reducing computation.
+Our contributions are as follows: (1) We introduce MoRAM, a novel CL framework that treats Rank-1 updates as atomic associative memory experts, enabling fine-grained knowledge reuse and minimizing interference.
+(2) We propose a sparse self-activation routing mechanism that eliminates external routers, allowing the model to sparsely retrieve relevant memory atoms based on input content alone.
+(3) We conduct extensive evaluations on both vision–language (CLIP) and large language model (LLM) benchmarks.
+Empirical results demonstrate that MoRAM significantly outperforms state-of-the-art methods, achieving superior plasticity-stability trade-offs, while effectively preserving pre-trained generalization capabilities.
 2 Related Work
 Continual learning enables sequential knowledge acquisition without forgetting.
-Experience replay (ER) methods (Luo et al., 2023; Aljundi et al., 2019b; Chaudhry et al., 2018a; Liu et al., 2020; Chaudhry et al., 2018b; Yan et al., 2022; 2021) interleave past examples with new data.
-Parameter regularization (Kirkpatrick et al., 2017; Aljundi et al., 2018; Zenke et al., 2017; Aljundi et al., 2019a; Jha et al., 2023) penalizes updates to critical weights.
-Dynamic networks (Zhou et al., 2024; Wang et al., 2022a; b; Zhou et al., 2022; Wang et al., 2024; McDonnell et al., 2024; Liang & Li, 2024; Wang et al., 2022e; d; Smith et al., 2023; Wang et al., 2022c; 2024) allocate new capacity on the fly and preserve dedicated pathways for prior tasks.
+Experience replay (ER) methods [52, 4, 9, 46, 10, 92, 93] interleave past examples with new data.
+Parameter regularization [36, 2, 97, 3, 35] penalizes updates to critical weights.
+Dynamic networks [106, 76, 77, 107, 78, 55, 45, 85, 84, 69, 82, 78] allocate new capacity on the fly and preserve dedicated pathways for prior tasks.
 Continual learning of PTMs.
-For CL on vision–language CLIP model (Jha et al., 2024; Zhang et al., 2024b; Garg et al., 2023), methods like ZSCL (Zheng et al., 2023) retain zero-shot performance during adaptation, and follow-up work (Yu et al., 2024; Tang et al., 2025; Xu et al., 2024; Lu et al., 2024) continually fine-tunes while leveraging frozen pre-trained predictions.
-The X-TAIL benchmark (Xu et al., 2024) further challenges models by mixing domain labels at test time.
-In language models (LMs) (de Masson D’Autume et al., 2019; Qin & Joty, 2021; Razdaibiedina et al., 2023; Wang et al., 2023; Qiao & Mahdavi, 2024), continual learning uses capacity expansion or task-specific submodules to reduce interference.
-Low‐rank adaptation (LoRA) (Hu et al., 2021) is widely used for parameter‐efficient fine‐tuning of large pre‐trained models.
-Building on this foundation, recent methods reformulate LoRA’s updates via SVD‐based initialization and dynamic rank scheduling (Zhang et al., 2024a; Meng et al., 2024; Ding et al., 2023; Zhang et al., 2023; Liu et al., 2024; Wu et al., 2024a), showing that task adaptation primarily lies in a compact subspace.
-In this work, we offer a complementary perspective by viewing both pre‐trained weight matrix and its low‐rank updates as a linear associative memory (Li et al., 2018; Aghajanyan et al., 2020; Kohonen, 1972; Anderson, 1972).
-Under this lens, a rank-$r$ update corresponds to writing $r$ new key–value entries into the pre-trained memory matrix, with each rank-one component acting as an independent memory slot.
-Mixture‐of‐experts (MoE) in LoRA fine‐tuning.
-MoE increases model capacity by replacing the Transformer’s feed-forward layer with expert subnetworks, routing each input to a small subset and using load‐balancing losses to even out usage (Shazeer et al., 2017; Lepikhin et al., 2020; Fedus et al., 2022; Dai et al., 2024).
-This paradigm has been adapted for standard fine-tuning (Chen et al., 2023; Li et al., 2024; Dou et al., 2023) and CL (Yu et al., 2024; Wang et al., 2024; Yang et al., 2024a; Chen et al., 2024) with LoRA, treating each low‐rank adapter as an expert: new adapters are trained and frozen per task to prevent forgetting.
-In contrast, we decompose each rank‐$r$ update into $r$ rank‐11 components and compute an input‐dependent mixture over these fine‐grained experts, greatly enhancing both expert specialization and mixture diversity.
+For CL on vision–language CLIP model [34, 101, 27], methods like ZSCL [105] retain zero-shot performance during adaptation, and follow-up work [96, 70, 91, 50, 89] continually fine-tunes while leveraging frozen pre-trained predictions.
+The X-TAIL benchmark [91] further challenges models by mixing domain labels at test time.
+In language models (LMs) [17, 62, 65, 80, 61], continual learning uses capacity expansion or task-specific submodules to reduce interference.
+Low‐rank adaptation (LoRA) [33] is widely used for parameter-efficient fine-tuning of large pre-trained models.
+Building on this foundation, recent methods have reformulated LoRA’s updates via SVD-based initialization and dynamic rank scheduling [98, 56, 19, 100, 47, 87], demonstrating that task adaptation primarily relies on finding a compact subspace.
+In this work, we offer a complementary perspective by treating both the pre-trained weight matrix and its low-rank updates through the lens of linear associative memory [41, 1, 37, 5].
+Under this formulation, a rank-$r$ update corresponds to $r$ new memory entries into the matrix, where each rank-1 component is an atomic memory slot.
+Mixture-of-Experts (MoE) with LoRA.
+MoE scales capacity by routing inputs to sparse expert subnetworks via load-balancing [68, 40, 24, 15].
+This paradigm has been adapted for LoRA fine-tuning [13, 42, 22] and CL [96, 78, 94, 12] by treating adapters as experts, typically frozen per task to prevent forgetting.
+In contrast, we decompose rank-$r$ updates into $r$ atomic rank-1 components and compute an input-dependent mixture over these fine-grained experts, significantly enhancing specialization and diversity.
 3 Methods
@@
 In CL, a model sequentially learns $T$ tasks.
-For task $t\in{1,\dots,T}$, let $\mathcal{D}^{t}={(\mathbf{x}^{t}_{i},y^{t}_{i})}_{i=1}^{N^{t}}$, where $\mathbf{x}^{t}_{i}\in\mathbb{R}^{n\times d}$, $y^{t}_{i}\in\mathcal{C}^{t}$, and $N^{t}$ is the number of examples.
+For task $t\in{1,\dots,T}$, let $\mathcal{D}^{t}={\{}(\mathbf{x}^{t}_{i},y^{t}_{i}){\}}_{i=1}^{N^{t}}$, where $\mathbf{x}^{t}_{i}\in\mathbb{R}^{n\times d}$, $y^{t}_{i}\in\mathcal{C}^{t}$, and $N^{t}$ is the number of examples.
 In the memory-free setting, the model may access only $\mathcal{D}^{t}$ and cannot access data from any $\mathcal{D}^{u}$ with $u<t$.
 Low-rank adaptation.
-LoRA (Hu et al., 2021) parameterizes a low‐rank update to a pre‐trained weight matrix $\mathbf{W}_{0}\in\mathbb{R}^{d_{out}\times d_{in}}$ by introducing two factors $\mathbf{B}\in\mathbb{R}^{d_{out}\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times d_{in}}$, such that $\Delta\mathbf{W}=\mathbf{B}\mathbf{A}$, where $r\ll\min(d_{in},d_{out})$.
+LoRA [33] parameterizes a low‐rank update to a pre‐trained weight matrix $\mathbf{W}_{0}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$ by introducing two factors $\mathbf{B}\in\mathbb{R}^{d_{\text{out}}\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times d_{\text{in}}}$, such that $\Delta\mathbf{W}=\mathbf{B}\mathbf{A}$, where $r\ll\min(d_{\text{in}},d_{\text{out}})$.
 The updated weight matrix is then defined as:
 $\mathbf{W}=\mathbf{W}_{0}+\Delta\mathbf{W}=\mathbf{W}_{0}+\mathbf{B}\mathbf{A}.$ (1)
-In this formulation, the original weights $\mathbf{W}_{0}$ remain fixed, and only $\mathbf{B}$ and $\mathbf{A}$ are trained, reducing the number of trainable parameters from $d_{in}d_{out}$ to $r(d_{in}+d_{out})$.
+In this formulation, the original weights $\mathbf{W}_{0}$ remain fixed, and only $\mathbf{B}$ and $\mathbf{A}$ are trained, reducing the number of trainable parameters from $d_{\text{in}}d_{\text{out}}$ to $r(d_{\text{in}}+d_{\text{out}})$.
+While parameter-efficient, standard LoRA applies the rank-$r$ update densely to every input.
+This would result in updates for a new task that inevitably perturbs the subspaces used by prior tasks, leading to interference.
 Mixture‐of‐Experts LoRA.
-Building on the Mixture‐of‐Experts (MoE) paradigm, a common generic framework of MoE‐LoRA (Yu et al., 2024; Wang et al., 2024) treats each LoRA as an independent expert.
+Building on the Mixture‐of‐Experts (MoE) paradigm, a common generic framework of MoE‐LoRA [96, 78] treats each LoRA as an independent expert.
 Suppose after $T$ tasks, we have $T$ LoRA experts $\{(\mathbf{A}^{1},\mathbf{B}^{1}),\dots,(\mathbf{A}^{T},\mathbf{B}^{T})\}$.
@@
 Each LoRA’s contribution is weighted by the learnable router, enabling the model to dynamically select and combine the most relevant low‐rank updates for each token.
-In practice, a Top-$k$ masking is typically applied to mixture weights to enforce sparsity, activating only the $k$ most relevant experts and controlling computational overhead.
-3.2 Model Weights and Updates as Linear Associative Memory Model
-Weight matrix as linear associative memory.
-Large pre‐trained models (PTMs) have been shown to inhabit a low “intrinsic‐dimension” manifold compared to their full parameter count (Li et al., 2018; Aghajanyan et al., 2020).
-Any weight matrix of a pre-trained model $\mathbf{W}_{0}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$ can be treated as a linear associative memory model with a key-value structure (Kohonen, 1972; Anderson, 1972; Bau et al., 2020; Meng et al., 2022).
-Concretely, for any matrix $\mathbf{W}\in\mathbb{R}^{d_{\mathrm{out}}\times d_{\mathrm{in}}}$ and input $\mathbf{x}\in\mathbb{R}^{d_{\mathrm{in}}}$, the product $\mathbf{W}\,\mathbf{x}$ implements a content‐addressable memory read over $m$ stored key–value pairs $\{(\mathbf{K}_{i,:},\mathbf{V}_{:,i})\}_{i=1}^{m}$, with $m$ potentially large and varied across each weight matrix.
-Here, ${\mathbf{K}}=\begin{bmatrix}{\mathbf{K}}_{1,:}^{\top}&\cdots&{\mathbf{K}}_{m,:}^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{m\times d_{\text{in}}}$ and $\mathbf{V}$$\begin{bmatrix}{\mathbf{V}}_{:,1}&\cdots&{\mathbf{V}}_{:,m}\end{bmatrix}\in\mathbb{R}^{d_{\text{out}}\times m}$ with $\mathbf{K}_{i,:}\in\mathbb{R}^{1\times d_{\text{in}}}$ is the $i$-th row of $\mathbf{K}$} and ${\mathbf{V}}_{:,i}\in\mathbb{R}^{d_{\text{out}}\times 1}$ is the $i$-th column of $\mathbf{V}$}.
-Given any input $\mathbf{x}$, the output $\mathbf{y}\in\mathbb{R}^{d_{\text{out}}}$ is computed as
-$\mathbf{y}=\mathbf{W}\mathbf{x}\approx\sum\nolimits_{i=1}^{m}\mathbf{V}_{:,i}(\mathbf{K}_{i,:}\mathbf{x}).$ (2)
-In this formulation, the forward pass through the weight matrix with the input decomposes into two steps: (1) computing response scores or assignment weights via the key vectors $\mathbf{K}_{i,:}$, and (2) aggregating the corresponding value vectors $\mathbf{V}_{:,i}$ according to those weights.
-Figure 2: Overview of MoRA.
+In practice, a top-k masking is typically applied to mixture weights to enforce sparsity, activating only the $k$ most relevant experts.
+However, this design remains coarse-grained.
+Each expert is a dense rank-$r$ block, and the external router $R(\mathbf{x})$ must learn to map inputs to these overlapping experts.
+As the number of tasks grows, this external mapping becomes ambiguous, leading to routing collapse and forgetting.
+Figure 2: Overview of MoRAM.
 For each new task, we freeze the ranks learned on previous tasks and introduce $r$ new ranks of updates.
 Our sparse self-activated mixture‐of‐ranks framework jointly considers all old and new ranks, adaptively inferring a sparse mixture weight for each rank.
-Panels (a,c) illustrate MoRA conceptually and (b,d) detail its computation for tasks $t$ and $t+1$, respectively.
-Parameter update as modifications to the linear associative memory.
-Building on the associative-memory view of weights, we can likewise interpret weight updates during training as key–value insertions and modifications into the existing memory.
-During training, for an arbitrary update $\Delta\mathbf{W}$ on a given $\mathbf{W}_{0}$ as the starting point, the weight update process can be represented as
-$\mathbf{W}_{\text{new}}=\mathbf{W}_{0}+\Delta\mathbf{W},$ (3)
-where $\mathbf{W}_{0}$ can be the pre-trained model weights.
-We can also formulate the weight updates $\Delta\mathbf{W}$ with the key-value memory model $\{\Delta\mathbf{K},\Delta\mathbf{V}\}$ as an update to the original memory model.
-With the operation on an input $\mathbf{x}$ via
-$\Delta\mathbf{W}\mathbf{x}=\Delta\mathbf{V}\Delta\mathbf{K}\mathbf{x},$ (4)
-the updates on weights can be seen as a retrievable key-value system for the updating of $\mathbf{W}\mathbf{x}$.
-LoRA as low-rank key–value memory updates.
-LoRA (Hu et al., 2021) performs parameter‐efficient fine‐tuning by adding a rank‐$r$ update $\Delta\mathbf{W}=\mathbf{B}\mathbf{A}$ to the pre‐trained weight matrix $\mathbf{W}_{0}$, where $\mathbf{B}\in\mathbb{R}^{d_{\text{out}}\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times d_{in}}$.
-With the model update based on LoRA, given an input $\mathbf{x}$, the computation of updated weight becomes
-$\mathbf{y}_{\text{new}}=\mathbf{W}_{\text{new}}\mathbf{x}=\mathbf{W}_{0}\mathbf{x}+\Delta\mathbf{W}\mathbf{x}=\mathbf{W}_{0}\mathbf{x}+\mathbf{B}\mathbf{A}\mathbf{x}.$ (5)
-Decomposing the low-rank matrices into a collection of its $r$ rank-one components gives
-$\Delta\mathbf{W}\mathbf{x}=\mathbf{B}\mathbf{A}\mathbf{x}=\sum\nolimits_{i=1}^{r}\mathbf{B}_{:,i}(\mathbf{A}_{i,:}\mathbf{x}),$ (6)
-with $\mathbf{A}=\begin{bmatrix}\mathbf{A}_{1,:}^{\top}&\cdots&\mathbf{A}_{r,:}^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{r\times d_{\text{in}}}$ where $\mathbf{A}_{i,:}$ is the $i$-th row of $\mathbf{A}$, and $\mathbf{B}=\begin{bmatrix}\mathbf{B}_{:,1}&\cdots&\mathbf{B}_{:,r}\end{bmatrix}\in\mathbb{R}^{d_{\text{out}}\times r}$ where $\mathbf{B}_{:,i}$ is the $i$-th column of $\mathbf{B}$.
-Considering the connection between the formulation of LoRA Eq.
-(1) and associative‐memory view (Anderson, 1972; Meng et al., 2022) of weight update Eq.
-(4), we can intuitively interpret LoRA as obtaining a small key–value store in a low‐rank subspace, where the rows of $\mathbf{A}$ serve as keys and the columns of $\mathbf{B}$ serve as their corresponding values.
-After training the LoRA module on a base weight matrix $\mathbf{W}_{0}$, the modified forward pass in Eq.
-(5) with Eq.
-(6) can be seen as: first, the original pre‐trained computation $\mathbf{W}_{0}\,\mathbf{x}$ (Eq.
-(2)) is performed; then, each learned key $\mathbf{A}_{i,:}$ measures its relevance to the input $\mathbf{x}$ via the product $\mathbf{A}_{i,:}\mathbf{x}$, and each learned value $\mathbf{B}_{:,i}$ is scaled by this relevance score and aggregated, in a same way as shown in Eq.
-(2).
-In this way, LoRA effectively inserts or updates a small set of key–value memories that augment the pre‐trained model’s capabilities.
-3.3 Mixture of Ranks Adaptation
-In the key–value associative‐memory view of LoRA (Eq.
-(5) and Eq.
-(6)), a standard LoRA or its variants (Hu et al., 2021; Meng et al., 2024; Zhang et al., 2024a) apply a dense update: every rank’s subspace contributes to $\Delta\mathbf{W}\,\mathbf{x}$ for any input $\mathbf{x}$.
-Even in MoE-LoRA, where each LoRA module is treated as a separate expert, all key–value pairs within a LoRA module are typically used densely as a whole (either fully activated or not), which creates redundancy across experts and underutilizes the expressive potential of the key vectors $\mathbf{A}$ to indicate input relevance.
-Rather than by default activating all key–value memories for every input, many of which may encode patterns irrelevant to the current token, we introduce a gated, sparse mixture-of-ranks update that selectively activates only the most relevant rank-one key–value pairs.
-Concretely, we associate each rank-one key–value pair with a scalar weight ${\textnormal{w}}_{i}\in\mathbb{R}$.
-For an input token $\mathbf{x}\in\mathbb{R}^{d_{\mathrm{in}}}$ from task $t$, with $r_{t}$ total ranks, the update becomes
+Panels (a,c) illustrate MoRAM conceptually and (b,d) detail its computation for tasks $t$ and $t+1$, respectively.
+3.2 Weights as Linear Associative Memory
+To resolve the limitations of coarse granularity and routing ambiguity inherent in dense rank-$r$ updates, we depart from the conventional view of LoRA as monolithic blocks, instead reconceptualizing both the weight matrix and its updates as a Linear Associative Memory.
+Definition 1 (Weight Matrix as Linear Associative Memory).
+Consider a weight matrix $\mathbf{W}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$ of rank $m$.
+Through the lens of linear associative memory [37, 5], a weight matrix $\mathbf{W}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$ of rank $m$ acts as a retrieval system composed of $m$ atomic key-value pairs $\{(\mathbf{k}_{i},\mathbf{v}_{i})\}_{i=1}^{m}$, where $\mathbf{k}_{i}\in\mathbb{R}^{d_{\text{in}}}$ and $\mathbf{v}_{i}\in\mathbb{R}^{d_{\text{out}}}$ such that $\mathbf{W}\approx\sum_{i=1}^{m}\mathbf{v}_{i}\mathbf{k}_{i}^{\top}$.
+For an input hidden state for a token $\mathbf{x}\in\mathbb{R}^{d_{\text{in}}}$, the matrix-vector product effectively performs a content-addressable read operation:
+$\mathbf{y}=\mathbf{W}\mathbf{x}\approx\sum\nolimits_{i=1}^{m}\mathbf{v}_{i}(\mathbf{k}_{i}^{\top}\mathbf{x}),$ (2)
+where the inner product $(\mathbf{k}_{i}^{\top}\mathbf{x})$ computes the relevance (or activation strength) of the $i$-th memory slot to the input, which weights the retrieval of the value vector $\mathbf{v}_{i}$.
+Remark 2 (Key-Value Memory of Weight Matrix).
+It is crucial to distinguish Definition 1 from the key-value mechanism in Self-Attention [73].
+In attention, keys and values are dynamic projections of the input sequence generated at runtime.
+In contrast, under the linear associative memory view, the keys $\mathbf{k}_{i}$ and values $\mathbf{v}_{i}$ are static parameters intrinsic to the weight matrix, representing knowledge patterns acquired during pre-training.
+3.3 Proposed MoRAM
+3.3.1 Mixture of Rank-1 Memory Experts
+Fine-grained rank-1 memory augmentation.
+Grounded by the associative memory view (Definition 1), we reconceptualize the fine-tuning process.
+Instead of formulating the update $\Delta\mathbf{W}$ as a single-unit rank-$r$ matrix, we treat it as a collection of retrievable rank-1 memory augmentations.
+Formally, we define the update as the aggregation of $r$ atomic rank-1 key-value pair updates:
+$\Delta\mathbf{W}\mathbf{x}=\sum\nolimits_{i=1}^{r}\underbrace{\mathbf{B}_{:,i}}_{\text{Value }\mathbf{v}_{i}}(\underbrace{\mathbf{A}_{i,:}}_{\text{Key }\mathbf{k}_{i}^{\top}}\mathbf{x}),$ (3)
+where the row vector $\mathbf{A}_{i,:}$ acts as the Key determining the relevance of the atom to the input, and the column vector $\mathbf{B}_{:,i}$ acts as the Value storing the retrieved knowledge.
+Crucially, this shifts the paradigm from matrix adaptation to memory expansion: the update is no longer a rigid block, but a flexible set of fine-grained atomic memories.
+Despite this granular potential, standard LoRA and its variants [33, 56, 98] apply updates densely: every rank contributes to every input.
+Even in MoE-LoRA, while experts are separated at rank-$r$ adapter level, the constituent rank-1 memories within each expert remain entangled: (1) Interference: Irrelevant knowledge subspaces within a chosen expert are forced to activate for mismatched inputs.
+(2) Routing collapse: It overlooks the intrinsic capacity of the key vectors $\mathbf{A}$ to act as content retrieval keys, instead relying on indiscriminate dense activation or redundant external routers.
+MoRAM formulation.
+To address this, we propose MoRAM (Mixture of Rank-1 Associative Memory).
+We dismantle the rigid adapter structure, redefining the model’s adaptation parameters as a dynamic collection of atomic memory experts $\mathcal{M}_{t}=\{(\mathbf{B}_{:,i},\mathbf{A}_{i,:})\}_{i=1}^{r_{t}}$, where $r_{t}$ denotes the total accumulated rank-1 pairs avaliable at task $t$.
+For an input $\mathbf{x}$ during task $t$ (with accumulated memory atoms $r_{t}$), the effective update is defined as a sparse, input-dependent mixture:
 $\Delta\mathbf{W}^{t}=\sum\nolimits_{i=1}^{r_{t}}{\textnormal{w}}_{i}\mathbf{B}_{:,i}\mathbf{A}_{i,:},$ (4)
-where ${\textnormal{w}}_{i}$ modulates whether and how strongly each rank $i$ contributes to the updated weight matrix.
-3.4 Self-Activated Adaptive Sparse Mixture of Ranks
-A key challenge for MoE-style CL is routing ambiguity, leading to catastrophic forgetting.
-Conventional routers assign examples to coarse-grained experts, but adding new experts often makes routing suboptimal, causing over-activation or underuse, i.e., the three challenges discussed in Sec.
-1.
-To avoid these pitfalls, we treat each rank-one LoRA (or rank) as an expert based on the rank-level mixture model (Eq.
-(7)), enabling sparse activation of fine-grained experts for specific tasks.
-Thus, CL with incrementally added rank-one experts becomes a process of accumulating atomic key–value knowledge that can be reused in future tasks.
-Unlike conventional MoE-LoRA (Yang et al., 2024a; Wang et al., 2024; Yu et al., 2024), we use the rank-one $\mathbf{A}_{i}$’s (keys) as routers to their corresponding $\mathbf{B}_{i}$’s (values) through self-activation, with each $\mathbf{A}_{i}$$\mathbf{B}_{i}$ pair functioning as a router-expert pair.
-Mixture of self-activated rank-one experts.
-Rather than relying on an external router, our framework derives mixture weights for each rank directly from that rank’s activation.
-Recall from Eq.
-(6) that each rank-one update is parameterized by a key–value memory pair $(\mathbf{A}_{i,:},\mathbf{B}_{:,i})_{i=1}^{r_{t}}$; the key $\mathbf{A}_{i,:}$ therefore provides a natural, input-dependent relevance signal.
-Specifically, given an input token $x\in\mathbb{R}^{d_{\mathrm{in}}}$ from task $t$, and treating $\mathbf{A}$ as the set of keys, we compute a raw score for all $r_{t}$ ranks accumulated up to task $t$:
-${\textnormal{s}}_{i}\;=\;\frac{\mathbf{A}_{i}x}{\sqrt{\sum_{j=1}^{r_{t}}\|\mathbf{A}_{j}x\|_{2}^{2}}}\quad\in\mathbb{R},$ (8)
-where the $\ell_{2}$-normalization across all $r_{t}$ ranks ensures numerical stability and compensates for activation-scale variations between layers and pre-trained weights.
-Using raw scores alone matches or surpasses a separately learned router (with 1.5 times parameters) without self-activation (Table 4), showing that each rank can infer its own relevance from activation.
-Self-activated rank selection.
-To ensure a stable and gated, sparse activation pattern, we impose a fixed sparse rank‐activation budget via top-k masking on the $\ell_{2}$–normalized scores ${\textnormal{s}}_{i}:=\mathrm{TopK}({\textnormal{s}},\,k),$ where
+where ${\textnormal{w}}_{i}\in\mathbb{R}$ represents the computed retrieval confidence of the $i$-th memory atom on each input token hidden state $\mathbf{x}$}.
+This formulation allows the model to freeze specific “atomic” memories (preserving old tasks) while inserting new ones on demand of input, or to jointly activate a combination of old and new memory slots to handle shared concepts.
+3.3.2 Self-Activation for MoRAM Routing
+Relying on the key-value form of the rank-1 adapter, we propose a self-activation mechanism for routing the memory experts.
+We derive the mixing weights directly from the intrinsic activation of each atomic memory, leveraging the key vectors $\mathbf{A}_{i,:}$ for memory retrieval.
+Unlike standard MoE approaches that rely on additional routing networks, our self-activation routing performs routing as a content-addressable retrieval, reducing the possibility of forgetting caused by an additional router.
+Self-activated relevance scoring.
+By eliminating the external router, MoRAM avoids the optimization instability and forgetting associated with auxiliary networks.
+Instead, we derive mixing weights directly from the intrinsic alignment between the input and the memory keys.
+Formally, for an input token $\mathbf{x}\in\mathbb{R}^{d_{\text{in}}}$ and the set of accumulated memory atoms up to task $t$, we compute the raw relevance score ${\textnormal{s}}_{i}$ for the $i$-th atom as:
+${\textnormal{s}}_{i}\;=\;\frac{\mathbf{A}_{i,:}\mathbf{x}}{\sqrt{\sum_{j=1}^{r_{t}}(\mathbf{A}_{j,:}\mathbf{x})^{2}}},$ (5)
+where the numerator represents the relevance to the memory key $\mathbf{A}_{i,:}$, the denominator performs an $\ell_{2}$-normalization across the entire memory ensemble for numerical stability.
+Empirically, we find that this intrinsic scoring mechanism matches or exceeds the performance of external routers (see Table 4), confirming that the memory keys alone contain sufficient information to determine their own utility.
+3.3.3 Sparse Expert Routing and Mixture
+While Eq.
+(5) measures relevance, naive dense activation leads to low specialization and induces interference and computational overhead.
+We therefore employ sparse routing to enhance fine-grained specialization of the fine-grained experts and routers.
+Sparse rank selection.
+To prevent interference between tasks and ensure a budgeted computational cost, we enforce a sparsity constraint via top-k masking.
+Given the raw relevance scores s, we retain only the $k$ entries with the highest activation and mask the rest:
 […]
-ensuring that at most $k$ out of the total $r_{t}$ ranks are eligible to be activated for task $t$.
-This encourages rank specialization: only a small set of the most relevant ranks are emphasized and trained to capture each kind of input-specific dynamics, and it prevents tiny, noisy activations from affecting gradients and learning dynamics.
-Sharpness control on rank activations. To further encourage rank specialization and concentrate the update on the most relevant ranks, we temperature-scale and normalize the scores:
-${\textnormal{w}}_{i}\;=\;\mathrm{softmax}\!\bigl(\tfrac{s}{\tau_{\text{MoRA}}}\bigr)_{i},$ (10)
-where the softmax is taken over the vector ${\textnormal{s}}=[{\textnormal{s}}_{1},\dots,{\textnormal{s}}_{r_{t}}]$ and $\tau_{\text{MoRA}}$ controls the distribution’s sharpness.
-A lower $\tau_{\text{MoRA}}$ concentrates mass on a few ranks (specialists), sharpens the distribution and routes larger learning signals to top ranks, which accelerates specialization and reduces noisy updates to low-relevance ranks, while a higher $\tau_{\text{MoRA}}$ spreads mass (shared representations).
-Threshold-based test-time rank selection.
-To further suppress low-signal ranks at inference, we optionally apply a threshold $\delta$ to the $\ell_{2}$–normalized raw scores and zero out gates below it.
-We do not hard-threshold during training such that gradients can still flow to near-threshold ranks and allow exploration and adaptation; thresholding only at test time reduces noise and runtime cost:
+This masking ensures that for any given input, at most $k$ out of the total $r_{t}$ accumulated memory atoms are eligible for activation.
+This encourages rank specialization: only a small set of the most relevant ranks are emphasized and trained to capture each kind of input-specific dynamics, and it prevents tiny, noisy activations from spuriously triggering unrelated memory atoms (e.g., those frozen from prior tasks).
+Sharpness enhancement.
+To further encourage rank specialization and concentrate the update on the most relevant ranks, we apply temperature-scaled softmax normalization to enhance the sharpness of the mixture weights:
+${\textnormal{w}}_{i}\;=\;\text{softmax}\Bigl(\frac{\mathrm{TopK}({\textnormal{s}},k)}{\tau_{\text{MoRAM}}}\Bigr)_{i},$ (7)
+where $\tau_{\text{MoRAM}}$ is a scalar temperature hyperparameter.
+In the forward pass, a lower $\tau_{\text{MoRAM}}$ acts as a contrast enhancer, concentrating probability mass on the definitive specialist ranks.
+In the backward pass, it functions as a gradient router: by sharpening the distribution, stronger learning signals are directed exclusively to the winning ranks.
+This effectively isolates them from irrelevant updates and accelerates specialization, ensuring that memory atoms only evolve when they are truly relevant to the current data distribution.
+Threshold-based expert selection.
+To maximize retrieval precision during inference, we apply an additional threshold-based expert selection, via a relevance threshold $\delta$ to the normalized scores, enabling adaptively pruning memory atoms with weak activation signals.
+Since the number of useful experts may differ depending on layers and input, at inference time, we apply this threshold-based filter to further adaptively select experts from the top-k experts, while reducing both computational overhead and noise:
 ${\textnormal{w}}_{i}:=\mathbbm{1}\{{\textnormal{s}}_{i}\geq\delta\}\odot{\textnormal{w}}_{i}.$ (8)
-This yields a sparse, input-dependent mixture of only the most significant rank-one experts.
-Our fine-grained mixture framework offers two key advantages: (1) It reduces redundancy and enables shared-knowledge reuse: ranks that capture common, cross-task features (such as generic semantics) can be reactivated across multiple tasks, preventing those patterns from being relearned redundantly or interfered; (2) It minimizes routing ambiguity and prevents forgetting: each rank can specialize on more nuanced patterns, yielding an optimized, rank‐wise mixture that integrates new knowledge without disrupting existing representations or hindering future updates.
-Incremental mixture of ranks.
-Unlike methods that add a fixed number of LoRA experts per task (Yu et al., 2024), MoRA incrementally adds $r$ new ranks for each task, freezes all prior ranks, and adaptively selects a sparse subset of the most relevant ranks during both training and inference.
-This design seamlessly incorporates new task-specific knowledge into freshly initialized ranks while preserving earlier ranks unchanged to prevent forgetting.
-Moreover, it facilitates common knowledge reuse by allowing ranks that encode common patterns to be activated in later tasks.
+This operation yields a highly sparse, input-dependent set comprising only the most significant memory experts.
+3.3.4 Continual Learning with MoRAM
+Incremental memory expansion.
+Unlike methods that add fixed-size adapter blocks [96], MoRAM expands memory “little by little.” For each new task, we introduce a set of $r$ new atomic rank-1 pairs (Keys $\mathbf{A}$ and Values $\mathbf{B}$), freeze all prior atoms, and let the self-activation mechanism jointly route across the union of all old and new memories.
+This seamlessly integrates new knowledge while preserving the reusability of prior tasks.
+To enhance efficiency and enable sub-linear expansion rate, MoRAM supports low-utility atoms to be pruned after training with minimal performance degradation (more in Appendix A.10).
 Training objectives.
 In our experiments, we only use the model’s standard training objective, without any extra regularization or load‐balancing constraints.
-In practice, we empirically observe that the conventional load-balancing loss widely employed in the Mixture-of-Experts literature (Shazeer et al., 2017; Lepikhin et al., 2020; Fedus et al., 2022; Dai et al., 2024) is not necessary, thanks to our fine-grained self-activated gating design.
-Imposing a strict load-balancing constraint might hinder the learning of the key vectors (i.e., matrix $\mathbf{A}$), since each rank should remain free to capture its own characteristic input distribution.
-Fig.
-3 illustrates how each rank learns distinct characteristics.
+Benefiting from the fine-grained modeling with rank-1 experts, the experts tend to be special and the routing tends to be balanced and stable naturally.
+Even without additional regularization, the specialized experts and router reduce router degradation and forgetting, while we believe further regularization might be more helpful.
+In practice, we empirically observe that conventional load-balancing regularization [68, 40, 24, 15] for MoE may not be necessary for the experimental setup.
+And imposing regularization on the self-activation routing improperly tends to hinder the learning of the key vectors, since the key vectors for routing are also crucial for information representation.
 Table 1: Comparisons on X-TAIL for each domain in terms of “Transfer”, “Average”, and “Last” scores (%). The best and the second best results are highlighted in red and blue, respectively.
-Method
-Aircraft Caltech DTD EuroSAT Flowers Food MNIST OxPet Cars SUN397 Average
+    Method Aircraft Caltech DTD EuroSAT Flowers Food MNIST OxPet Cars SUN397 Average
 CLIP
@@
 Transfer
-   Zero-shot (Radford et al., 2021) 76.8 37.3 36.7 63.6 84.0 46.7 86.7 66.1 63.7 62.4
-   LwF (Li & Hoiem, 2017) 66.6 26.9 19.5 51.0 78.4 26.6 68.9 35.5 56.1 47.7
-   WiSE-FT (Wortsman et al., 2022) 70.1 31.9 25.3 56.3 79.8 29.9 74.9 45.6 56.8 52.3
-   iCaRL (Rebuffi et al., 2017) 71.7 35.0 43.0 63.4 86.9 43.9 87.8 63.7 60.0 61.7
-   ZSCL (Zheng et al., 2023) 73.3 32.6 36.8 62.1 83.8 42.1 83.6 56.5 60.2 59.0
-   MoE-Adapter (Yu et al., 2024) 71.0 34.9 19.2 63.0 86.6 20.0 87.2 63.7 58.6 56.0
-   RAIL-Primal (Xu et al., 2024) 76.8 37.3 36.7 63.6 84.0 46.7 86.7 66.1 63.7 62.4
-   CoDyRA (Lu et al., 2024) 74.3 36.8 44.2 69.9 83.5 42.8 88.9 64.6 63.4 63.2
-   MoRA 74.5 38.1 46.9 65.3 82.9 45.8 88.2 65.1 62.9 63.3
+   Zero-shot [63] 76.8 37.3 36.7 63.6 84.0 46.7 86.7 66.1 63.7 62.4
+   LwF [44] 66.6 26.9 19.5 51.0 78.4 26.6 68.9 35.5 56.1 47.7
+   WiSE-FT [86] 70.1 31.9 25.3 56.3 79.8 29.9 74.9 45.6 56.8 52.3
+   iCaRL [66] 71.7 35.0 43.0 63.4 86.9 43.9 87.8 63.7 60.0 61.7
+   ZSCL [105] 73.3 32.6 36.8 62.1 83.8 42.1 83.6 56.5 60.2 59.0
+   MoE-Adapter [96] 71.0 34.9 19.2 63.0 86.6 20.0 87.2 63.7 58.6 56.0
+   RAIL-Primal [91] 76.8 37.3 36.7 63.6 84.0 46.7 86.7 66.1 63.7 62.4
+   CoDyRA [50] 74.3 36.8 44.2 69.9 83.5 42.8 88.9 64.6 63.4 63.2
+   MoRAM 74.5 38.1 46.9 65.3 82.9 45.8 88.2 65.1 62.9 63.3
 Average
-   LwF (Li & Hoiem, 2017) 24.7 79.7 38.3 36.9 63.9 81.0 36.5 71.9 42.7 56.7 53.2
-   WiSE-FT (Wortsman et al., 2022) 27.1 76.5 40.9 31.3 68.7 81.6 31.4 74.7 51.7 58.4 54.2
-   iCaRL (Rebuffi et al., 2017) 25.4 72.1 37.5 51.6 65.1 87.1 59.1 88.0 63.7 60.1 61.0
-   ZSCL (Zheng et al., 2023) 36.0 75.0 40.7 40.5 71.0 85.3 46.3 83.3 60.7 61.5 60.0
-   MoE-Adapter (Yu et al., 2024) 43.6 77.9 52.1 34.7 75.9 86.3 45.2 87.4 66.6 60.2 63.0
-   RAIL-Primal (Xu et al., 2024) 42.4 89.8 55.7 68.5 84.0 83.3 65.3 85.8 67.9 64.5 70.7
-   CoDyRA (Lu et al., 2024) 41.4 81.0 58.7 77.8 83.4 84.6 64.5 90.4 67.2 64.4 71.3
-   MoRA 44.1 81.6 64.6 79.6 83.9 84.4 66.5 89.7 68.4 64.1 72.7
+   LwF [44] 24.7 79.7 38.3 36.9 63.9 81.0 36.5 71.9 42.7 56.7 53.2
+   WiSE-FT [86] 27.1 76.5 40.9 31.3 68.7 81.6 31.4 74.7 51.7 58.4 54.2
+   iCaRL [66] 25.4 72.1 37.5 51.6 65.1 87.1 59.1 88.0 63.7 60.1 61.0
+   ZSCL [105] 36.0 75.0 40.7 40.5 71.0 85.3 46.3 83.3 60.7 61.5 60.0
+   MoE-Adapter [96] 43.6 77.9 52.1 34.7 75.9 86.3 45.2 87.4 66.6 60.2 63.0
+   RAIL-Primal [91] 42.4 89.8 55.7 68.5 84.0 83.3 65.3 85.8 67.9 64.5 70.7
+   CoDyRA [50] 41.4 81.0 58.7 77.8 83.4 84.6 64.5 90.4 67.2 64.4 71.3
+   MoRAM 44.1 81.6 64.6 79.6 83.9 84.4 66.5 89.7 68.4 64.1 72.7
 Last
-   LwF (Li & Hoiem, 2017) 25.5 72.1 38.9 55.4 65.5 87.3 81.9 88.6 63.6 61.5 64.0
-   WiSE-FT (Wortsman et al., 2022) 21.8 76.8 42.9 20.8 77.5 84.9 30.7 76.6 75.8 72.5 58.0
-   iCaRL (Rebuffi et al., 2017) 25.5 72.1 38.9 55.4 65.5 87.3 81.9 88.6 63.6 61.5 64.0
-   ZSCL (Zheng et al., 2023) 33.1 75.3 43.5 35.2 74.6 87.4 50.4 84.2 77.3 73.4 63.4
-   MoE-Adapter (Yu et al., 2024) 43.2 78.7 57.6 32.8 79.4 86.0 86.7 87.8 78.2 74.2 70.5
-   RAIL-Primal (Xu et al., 2024) 41.7 94.0 66.0 86.4 97.2 82.4 93.1 83.6 75.0 71.3 79.1
-   CoDyRA (Lu et al., 2024) 37.7 81.5 65.1 89.9 91.4 85.5 96.8 93.3 77.3 73.5 79.2
-   MoRA 37.7 81.5 70.7 92.4 95.0 86.0 97.6 92.6 81.0 74.7 80.9
+   LwF [44] 25.5 72.1 38.9 55.4 65.5 87.3 81.9 88.6 63.6 61.5 64.0
+   WiSE-FT [86] 21.8 76.8 42.9 20.8 77.5 84.9 30.7 76.6 75.8 72.5 58.0
+   iCaRL [66] 25.5 72.1 38.9 55.4 65.5 87.3 81.9 88.6 63.6 61.5 64.0
+   ZSCL [105] 33.1 75.3 43.5 35.2 74.6 87.4 50.4 84.2 77.3 73.4 63.4
+   MoE-Adapter [96] 43.2 78.7 57.6 32.8 79.4 86.0 86.7 87.8 78.2 74.2 70.5
+   RAIL-Primal [91] 41.7 94.0 66.0 86.4 97.2 82.4 93.1 83.6 75.0 71.3 79.1
+   CoDyRA [50] 37.7 81.5 65.1 89.9 91.4 85.5 96.8 93.3 77.3 73.5 79.2
+   MoRAM 37.7 81.5 70.7 92.4 95.0 86.0 97.6 92.6 81.0 74.7 80.9
 4 Experiments
-In this paper, we conduct experiments across a diverse set of tasks, including continual learning for both vision-language CLIP models and LMs, and analyze catastrophic forgetting during fine‐tuning.
+In this paper, we conduct experiments across a diverse set of tasks, including continual learning for both vision-language CLIP models and LLMs, and analyze catastrophic forgetting during fine‐tuning.
 Detailed implementation settings and more experiment results are provided in Appendix A.1.
-CL of CLIP.
-We evaluate on two benchmarks using CLIP-ViT/B-16 (Radford et al., 2021; Ilharco et al., 2021) on MTIL (Zheng et al., 2023; Yu et al., 2024) and X-TAIL (Xu et al., 2024; Lu et al., 2024), with a range of image classification datasets.
-Following prior work, we report Transfer (zero-shot on unseen tasks), Last (retention on earlier tasks), and Average (mean accuracy across all datasets of all learning tasks) to assess continual learning performance.
-CL of LM.
-We follow prior works (Qin & Joty, 2021; Razdaibiedina et al., 2023; Wang et al., 2023; Qiao & Mahdavi, 2024) to continually fine-tune T5-large (Raffel et al., 2020) and LLaMA2-7B (Touvron et al., 2023) on five text-classification benchmarks, under three different task orderings.
-We report the average final accuracy across all tasks after completing the last task.
-Generalization and forgetting on unseen tasks.
-To assess effects on pre-trained general knowledge, we fine-tune Llama3.1-8B (Grattafiori et al., 2024) on the CodeAlpaca code-generation dataset (Chaudhary, 2023) and evaluate zero-shot in-domain performance on HumanEval (Chen et al., 2021), as well as out-of-domain accuracy on a broad selection of MMLU (Hendrycks et al., 2021) subjects.
-Table 2: Zero-shot in-domain performance on HumanEval (Pass@1) for the code generation domain and out-of-domain accuracy on selected MMLU subjects (formal logic, philosophy, world religions, economics, public relations, STEM, physics, machine learning) after fine-tuning Llama-3.1-8B on CodeAlpaca.
-The last two columns report trainable parameters (for MoRA: added / activated).
-Method
-HumanEval
-(Pass@1)
-Out-of-Domain (Acc.)
-Params
+4.1 Experimental Results
+Continual Learning of CLIP.
+We evaluate X-TAIL performance in Table 1, reporting Transfer, Last, and Average accuracy (MTIL results in Appendix Table 12).
+While prior methods [96, 91] are constrained by the base model’s zero-shot ceiling, MoRAM follows [105, 50] to continuously adapt, surpassing this limit and achieving superior Last and Average scores.
+Critically, MoRAM enhances representation without the external domain-IDs or feature banks required by [96, 91], and improves upon the fixed weighting in [50] by utilizing dynamic, input-dependent gating for finer expert specialization.
+Continual Learning of LLMs.
+In Table 2, we evaluate MoRAM on LLaMA [23] and Gemma [71] using the TRACE benchmark [81].
+Traditional regularization [36], rehearsal [48], and prompt-based [85] methods struggle to scale, often underperforming simple In-Context Learning.
+Among LoRA variants, naive SeqLoRA suffers significant forgetting, while recent approaches like O-LoRA [80], HiDeLoRA [79], and TreeLoRA [60] rely on rigid orthogonality or complex structural expansions.
+In contrast, MoRAM performs robustly on LLMs, achieving results competitive with existing methods.
+We provide additional evaluations on language classification tasks in Tables 8 and 9 of the Appendix.
+Table 2: Comparison with a broad range of CL methods on the TRACE benchmark.
+We report Overall Performance (OP (%) $\uparrow$) and Backward Transfer (BWT (%) $\downarrow$).
+Results are averaged over three runs with standard deviations.
+The best results are highlighted in bold.
+FIX(ICL) SeqLoRA OGD GEM EWC L2P DualPrompt HiDeLoRA O-LoRA TreeLoRA MoRAM
+meta-llama / LLaMA-2-7B-Chat
+OP 38.94 $\pm$ 0.3 34.3 $\pm$ 1.2 42.09 $\pm$ 1.6 40.08 $\pm$ 1.6 42.36 $\pm$ 1.2 36.23 $\pm$ 0.8 37.69 $\pm$ 1.2 41.60 $\pm$ 0.8 42.78 $\pm$ 0.8 43.52 $\pm$ 1.0 44.54 $\pm$ 0.9
+BWT 18.5 $\pm$ 0.8 8.06 $\pm$ 1.2 6.77 $\pm$ 1.2 5.97 $\pm$ 0.8 8.25 $\pm$ 0.8 8.03 $\pm$ 0.8 7.12 $\pm$ 0.4 7.16 $\pm$ 0.4 3.46 $\pm$ 0.4 1.37 $\pm$ 0.3
+google / Gemma-2B-it
+OP 32.3 $\pm$ 0.2 31.89 $\pm$ 0.8 32.85 $\pm$ 1.4 26.48 $\pm$ 1.5 28.35 $\pm$ 1.6 31.14 $\pm$ 1.2 32.42 $\pm$ 1.0 33.25 $\pm$ 0.9 33.73 $\pm$ 0.8 33.41 $\pm$ 0.9 36.27 $\pm$ 0.7
+BWT 15.28 $\pm$ 0.4 12.27 $\pm$ 0.9 18.25 $\pm$ 0.9 16.96 $\pm$ 1.2 15.77 $\pm$ 0.7 14.25 $\pm$ 0.5 13.66 $\pm$ 0.5 12.36 $\pm$ 0.4 8.50 $\pm$ 0.5 2.74 $\pm$ 0.4
+meta-llama / LLaMA-3-1B-Instruct
+OP 31.16 $\pm$ 0.4 29.73 $\pm$ 1.6 30.12 $\pm$ 2.0 32.19 $\pm$ 2.0 31.96 $\pm$ 1.6 29.38 $\pm$ 1.2 30.76 $\pm$ 1.2 33.73 $\pm$ 1.2 32.94 $\pm$ 0.8 36.14 $\pm$ 0.7 37.77 $\pm$ 0.8
+BWT 17.03 $\pm$ 1.2 15.2 $\pm$ 1.6 10.74 $\pm$ 1.6 11.62 $\pm$ 1.2 13.57 $\pm$ 0.8 11.34 $\pm$ 0.8 12.36 $\pm$ 0.8 12.89 $\pm$ 1.2 7.36 $\pm$ 0.8 3.12 $\pm$ 0.8
+Table 3: Standard fine-tuning of Llama-3.1-8B on CodeAlpaca.
+We report zero-shot in-domain performance on HumanEval (Pass@1) for the code generation and out-of-domain accuracy on selected MMLU subjects (formal logic, philosophy, world religions, economics, public relations, STEM, physics, machine learning).
+The last two columns report trainable parameters (for MoRAM: added / activated).
 […]
@@
 LoRA ($r=32$) 47.56 42.85 69.45 81.87 72.27 66.36 55.44 45.10 39.29 63.59 83.9 1.03%
-MoRA 47.56 48.41 70.09 82.46 73.53 68.18 55.53 46.08 41.96 63.70 41.9/26.2 0.52%/0.32%
-4.1 Continual Learning of CLIP
-Table 1 reports X-TAIL results (MTIL results in Appendix Table 11).
-Prior methods (Yu et al., 2024; Xu et al., 2024) preserve pre-trained capabilities via zero-shot predictions and are therefore capped by the base model’s zero-shot performance.
-Following recent practice (Zheng et al., 2023; Lu et al., 2024), MoRA continuously trains the pre-trained model so it can surpass that zero-shot ceiling and better generalize to unseen domains, while still preserving existing capabilities.
-Across both Last and Average metrics, MoRA substantially outperforms prior works.
-Unlike methods that rely on domain-ID prediction (Yu et al., 2024), high-dimensional projection layers or feature-banks (Xu et al., 2024), MoRA improves the pre-trained representation without external mechanisms; compared with (Lu et al., 2024), which learns fixed weights per rank-$r$ update, MoRA uses input-dependent gates on individual rank-one updates for finer expert specialization and dynamic utilization.
-Table 3: Summary of results on standard CL benchmarks with T5-large. We report averaged accuracy after training on the last task across three task orderings.
-Standard CL Benchmark
-Method Order-1 Order-2 Order-3 Avg.
-MTL 80.0
-SeqFT 18.9 24.9 41.7 28.5
-SeqLoRA 44.6 32.7 53.7 43.7
-IncLoRA 66.0 64.9 68.3 66.4
-Replay 55.2 56.9 61.3 57.8
-EWC 48.7 47.7 54.5 50.3
-LwF 54.4 53.1 49.6 52.3
-L2P 60.3 61.7 61.1 60.7
-LFPT5 67.6 72.6 77.9 72.7
-InfLoRA 75.2 75.4 75.8 75.5
-O-LoRA 75.4 75.7 76.3 75.8
-LB-CL 76.9 76.5 76.8 76.7
-MoRA 77.4 77.5 77.9 77.6
-4.2 Continual Learning of LM
-Table 3 reports results across three task orderings: MoRA consistently outperforms prior methods and closely approaches the multi-task learning (MTL) upper bound.
-Unlike O-LoRA (Wang et al., 2023) and LB-CL (Qiao & Mahdavi, 2024), which rely on orthogonality constraints or gradient projections between per-task LoRA adapters (potentially limiting adapter capacity), MoRA needs no extra regularization.
-By decomposing each rank-r update into rank-one components and applying self-activated, sparse gating, MoRA lets each component specialize on its own input distribution, reducing interference and more effectively capturing diverse patterns.
-We further evaluate MoRA on LLaMA2-7B (Touvron et al., 2023) under the same continual-learning setup (Table 9 in the Appendix), MoRA also outperforms O-LoRA by 2.3% averaged over 3 task orders.
-4.3 Forgetting and Generalization during Fine‐Tuning
+MoRAM 47.56 48.41 70.09 82.46 73.53 68.18 55.53 46.08 41.96 63.70 41.9/26.2 0.52%/0.32%
+Forgetting and generalization after standard fine‐tuning.
 Table 3 examines how standard fine-tuning affects both in-domain performance and out-of-domain generalization.
-We observe two key phenomena.
-(1) Forgetting of unrelated knowledge: conventional LoRA fine-tuning causes drops on topics that are unrelated to the fine-tuning data (e.g., Philosophy, World Religions, Public Relations), indicating that pre-trained general knowledge is overwritten during adaptation.
-(2) Improved generalization: fine-tuning on code data increases accuracy on logically and quantitatively oriented tasks (Formal Logic, Economics, STEM, Physics, Machine Learning), showing that representations learned from code transfer to related symbolic domains.
-In contrast, MoRA both preserves unrelated pre-trained knowledge and improves in- and out-of-domain generalization.
-Decomposing each low-rank update into specialized rank-one components and activating only the most relevant subset prevents catastrophic overwriting of prior knowledge while allowing targeted adaptation.
-At the same time, MoRA achieves strong code-generation performance using roughly one-third of the activated parameters required by a rank-32 LoRA and substantially improves out-of-domain accuracy.
-4.4 Visualizations of Fine-grained Rank Activations
+(1) Forgetting: Dense LoRA updates overwrite pre-trained knowledge, degrading performance on semantically distant topics (e.g., Religions).
+(2) Positive transfer: Conversely, code adaptation improves logically related tasks (e.g., STEM, Logic), indicating valid transfer potential.
+In contrast, MoRAM leverages its atomic memory structure, treating updates as independent rank-1 memory atoms rather than entangled blocks, the model isolates new knowledge accumulation from existing representations.
+This granular independence prevents the catastrophic overwriting of unrelated concepts (e.g., Religions) while the content-addressable self-activation ensures that relevant atoms are reused for logical tasks (e.g., STEM, Logic).
+Consequently, MoRAM achieves superior out-of-domain accuracy using roughly one-third of the active parameters required by a standard rank-32 LoRA.
+[…]
+(a) Rank Activations of MoRAM on data from Task 1 after learning Task 1.
+[…]
+(b) Rank Activations of MoRAM on data from Task 1 after learning Task 2.
+[…]
+(c) Rank Activations of MoRAM on data from Task 2 after learning Task 2.
+Figure 3: Visualization of MoRAM rank activations during Task 1 and Task 2 training.
+Activations are extracted from the K projection in the attention module (layer 8) of the image encoder.
+Corresponding image patches are shown below each activation map, with regions relevant to each class marked by orange bounding boxes.
+Zoom in for details.
+More visualizations are in Figs.
+5 and 6 of the Appendix, demonstrating forgetting mitigation and knowledge reuse.
+Table 4: Routing Strategies.
+Routing Transfer Average Last
+MoE-LoRA (Coarse Baseline) 62.56 69.45 74.53
+External Router (Learned $W_{r}$) 60.09 65.97 69.76
+Self-Activated Retrieval 60.26 65.94 69.85
+    w/ Sparsity Constraint (Top-$k$) 60.69 66.52 70.62
+    w/ Temperature Scaling ($\tau_{\text{MoRAM}}$) 62.07 71.15 79.62
+    w/ Threshold-based expert selection ($\delta$) 60.78 66.83 71.08
+    MoRAM (Full) 63.30 72.70 80.90
+[…]
+(a) Acti. Budget
+(b) Temp. $\tau_{\text{MoRAM}}$
+(c) Thre. $\delta$
+Figure 3: Ablation on (a) rank activation budget, (b) temperature $\tau_{\text{MoRAM}}$, and (c) threshold $\delta$.
+4.2 Visualizations of Rank-1 Memory Activations
 Figure 3 shows rank activations recorded during continual learning.
-These visualizations illustrate two properties of MoRA: (1) individual ranks specialize on distinct input patterns, and (2) the model substantially reduces cross-task interference, thereby mitigating forgetting.
+These visualizations illustrate two properties of MoRAM: (1) individual ranks specialize on distinct input patterns, and (2) the model substantially reduces cross-task interference, thereby mitigating forgetting.
 Extended visualizations across more tasks and scenarios appear in Fig.
@@
 3(c) (e.g., leaves, shadows) yield a richer, more distributed pattern than the simple blue sky in Fig.
-3(a), highlighting MoRA’s capacity to model contextual complexity.
+3(a), highlighting MoRAM ’s capacity to model contextual complexity.
 Some ranks learned earlier are also reused on later tasks (Fig.
-5, Appendix), indicating transfer of shared semantics.
-Reduced cross-task interference and mitigated forgetting.
+6, Appendix), indicating transfer of shared semantics.
+Reduced interference and mitigated forgetting.
 Comparing the same input after Task 1 (Fig.
@@
 3(b)) shows almost identical activations (more in Fig.
-4, Appendix): Rank 0 still responds to airplane semantics and Rank 11 to blue-sky patches.
-This stability indicates that our self-activated, sparse mixture-of-ranks prevents later updates from overwriting earlier task representations, reducing inter-task interference and mitigating catastrophic forgetting.
-[…]
-(a) Rank Activations of MoRA on data from Task 1 after learning Task 1.
-[…]
-(b) Rank Activations of MoRA on data from Task 1 after learning Task 2.
-[…]
-(c) Rank Activations of MoRA on data from Task 2 after learning Task 2.
-Figure 3: Visualization of MoRA rank activations during Task 1 and Task 2 training.
-Activations are extracted from the K projection in the attention module (layer 8) of the image encoder.
-Corresponding image patches are shown below each activation map, with regions relevant to each class marked by orange bounding boxes.
-Zoom in for details.
-More visualizations are in Figs.
-4 and 5 of the Appendix, demonstrating forgetting mitigation and knowledge reuse.
+5, Appendix): Rank 0 still responds to airplane semantics and Rank 11 to blue-sky patches.
+This stability indicates that our self-activated, sparse mixture-of-ranks prevents later updates from overwriting earlier task representations, reducing interference and mitigating forgetting.
 4.3 Ablation Studies
-Routing/activation strategy.
-Table 4 compares different strategies for computing mixture weights, under a controlled setup with one LoRA per task, standard training loss, and no additional regularization: (1) MoE-LoRA trains a router to assign weights to each LoRA adapter; its coarse-grained design and lack of mechanisms that encourage expert specialization lead to greater interference and increased forgetting compared to MoRA.
-(2) Non-self-activation trains a finer-grained router for mixing each rank-one component.
-However, with a much larger number of experts (ranks), the router becomes increasingly prone to confusion, resulting in higher training complexity and degraded performance compared with MoE-LoRA.
-(3) Self-activation infers weights directly from softmax-normalized activations (Eq.
-8), matching or exceeding Non-self-activation and demonstrating that each rank can infer its own relevance without extra trainable router modules.
-(4) Top-k rank selection restricts updates to the k most relevant ranks, reducing interference and improving accuracy.
-(5) Sharpness control concentrates activation mass and substantially boosts specialization.
-(6) Test-time rank thresholding removes low-importance ranks at inference, yielding a slight gain by eliminating noisy contributions.
-Table 4: Routing/activation strategy
-Mixture Calculation Transfer Average Last
-MoE-LoRA 62.56 69.45 74.53
-Non-self-activation 60.09 65.97 69.76
-Self-activation 60.26 65.94 69.85
-    w/ top-k rank selection 60.69 66.52 70.62
-    w/ sharpness control with $\tau_{\text{MoRA}}$ 62.07 71.15 79.62
-    w/ test-time rank thresholding 60.78 66.83 71.08
-    MoRA 63.30 72.70 80.90
-[…]
-(a) Acti. Budget
-(b) Temp. $\tau_{\text{MoRA}}$
-(c) Thre. $\delta$
-Figure 3: Ablation on (a) rank activation budget, (b) temperature $\tau_{\text{MoRA}}$, and (c) threshold $\delta$.
+Ablation of memory retrieval strategies.
+Table 4 analyzes the impact of different retrieval mechanisms under a controlled setup.
+(1) MoE-LoRA (Coarse Baseline) employs a router to weight entire adapters.
+Its coarse granularity forces the simultaneous activation of conflicting subspaces, leading to significant interference and forgetting.
+(2) External Router applies a separately learned router to individual rank-1 atoms.
+However, decoupling the routing logic from the memory content leads to retrieval collapse: as the number of atoms grows, the external router struggles to index them precisely, resulting in worse performance than the coarse baseline.
+(3) Self-Activated Retrieval replaces the external router with intrinsic key-value matching (Eq.
+5).
+By ensuring the retrieval condition is strictly aligned with the expert’s content, it resolves the routing ambiguity without extra parameters.
+(4) Sparsity Constraint (Top-$k$) enforces a strict activation budget.
+This serves as a gating mechanism that prevents memory collisions and reduces interference during training.
+(5) Temperature Scaling sharpens the retrieval distribution.
+This concentrates gradient flow on the most relevant “specialist” atoms, accelerating their adaptation while protecting shared memories.
+(6) Threshold-based selection prunes weak activation signals ($\delta$) strictly at test time, maximizing retrieval precision by eliminating noise.
 Rank Activation Budget.
-We probe how many ranks are necessary by varying the activation budget (Fig.
+We probe rank usage by varying the activation budget on a 10-task checkpoint (Fig.
 4(a)).
 To isolate this effect, we disable other sparsity mechanisms (e.g., thresholding).
-After training on 10 tasks (160 ranks total), increasing the budget from 2 (1.25%) to 16 active ranks (10%) yields a large improvement in final accuracy and then gradually plateaus.
-This suggests that each rank specializes in distinct input patterns and that a small, sparse subset of ranks is sufficient to cover diverse inputs.
-Even with a high budget, the self-activating gating remains robust: it still privileges the most relevant ranks per token while suppressing lower-relevance contributions.
-Sharpness control of $\tau_{\text{MoRA}}$.
-We use temperature scaling (Eq.
-(10)) to control mixture sharpness: lower $\tau_{\text{MoRA}}$ yields sharper activations, while higher $\tau_{\text{MoRA}}$ flattens the distribution and engages more ranks.
+Performance improves as the budget increases from 2 to 16 (10% of total ranks) before plateauing, confirming that a small, sparse subset of specialized ranks suffices to cover diverse inputs.
+Crucially, accuracy remains robust even at high budgets, demonstrating that self-activation naturally prioritizes relevant signals while suppressing noise, even without strict sparsity constraints.
+Sharpness enhancement.
+Temperature scaling (Eq.
+(7)) modulates activation sparsity: lower $\tau_{\text{MoRAM}}$ concentrates probability mass, while higher $\tau_{\text{MoRAM}}$ broadens expert participation.
 Fig.
-4(b) shows that increasing $\tau_{\text{MoRA}}$ improves transfer to unseen tasks, even surpassing the reported state-of-the-art Transfer in Table 1 when $\tau_{\text{MoRA}}$ is high.
-A moderate $\tau_{\text{MoRA}}$ around 0.01 balances specificity and generalization, delivering strong Last and Average scores.
+4(b) demonstrates that higher $\tau_{\text{MoRAM}}$ significantly boosts transfer to unseen tasks (even surpassing SOTA in Table 1).
+However, we select a moderate $\tau_{\text{MoRAM}}$ of 0.01 to optimally balance specificity (retention) and generalization (transfer), ensuring strong overall performance.
 Threshold-based rank selection.
@@
 5 Conclusion
-We present MoRA, a self-activated sparse mixture-of-ranks framework that decomposes each LoRA update into rank‐1 experts and self‐infer their sparse mixture weights.
-By reducing knowledge redundancy and resolving router–expert mismatch during CL, MoRA delivers superior continual‐learning performance, significantly reduces forgetting, and enhances generalization.
+We presented MoRAM, a framework that challenges the coarse-grained design of standard MoE-LoRA.
+By redefining adaptation as the accumulation of atomic associative memories, we demonstrated that high-performing CL does not require complex external routers, but rather precise, content-addressable retrieval of low-rank atomic memories.
 Future works.
-MoRA controls mixture sharpness and sparsity via top-k selection, temperature scaling, and thresholding.
+MoRAM controls mixture sharpness and sparsity via top-k selection, temperature scaling, and thresholding.
 While effective, the sparse activation mechanism can be further improved by incorporating awareness of the input data distribution.
@@
 We also aim to extend the method’s applicability to a broader range of PTMs and applications.
-Reproducibility Statement
-Further implementation details and dataset descriptions are provided in Sec. 4 and Appendix A.1. Our source code will be publicly released upon paper acceptance.
+Impact Statement
+MoRAM’s sparse mixture-of-ranks method makes continual adaptation of large vision–language and language models far more efficient and reliable.
+By updating only the most relevant low-rank subspaces, it reduces compute and memory requirements, enabling on-device personalization in domains like education, healthcare, and finance.
+This democratizes access to powerful, continually evolved models for smaller teams and resource-constrained settings, driving innovation and improving quality of life.
+Preserving pre-trained knowledge also ensures stability in safety-critical systems, such as medical diagnostics and autonomous vehicles, where forgetting of pre-trained capabilities can be dangerous.
+The focus of this paper is fundamental research, and is broadly applicable to model fine-tuning techniques.
+Thus, it is possible that MoRAM’s lightweight updates could be misused to insert stealthy backdoors or reinforce hidden biases that are hard to detect.
+Because it leaves most pre-trained parameters untouched, existing biases may persist or become entrenched.
+To mitigate these risks, we recommend strict audit logging of rank-level updates, anomaly detection on sparse changes, robust access controls, and routine bias and fairness assessments alongside any deployment.
+Acknowledgments and Disclosure of Funding
+This work was partially supported by the ARC DECRA Fellowship (DE230101591) awarded to D.
+Gong.
+H.
+Lu is affiliated with CSIRO Data61 through a PhD scholarship and acknowledges the support of the Google PhD Fellowship.
 Appendix A Additional Details and Results of Experiments
+A.1 Detailed Experiment Settings
+Continual Learning of CLIP on X-TAIL and MTIL.
+The MTIL setting consists of 1,201 classes drawn from 11 diverse datasets: Aircraft [53], Caltech101 [25], CIFAR100 [39], DTD [14], EuroSAT [31], Flowers [58], Food [7], MNIST [18], OxfordPet [59], Cars [38], and SUN397 [90].
+In the X-TAIL setting, a total of 10 datasets are used, with CIFAR100 [25] excluded to prevent domain overlap, following the protocol in [91].
+In line with [91], we use a 5-shot split for MTIL and a 16-shot split for X-TAIL.
+We follow the experimental setups in [105, 96, 91] and use the CLIP model with a ViT-B/16 backbone [63] for all experiments.
+By default, MoRAM is applied to every pre-trained weight matrix in both the vision and text encoders, with an initial rank of 16 per update.
+Each task is trained for 500 iterations using AdamW [49] with a learning rate of 5e45e-4.
+During continual learning, we freeze the ranks learned from previous tasks and initialize new $r=16$ ranks for each incoming task.
+The rank activation budget is set to 16 throughout all tasks.
+We set the temperature $\tau_{\text{MoRAM}}=0.1$ and the threshold $\delta=0.2$.
+Continual Learning of LLMs on TRACE.
+We evaluate MoRAM on the TRACE benchmark [81], a comprehensive suite designed to assess continual learning in LLMs such as LLaMA [72] and Gemma [71].
+TRACE standardizes eight diverse datasets spanning domain-specific understanding (C-STANCE, FOMC), reasoning (ScienceQA, NumGLUE-cm/ds), summarization (MeetingBank), code generation (Py150), and multilingual simplification (20Minuten).
+Following standard protocols, each task consists of 5,000 training and 2,000 testing examples.
+We employ task-specific metrics to capture performance nuances: Accuracy for classification and reasoning tasks; ROUGE-L for summarization; SARI for simplification; and similarity scores for code generation.
+We adopt the variable epoch schedule from Qian et al.
+[60] ({2, 1, 3, 2, 1, 2, 2, 3} epochs).
+For each incoming task, we initialize $r=16$ new ranks while maintaining a fixed rank activation budget of $k=16$ across all tasks.
+We set the softmax temperature $\tau_{\text{MoRAM}}=0.03$ and the inference threshold $\delta=0.2$.
+Optimization proceeds with a learning rate of $1\times 10^{-4}$ and a batch size of 4, utilizing a maximum context length of 1024 tokens.
+All experiments leverage DeepSpeed ZeRO-2 with BF16 mixed-precision on a cluster of four Nvidia H100 GPUs.
+Continual Learning of LMs on language classification tasks.
+We follow the protocol of previous work in continually fine-tuning the T5-large [64] and LLaMA2-7B [72] on a suite of text-classification tasks.
+We train on five standard benchmarks—AG News, Amazon Reviews, Yelp Reviews, DBpedia, and Yahoo Answers—using three distinct task orderings drawn from [62, 65, 80, 61].
+To probe longer sequences, we extend this to a 15-dataset stream (Table 5), incorporating tasks from the original CL benchmark [102], GLUE [75], SuperGLUE [74], and the IMDB movie reviews corpus.
+Natural language prompts for each task are presented in Table 6, with NLI tasks (MNLI, RTE, CB), sentiment classification (Amazon, Yelp, SST-2, IMDB), and topic classification (AG News, DBpedia, Yahoo).
+We evaluate three distinct task sequences for both the standard CL and 15-task benchmarks (Table 7).
+After completing the final task in each stream, we report the average accuracy across all tasks.
+All experiments use one epoch per task with DeepSpeed, a fixed learning rate of 1e31e-3, batch size 64, and dropout of 0.1.
+MoRAM is applied to both the query and key projection matrices within attention layers, initializing $r=8$ new ranks for each incoming task similarly as in [80, 61].
+We maintain a constant activation budget of 4 ranks throughout continual learning, set the temperature $\tau_{\text{MoRAM}}=0.1$, and the threshold $\delta=0.2$.
+Generalization and forgetting on unseen tasks after standard fine-tuning.
+To assess effects on pre-trained general knowledge, we fine-tune Llama3.1-8B [28] on the CodeAlpaca code-generation dataset [8] using llama-Factory [104] and evaluate using lm-eval-harness [26] on zero-shot in-domain performance on HumanEval [11], as well as out-of-domain accuracy on a broad selection of MMLU [32] subjects—Formal Logic, Philosophy, World Religions, Economics, Public Relations, STEM, Physics, and Machine Learning.
+In this experiment, MoRAM is applied to all linear weight matrices of the pre-trained model.
+We fine-tune on CodeAlpaca with a batch size of 32 over 3 epochs and a cosine learning-rate schedule starting at 5e45e-4.
+We train with $r=16$ ranks and enforce a constant activation budget of 4 ranks.
+The self-routed gating uses a temperature $\tau_{\text{MoRAM}}=0.5$ and a threshold $\delta=0.2$.
+We observe that, due to variations in hidden representations across architectures, the optimal temperature setting can differ across each pre-trained models.
 Table 5: Details of the 15 datasets used in our continual‐learning experiments using LMs.
@@
 Sequences 1–3 follow the standard CL benchmarks employed in prior work.
-Sequences 4–6 extend to longer 15‐task streams, as introduced in (Razdaibiedina et al., 2023).
+Sequences 4–6 extend to longer 15‐task streams, as introduced in [65].
 Order Task Sequence
@@
 sst-2 → dbpedia → ag → yahoo → multirc → boolqa → wic
-A.1 Detailed Experiment Settings
-CL on CLIP.
-We follow the experimental setups in (Zheng et al., 2023; Yu et al., 2024; Xu et al., 2024) and use the CLIP model with a ViT-B/16 backbone (Radford et al., 2021) for all experiments.
-By default, MoRA is applied to every pre-trained weight matrix in both the vision and text encoders, with an initial rank of 16 per update.
-Each task is trained for 500 iterations using AdamW (Loshchilov & Hutter, 2017) with a learning rate of 5e45e-4.
-During continual learning, we freeze the ranks learned from previous tasks and initialize new $r=16$ ranks for each incoming task.
-The rank activation budget is set to 16 throughout all tasks.
-We set the temperature $\tau_{\text{MoRA}}=0.1$ and the threshold $\delta=0.2$.
-CL on LMs.
-We follow the protocol of previous work in continually fine-tuning the T5-large (Raffel et al., 2020) and LLaMA2-7B (Touvron et al., 2023) on a suite of text-classification tasks.
-We train on five standard benchmarks—AG News, Amazon Reviews, Yelp Reviews, DBpedia, and Yahoo Answers—using three distinct task orderings drawn from (Qin & Joty, 2021; Razdaibiedina et al., 2023; Wang et al., 2023; Qiao & Mahdavi, 2024).
-To probe longer sequences, we extend this to a 15-dataset stream (Table 5), incorporating tasks from the original CL benchmark (Zhang et al., 2015), GLUE (Wang et al., 2018), SuperGLUE (Wang et al., 2019), and the IMDB movie reviews corpus.
-Natural language prompts for each task are presented in Table 6, with NLI tasks (MNLI, RTE, CB), sentiment classification (Amazon, Yelp, SST-2, IMDB), and topic classification (AG News, DBpedia, Yahoo).
-We evaluate three distinct task sequences for both the standard CL and 15-task benchmarks (Table 7).
-After completing the final task in each stream, we report the average accuracy across all tasks.
-All experiments use one epoch per task with DeepSpeed, a fixed learning rate of 1e31e-3, batch size 64, and dropout of 0.1.
-MoRA is applied to both the query and key projection matrices within attention layers, initializing $r=8$ new ranks for each incoming task similarly as in (Wang et al., 2023; Qiao & Mahdavi, 2024).
-We maintain a constant activation budget of 4 ranks throughout continual learning, set the temperature $\tau_{\text{MoRA}}=0.1$, and the threshold $\delta=0.2$.
-Generalization and forgetting on unseen tasks.
-To assess effects on pre-trained general knowledge, we fine-tune Llama3.1-8B (Grattafiori et al., 2024) on the CodeAlpaca code-generation dataset (Chaudhary, 2023) using llama-Factory (Zheng et al., 2024) and evaluate using lm-eval-harness (Gao et al., 2024) on zero-shot in-domain performance on HumanEval (Chen et al., 2021), as well as out-of-domain accuracy on a broad selection of MMLU (Hendrycks et al., 2021) subjects—Formal Logic, Philosophy, World Religions, Economics, Public Relations, STEM, Physics, and Machine Learning.
-In this experiment, MoRA is applied to all linear weight matrices of the pre-trained model.
-We fine-tune on CodeAlpaca with a batch size of 32 over 3 epochs and a cosine learning-rate schedule starting at 5e45e-4.
-We train with $r=16$ ranks and enforce a constant activation budget of 4 ranks.
-The self-routed gating uses a temperature $\tau_{\text{MoRA}}=0.5$ and a threshold $\delta=0.2$.
-We observe that, due to variations in hidden representations across architectures, the optimal temperature setting can differ across each pre-trained models.
-Further details of the used datasets in MTIL and X-TAIL.
-The MTIL setting consists of 1,201 classes drawn from 11 diverse datasets: Aircraft (Maji et al., 2013), Caltech101 (Fei-Fei et al., 2004), CIFAR100 (Krizhevsky et al., 2009), DTD (Cimpoi et al., 2014), EuroSAT (Helber et al., 2019), Flowers (Nilsback & Zisserman, 2008), Food (Bossard et al., 2014), MNIST (Deng, 2012), OxfordPet (Parkhi et al., 2012), Cars (Krause et al., 2013), and SUN397 (Xiao et al., 2010).
-In the X-TAIL setting, a total of 10 datasets are used, with CIFAR100 (Fei-Fei et al., 2004) excluded to prevent domain overlap, following the protocol in (Xu et al., 2024).
-In line with (Xu et al., 2024), we use a 5-shot split for MTIL and a 16-shot split for X-TAIL.
-Table 8: Average accuracy on T5-large continual‐learning benchmarks after the final task, evaluated over extended 15-task sequences. Results for prior methods are taken from (Qiao & Mahdavi, 2024).
+A.2 Evaluation Metrics
+To strictly evaluate the plasticity and stability of our method, we employ the following metrics.
+Let $N$ denote the total number of learned tasks, and $A_{i,j}$ denote the performance (e.g., accuracy, ROUGE, or similarity score) on task $j$ after training on task $i$.
+CLIP Continual Learning Metrics.
+For the CLIP experiments (MTIL and X-TAIL), we report the following metrics as in [105, 96, 91]:
+  • Transfer Accuracy (Transfer): indicates the model’s zero-shot performance on future domains before they are learned, measures the extent to which the pre-trained zero-shot ability is preserved (or improved) throughout incremental learning.
+  • Average Accuracy (Average): indicates the average accuracy of all learning steps across all domains.
+It captures the comprehensive performance stability throughout the entire incremental training process.
+  • Last Accuracy (Last): represent the model’s performance on all seen domains after the training is fully completed.
+LLM Continual Learning Metrics.
+For the TRACE benchmark, we adopt the standard metrics as in [81]:
+  • Overall Performance (OP): The average performance across all learned tasks at the end of training.
+    $\text{OP}=\frac{1}{N}\sum_{j=1}^{N}A_{N,j},$ (9)
+    where $A_{i,j}$ represents the performance on task $j$ after learning task $i$.
+  • Backward Transfer (BWT): Measures the average performance degradation (forgetting) of previous tasks $j<N$ after learning new tasks. A lower BWT indicates less forgetting (better stability).
+    $\text{BWT}=\frac{1}{N-1}\sum_{j=1}^{N-1}(A_{j,j}-A_{N,j}).$ (10)
+A.3 Continual Learning of LMs on the Standard-CL benchmark.
+Table 8 reports results across three task orderings: MoRAM consistently outperforms prior methods and closely approaches the multi-task learning (MTL) upper bound.
+Unlike O-LoRA [80] and LB-CL [61], which rely on orthogonality constraints or gradient projections between per-task LoRA adapters (potentially limiting adapter capacity), MoRAM needs no extra regularization.
+By decomposing each rank-r update into rank-one components and applying self-activated, sparse gating, MoRAM lets each component specialize on its own input distribution, reducing interference and more effectively capturing diverse patterns.
+We further evaluate MoRAM on LLaMA2-7B [72] under the same continual-learning setup (Table 10 in the Appendix), MoRAM also outperforms O-LoRA by 2.3% averaged over 3 task orders.
+A.4 Continual Learning of LMs on Long Task Sequences
+In Table 8, we evaluate MoRAM on standard CL benchmarks [65].
+In Table 9, we extend the evaluation to challenged long task sequences using 15 datasets with 3 different orderings as in [80, 61].
+Consistent with the findings in Table 8, MoRAM outperforms previous methods in terms of averaged performance across three task orders, and largely closes the gap to multi-task learning.
+Two key design choices drive this robust performance in the long‐sequence regime.
+First, by decomposing each LoRA update into fine‐grained rank-1 components and enforcing a small, fixed activation budget, MoRAM encourages each rank to specialize on a narrow subspace of the data manifold.
+At inference time, only the most relevant subspaces are activated for a given input, which preserves earlier task representations and prevents catastrophic interference.
+Second, our self‐routed gating mechanism enables each rank to assess its own relevance on a per‐token basis, yielding stable mixture patterns as the expert pool grows.
+Coupling with our proposed rank pruning, these mechanisms ensure that MoRAM continually incorporates new knowledge only when needed while robustly maintaining prior capabilities.
+Table 8: Summary of results on standard CL benchmarks with T5-large. We report the averaged accuracy after training on the last task across three task orderings.
+Standard CL Benchmark
+Method Order-1 Order-2 Order-3 Avg.
+MTL 80.0
+SeqFT 18.9 24.9 41.7 28.5
+SeqLoRA 44.6 32.7 53.7 43.7
+IncLoRA 66.0 64.9 68.3 66.4
+Replay 55.2 56.9 61.3 57.8
+EWC 48.7 47.7 54.5 50.3
+LwF 54.4 53.1 49.6 52.3
+L2P 60.3 61.7 61.1 60.7
+LFPT5 67.6 72.6 77.9 72.7
+InfLoRA 75.2 75.4 75.8 75.5
+O-LoRA 75.4 75.7 76.3 75.8
+LB-CL 76.9 76.5 76.8 76.7
+MoRAM 77.4 77.5 77.9 77.6
+Table 9: Average accuracy on T5-large continual‐learning benchmarks after the final task, evaluated over extended 15-task sequences. Results for prior methods are taken from [61].
 Large Number of Tasks
@@
 LB-CL 68.4 67.3 71.8 69.2
-MoRA 68.91 68.32 71.95 69.72
+MoRAM 68.91 68.32 71.95 69.72
 Table 10: Continual learning results on standard CL benchmarks with the LLaMA2-7B model.
@@
 O-LoRA 76.8 75.7 75.7 76.1
-MoRA 77.8 78.0 79.3 78.4
-A.2 Continual Learning of LMs on Long Task Sequences
-In Sec.
-4.2 Table 3, we evaluate MoRA on standard CL benchmarks (Razdaibiedina et al., 2023).
-In Table 8, we extend the evaluation to challenged long task sequences using 15 datasets with 3 different orderings as in (Wang et al., 2023; Qiao & Mahdavi, 2024).
-Consistent with the findings in Table 3, MoRA outperforms previous methods in terms of averaged performance across three task orders, and largely close the gap to multi-task learning.
-Two key design choices drive this robust performance in the long‐sequence regime.
-First, by decomposing each LoRA update into fine‐grained rank-1 components and enforcing a small, fixed activation budget, MoRA encourages each rank to specialize on a narrow subspace of the data manifold.
-At inference time, only the most relevant subspaces are activated for a given input, which preserves earlier task representations and prevents catastrophic interference.
-Second, our self‐routed gating mechanism enables each rank to assess its own relevance on a per‐token basis, yielding stable mixture patterns as the expert pool grows.
-Coupling with our proposed rank pruning, these mechanisms ensure that MoRA continually incorporates new knowledge only when needed while robustly maintaining prior capabilities.
+MoRAM 77.8 78.0 79.3 78.4
 Table 11: Comparisons of trainable parameters for each pre-trained weight matrix during continual learning of each task.
 MoE-LoRA and MoE-Adapter trains additional router module with LoRA experts.
-In MoRA, $k$ denotes the rank activation budget (with $k\leq r$).
+In MoRAM, $k$ denotes the rank activation budget (with $k\leq r$).
 Method Trainable parameters per task
-LoRA (Hu et al., 2021) $r\,(d_{\text{in}}+d_{\text{out}})$
+LoRA [33] $r\,(d_{\text{in}}+d_{\text{out}})$
 MoE-LoRA (1 expert/task) $r\,(d_{\text{in}}+d_{\text{out}})\;+\;d_{\text{in}}$
-MoE-Adapter (2 experts/task) (Yu et al., 2024) $2(r\,(d_{\text{in}}+d_{\text{out}})+d_{\text{in}})$
-CoDyRA (Lu et al., 2024) $r\,(d_{\text{in}}+d_{\text{out}})+r$
-O-LoRA (Wang et al., 2023) $r\,(d_{\text{in}}+d_{\text{out}})$
-LB-CL (Qiao & Mahdavi, 2024) $r\,(d_{\text{in}}+d_{\text{out}})+r$
-MoRA $rd_{\text{in}}+kd_{\text{out}}$
-A.3 Multi‐domain task incremental learning.
-We evaluate MoRA in the few‐shot MTIL setting (Table 11) under the same protocols as (Yu et al., 2024; Xu et al., 2024; Lu et al., 2024).
+MoE-Adapter (2 experts/task) [96] $2(r\,(d_{\text{in}}+d_{\text{out}})+d_{\text{in}})$
+CoDyRA [50] $r\,(d_{\text{in}}+d_{\text{out}})+r$
+O-LoRA [80] $r\,(d_{\text{in}}+d_{\text{out}})$
+LB-CL [61] $r\,(d_{\text{in}}+d_{\text{out}})+r$
+MoRAM $rd_{\text{in}}+kd_{\text{out}}$
+A.5 Multi‐domain task incremental learning.
+We evaluate MoRAM in the few‐shot MTIL setting (Table 12) under the same protocols as [96, 91, 50].
 Consistent with the results observed in the X-TAIL setting, our method demonstrates clear superiority in this scenario.
@@
 These findings validate that our self-activated sparse mixture-of-ranks framework both facilitates continual acquisition of new knowledge and mitigates forgetting from the pre-trained model and earlier tasks.
-Table 11: Comparisons on 5-shot MTIL setting. Following the same protocol as in (Yu et al., 2024; Xu et al., 2024; Lu et al., 2024).
-Method
-Aircraft Caltech101 CIFAR100 DTD EuroSAT Flowers Food MNIST OxfordPet Cars SUN397 Average
+Table 12: Comparisons on 5-shot MTIL setting. Following the same protocol as in [96, 91, 50].
+    Method Aircraft Caltech101 CIFAR100 DTD EuroSAT Flowers Food MNIST OxfordPet Cars SUN397 Average
 CLIP
-   Zero-shot (Radford et al., 2021) 24.3 88.4 68.2 44.6 54.9 71.0 88.5 59.4 89.0 64.7 65.2 65.3
+   Zero-shot [63] 24.3 88.4 68.2 44.6 54.9 71.0 88.5 59.4 89.0 64.7 65.2 65.3
 Transfer
-   Zero-shot (Radford et al., 2021) 88.4 68.2 44.6 54.9 71.0 88.5 59.6 89.0 64.7 65.2 69.4
-   LwF (Li & Hoiem, 2017) 72.1 49.2 35.9 44.5 41.1 66.6 50.5 69.0 19.0 51.7 50.0
-   LwF-VR (Ding et al., 2022) 82.2 62.5 40.1 40.1 56.3 80.0 60.9 77.6 40.5 60.8 60.1
-   WiSE-FT (Wortsman et al., 2022) 77.6 60.0 41.3 39.4 53.0 76.6 58.1 75.5 37.3 58.2 57.7
-   ZSCL (Zheng et al., 2023) 84.0 68.1 44.8 46.8 63.6 84.9 61.4 81.4 55.5 62.2 65.3
-   MoE-Adapter (Yu et al., 2024) 87.9 68.2 44.1 48.1 64.7 88.8 69.0 89.1 64.5 65.1 68.9
-   RAIL-Primal (Xu et al., 2024) 88.4 68.2 44.6 54.9 71.0 88.5 59.6 89.0 64.7 65.2 69.4
-   CoDyRA (Lu et al., 2024) 92.4 68.4 45.8 54.5 69.6 87.4 65.2 88.5 64.2 64.5 69.9
-   MoRA 92.0 68.8 45.6 53.1 68.6 84.4 64.3 89.8 65.4 64.8 69.7
+   Zero-shot [63] 88.4 68.2 44.6 54.9 71.0 88.5 59.6 89.0 64.7 65.2 69.4
+   LwF [44] 72.1 49.2 35.9 44.5 41.1 66.6 50.5 69.0 19.0 51.7 50.0
+   LwF-VR [20] 82.2 62.5 40.1 40.1 56.3 80.0 60.9 77.6 40.5 60.8 60.1
+   WiSE-FT [86] 77.6 60.0 41.3 39.4 53.0 76.6 58.1 75.5 37.3 58.2 57.7
+   ZSCL [105] 84.0 68.1 44.8 46.8 63.6 84.9 61.4 81.4 55.5 62.2 65.3
+   MoE-Adapter [96] 87.9 68.2 44.1 48.1 64.7 88.8 69.0 89.1 64.5 65.1 68.9
+   RAIL-Primal [91] 88.4 68.2 44.6 54.9 71.0 88.5 59.6 89.0 64.7 65.2 69.4
+   CoDyRA [50] 92.4 68.4 45.8 54.5 69.6 87.4 65.2 88.5 64.2 64.5 69.9
+   MoRAM 92.0 68.8 45.6 53.1 68.6 84.4 64.3 89.8 65.4 64.8 69.7
 Average
-   LwF (Li & Hoiem, 2017) 23.5 77.4 43.5 41.7 43.5 52.2 54.6 63.4 68.0 21.3 52.6 49.2
-   LwF-VR (Ding et al., 2022) 24.9 89.1 64.2 53.4 54.3 70.8 79.2 66.5 79.2 44.1 61.6 62.5
-   WiSE-FT (Wortsman et al., 2022) 32.0 87.7 61.0 55.8 68.1 69.3 76.8 71.5 77.6 42.0 59.3 63.7
-   ZSCL (Zheng et al., 2023) 28.2 88.6 66.5 53.5 56.3 73.4 83.1 56.4 82.4 57.5 62.9 64.4
-   MoE-Adapter (Yu et al., 2024) 30.0 89.6 73.9 58.7 69.3 79.3 88.1 76.5 89.1 65.3 65.8 71.4
-   RAIL-Primal (Xu et al., 2024) 32.9 94.5 69.9 58.1 71.8 84.4 88.5 70.4 89.0 66.1 65.7 71.9
-   CoDyRA (Lu et al., 2024) 34.6 95.8 73.9 60.0 77.1 81.3 86.6 75.9 89.9 66.1 65.3 73.3
-   MoRA 36.7 95.4 74.9 61.9 77.1 82.6 85.3 76.0 90.5 67.0 65.6 73.9
+   LwF [44] 23.5 77.4 43.5 41.7 43.5 52.2 54.6 63.4 68.0 21.3 52.6 49.2
+   LwF-VR [20] 24.9 89.1 64.2 53.4 54.3 70.8 79.2 66.5 79.2 44.1 61.6 62.5
+   WiSE-FT [86] 32.0 87.7 61.0 55.8 68.1 69.3 76.8 71.5 77.6 42.0 59.3 63.7
+   ZSCL [105] 28.2 88.6 66.5 53.5 56.3 73.4 83.1 56.4 82.4 57.5 62.9 64.4
+   MoE-Adapter [96] 30.0 89.6 73.9 58.7 69.3 79.3 88.1 76.5 89.1 65.3 65.8 71.4
+   RAIL-Primal [91] 32.9 94.5 69.9 58.1 71.8 84.4 88.5 70.4 89.0 66.1 65.7 71.9
+   CoDyRA [50] 34.6 95.8 73.9 60.0 77.1 81.3 86.6 75.9 89.9 66.1 65.3 73.3
+   MoRAM 36.7 95.4 74.9 61.9 77.1 82.6 85.3 76.0 90.5 67.0 65.6 73.9
 Last
-   LwF (Li & Hoiem, 2017) 22.1 58.2 17.9 32.1 28.1 66.7 46.0 84.3 64.1 31.5 60.1 46.5
-   LwF-VR (Ding et al., 2022) 22.9 89.8 59.3 57.1 57.6 79.2 78.3 77.7 83.6 60.1 69.8 66.9
-   WiSE-FT (Wortsman et al., 2022) 30.8 88.9 59.6 60.3 80.9 81.7 77.1 94.9 83.2 62.8 70.0 71.9
-   ZSCL (Zheng et al., 2023) 26.8 88.5 63.7 55.7 60.2 82.1 82.6 58.6 85.9 66.7 70.4 67.4
-   MoE-Adapter (Yu et al., 2024) 30.1 89.3 74.9 64.0 82.3 89.4 87.1 89.0 89.1 69.5 72.5 76.1
-   RAIL-Primal (Xu et al., 2024) 32.9 95.1 70.3 63.2 81.5 95.6 88.5 89.7 89.0 72.5 71.0 77.2
-   CoDyRA (Lu et al., 2024) 31.6 95.5 72.8 63.5 85.0 89.7 85.0 94.7 93.2 73.6 73.0 78.0
-   MoRA 32.5 95.3 75.3 66.6 87.8 92.6 86.3 96.3 92.6 73.5 73.8 79.3
-A.4 More Comparison on HumanEval for Code Generation
-To further evaluate code generation performance, we compare MoRA against LoRI-D and LoRI-S (Zhang et al., 2025) on the HumanEval benchmark (Table 12).
-MoRA consistently outperforms both LoRI variants across all metrics, achieving notable improvements in Pass@1, Pass@5, and especially Pass@10.
+   LwF [44] 22.1 58.2 17.9 32.1 28.1 66.7 46.0 84.3 64.1 31.5 60.1 46.5
+   LwF-VR [20] 22.9 89.8 59.3 57.1 57.6 79.2 78.3 77.7 83.6 60.1 69.8 66.9
+   WiSE-FT [86] 30.8 88.9 59.6 60.3 80.9 81.7 77.1 94.9 83.2 62.8 70.0 71.9
+   ZSCL [105] 26.8 88.5 63.7 55.7 60.2 82.1 82.6 58.6 85.9 66.7 70.4 67.4
+   MoE-Adapter [96] 30.1 89.3 74.9 64.0 82.3 89.4 87.1 89.0 89.1 69.5 72.5 76.1
+   RAIL-Primal [91] 32.9 95.1 70.3 63.2 81.5 95.6 88.5 89.7 89.0 72.5 71.0 77.2
+   CoDyRA [50] 31.6 95.5 72.8 63.5 85.0 89.7 85.0 94.7 93.2 73.6 73.0 78.0
+   MoRAM 32.5 95.3 75.3 66.6 87.8 92.6 86.3 96.3 92.6 73.5 73.8 79.3
+A.6 More Comparisons on CL on LMs
+Results on SuperNI benchmark.
+To provide a complete comparison, we additionally evaluate MoRAM against SAPT [103] on the SuperNI benchmark [83] using the T5-Large backbone (Table 13).
+We report results in both replay-based and replay-free settings.
+Under the replay-based protocol introduced by SAPT, where pseudo-samples are generated using a trained generative model, MoRAM attains a stronger stability–plasticity trade-off (AP: 51.79% vs.
+51.54%; FT: 0.73% vs.
+0.91%).
+In the replay-free setting, which is the primary use case for MoRAM, our method achieves state-of-the-art performance (AP: 39.62%) and substantially outperforms parameter-efficient baselines such as O-LoRA (26.37%) and L2P (15.18%).
+These results highlight the complementary design choices.
+SAPT improves replay efficiency through shared attention prompts, while MoRAM relies on architectural isolation via sparse rank-1 experts.
+As a result, MoRAM remains effective without a memory buffer, yet can also incorporate replay to further improve performance beyond SAPT.
+Discussions on performance robustness.
+To assess order robustness, we incorporate the OPD metric proposed by [95], which measures a model’s sensitivity to the sequence of arriving tasks.
+Following standard practice for evaluating global performance stability, we compute the standard deviation of the final average accuracy over the $K$ task orders considered.
+In Table 14, using results on the Standard CL Benchmark with three distinct task orders, MoRAM demonstrates substantially improved robustness to task ordering.
+The disparity across orders is 0.26 for MoRAM, which is approximately half of that of O-LoRA (0.46).
+This indicates that the Self-Activated Sparse Mixture mechanism effectively reduces task interference and mitigates the unidirectional knowledge transfer effects identified in prior work.
+Table 13: Overall results on the SuperNI Benchmark using the T5-Large backbone.
+We report Average Performance (AP, $\uparrow$) and Forgetting (FT, $\downarrow$).
+The best results for the stability-plasticity trade-off are highlighted in bold for methods with and without memory replay, respectively.
+Methods Replay SuperNI
+AP FT
+Replay-Based Methods
+Replay 35.37 16.92
+SAPT 51.54 0.91
+MoRAM 51.79 0.73
+Replay-Free Methods
+L2P 15.18 3.65
+IncLoRA 12.33 41.93
+C-LoRA 22.69 24.25
+O-LoRA 26.37 19.15
+MoRAM 39.62 5.74
+Table 14: Order Robustness Analysis on standard CL benchmark with T5-Large. We report the accuracy for each order and the average accuracy $\pm$ the standard deviation.
+Method Order 1 Order 2 Order 3 Avg $\pm$ Std
+SeqFT 18.9 24.9 41.7 28.5 $\pm$ 11.82
+L2P 60.3 61.7 61.1 61.0 $\pm$ 0.70
+LFPT5 67.6 72.6 77.9 72.7 $\pm$ 5.15
+O-LoRA 75.4 75.7 76.3 75.8 $\pm$ 0.46
+MoRAM 77.4 77.5 77.9 77.6 $\pm$ 0.26
+A.7 More Comparison on HumanEval for Code Generation
+To further evaluate code generation performance, we compare MoRAM against LoRI-D and LoRI-S [99] on the HumanEval benchmark (Table 15).
+MoRAM consistently outperforms both LoRI variants across all metrics, achieving notable improvements in Pass@1, Pass@5, and especially Pass@10.
 Table 15: Performance comparison on the HumanEval benchmark, reported in terms of Pass@1, Pass@5, and Pass@10.
@@
 LoRI-S 41.3 54.4 59.6
-MoRA 47.6 60.9 70.1
-A.5 performance robustness
-MoRA employs a sparse mixture of previously learned and newly introduced rank-1 experts to capture both shared and task-specific knowledge, resulting in substantially improved Last performance.
+MoRAM 47.6 60.9 70.1
+A.8 Robustness
+MoRAM employs a sparse mixture of previously learned and newly introduced rank-1 experts to capture both shared and task-specific knowledge, resulting in substantially improved Last performance.
 To assess statistical significance and robustness, we report mean and standard deviation over three independent runs (Table 16).
-MoRA consistently outperforms competing methods across all metrics and exhibits lower variance, highlighting its effectiveness and stability in continual-learning scenarios.
-Table 13: Comparison to InfLoRA and performance robustness. We report mean and standard deviation across 3 independent runs. Best performances are marked in bold.
-Method
-Cars Aircraft OxfordPet Food SUN397 MNIST Flowers DTD Caltech101 EuroSAT Average
+MoRAM consistently outperforms competing methods across all metrics and exhibits lower variance, highlighting its effectiveness and stability in continual-learning scenarios.
+Table 16: Comparison to InfLoRA and performance robustness. We report the mean and standard deviation across 3 independent runs. Best performances are marked in bold.
+      Method Cars Aircraft OxfordPet Food SUN397 MNIST Flowers DTD Caltech101 EuroSAT Average
 Transfer
@@
      CoDyRA $74.3^{\pm 0.52}$ $36.8^{\pm 0.23}$ $44.2^{\pm 0.56}$ $\mathbf{69.9^{\pm 0.56}}$ $\mathbf{83.5^{\pm 0.23}}$ $42.8^{\pm 0.18}$ $\mathbf{88.9^{\pm 0.42}}$ $64.6^{\pm 0.47}$ $\mathbf{63.4^{\pm 0.56}}$ $63.2^{\pm 0.28}$
-   MoRA $\mathbf{74.5^{\pm 0.51}}$ $\mathbf{38.1^{\pm 0.24}}$ $\mathbf{46.9^{\pm 0.56}}$ $65.3^{\pm 0.44}$ $82.9^{\pm 0.18}$ $\mathbf{45.8^{\pm 0.31}}$ $88.2^{\pm 0.15}$ $\mathbf{65.1^{\pm 0.35}}$ $62.9^{\pm 0.10}$ $\mathbf{63.3^{\pm 0.26}}$
+     MoRAM $\mathbf{74.5^{\pm 0.51}}$ $\mathbf{38.1^{\pm 0.24}}$ $\mathbf{46.9^{\pm 0.56}}$ $65.3^{\pm 0.44}$ $82.9^{\pm 0.18}$ $\mathbf{45.8^{\pm 0.31}}$ $88.2^{\pm 0.15}$ $\mathbf{65.1^{\pm 0.35}}$ $62.9^{\pm 0.10}$ $\mathbf{63.3^{\pm 0.26}}$
 Average
@@
      CoDyRA $41.4^{\pm 0.28}$ $81^{\pm 0.38}$ $58.7^{\pm 0.26}$ $77.8^{\pm 0.47}$ $83.4^{\pm 0.39}$ $\mathbf{84.6^{\pm 0.28}}$ $64.5^{\pm 0.14}$ $\mathbf{90.4^{\pm 0.40}}$ $67.2^{\pm 0.23}$ $\mathbf{64.4^{\pm 0.47}}$ $71.3^{\pm 0.18}$
-   MoRA $\mathbf{44.1^{\pm 0.24}}$ $\mathbf{81.6^{\pm 0.34}}$ $\mathbf{64.6^{\pm 0.34}}$ $\mathbf{79.6^{\pm 0.37}}$ $\mathbf{83.9^{\pm 0.36}}$ $84.4^{\pm 0.15}$ $\mathbf{66.5^{\pm 0.24}}$ $89.7^{\pm 0.07}$ $\mathbf{68.4^{\pm 0.38}}$ $64.1^{\pm 0.09}$ $\mathbf{72.7^{\pm 0.17}}$
+     MoRAM $\mathbf{44.1^{\pm 0.24}}$ $\mathbf{81.6^{\pm 0.34}}$ $\mathbf{64.6^{\pm 0.34}}$ $\mathbf{79.6^{\pm 0.37}}$ $\mathbf{83.9^{\pm 0.36}}$ $84.4^{\pm 0.15}$ $\mathbf{66.5^{\pm 0.24}}$ $89.7^{\pm 0.07}$ $\mathbf{68.4^{\pm 0.38}}$ $64.1^{\pm 0.09}$ $\mathbf{72.7^{\pm 0.17}}$
 Last
@@
      CoDyRA $\mathbf{37.7^{\pm 0.42}}$ $81.5^{\pm 0.24}$ $65.1^{\pm 0.63}$ $89.9^{\pm 0.55}$ $91.4^{\pm 0.38}$ $85.5^{\pm 0.16}$ $96.8^{\pm 0.08}$ $\mathbf{93.3^{\pm 0.30}}$ $77.3^{\pm 0.66}$ $73.5^{\pm 0.21}$ $79.2^{\pm 0.18}$
-   MoRA $\mathbf{37.7^{\pm 0.28}}$ $81.5^{\pm 0.22}$ $\mathbf{70.7^{\pm 0.49}}$ $\mathbf{92.4^{\pm 0.20}}$ $\mathbf{95^{\pm 0.34}}$ $\mathbf{86^{\pm 0.13}}$ $\mathbf{97.6^{\pm 0.19}}$ $92.6^{\pm 0.10}$ $\mathbf{81^{\pm 0.35}}$ $\mathbf{74.7^{\pm 0.06}}$ $\mathbf{80.9^{\pm 0.12}}$
-A.6 More comparisons on X-TAIL
-To further validate MoRA’s robustness, we evaluated it under other continual-learning task orderings, i.e., X-TAIL (Order 2), as shown in Table 14.
-The results align with those in Table 1, confirming that MoRA consistently achieves state-of-the-art performance.
+     MoRAM $\mathbf{37.7^{\pm 0.28}}$ $81.5^{\pm 0.22}$ $\mathbf{70.7^{\pm 0.49}}$ $\mathbf{92.4^{\pm 0.20}}$ $\mathbf{95^{\pm 0.34}}$ $\mathbf{86^{\pm 0.13}}$ $\mathbf{97.6^{\pm 0.19}}$ $92.6^{\pm 0.10}$ $\mathbf{81^{\pm 0.35}}$ $\mathbf{74.7^{\pm 0.06}}$ $\mathbf{80.9^{\pm 0.12}}$
 Table 17: Comparisons on X-TAIL (Order 2) for each domain in terms of “Transfer”, “Average”, and “Last” scores (%).
-Method
-Cars Aircraft OxfordPet Food SUN397 MNIST Flowers DTD Caltech101 EuroSAT Average
+    Method Cars Aircraft OxfordPet Food SUN397 MNIST Flowers DTD Caltech101 EuroSAT Average
 CLIP
@@
 Transfer
-   Zero-shot (Radford et al., 2021) 23.5 86.7 84 63.7 46.7 63.6 37.3 76.8 36.7 57.7
-   LwF (Li & Hoiem, 2017) 20.0 74.1 79.6 58.1 34.1 48.9 27.7 64.4 15.1 46.9
-   WiSE-FT (Wortsman et al., 2022) 21.3 79.5 83.3 61.0 39.9 56.5 29.6 68.0 20.8 51.1
-   ZSCL (Zheng et al., 2023) 23.0 84.3 87.2 63.0 42.1 65.2 34.6 71.4 40.9 56.9
-   MoE-Adapter (Yu et al., 2024) 17.1 87.2 87.5 58.4 12.6 65.5 35.9 70.0 17.9 50.2
-   RAIL-Primal (Xu et al., 2024) 23.5 86.7 84 63.7 46.7 63.6 37.3 76.8 36.7 57.7
-   CoDyRA (Lu et al., 2024) 23.6 89.2 83 62 51 71.4 38 77.4 39 59.4
-   MoRA 23.6 88.7 83.4 62.6 51.2 69.9 39.3 77.5 39 59.5
+   Zero-shot [63] 23.5 86.7 84 63.7 46.7 63.6 37.3 76.8 36.7 57.7
+   LwF [44] 20.0 74.1 79.6 58.1 34.1 48.9 27.7 64.4 15.1 46.9
+   WiSE-FT [86] 21.3 79.5 83.3 61.0 39.9 56.5 29.6 68.0 20.8 51.1
+   ZSCL [105] 23.0 84.3 87.2 63.0 42.1 65.2 34.6 71.4 40.9 56.9
+   MoE-Adapter [96] 17.1 87.2 87.5 58.4 12.6 65.5 35.9 70.0 17.9 50.2
+   RAIL-Primal [91] 23.5 86.7 84 63.7 46.7 63.6 37.3 76.8 36.7 57.7
+   CoDyRA [50] 23.6 89.2 83 62 51 71.4 38 77.4 39 59.4
+   MoRAM 23.6 88.7 83.4 62.6 51.2 69.9 39.3 77.5 39 59.5
 Average
-   LwF (Li & Hoiem, 2017) 49.0 27.4 69.7 83.0 65.7 42.2 63.5 33.1 68.5 17.5 52.0
-   WiSE-FT (Wortsman et al., 2022) 57.9 29.6 77.8 85.4 68.0 51.6 69.3 35.5 71.0 23.0 56.9
-   ZSCL (Zheng et al., 2023) 74.4 36.4 86.7 88.7 68.9 50.0 75.1 40.1 72.5 43.7 63.6
-   MoE-Adapter (Yu et al., 2024) 74.4 38.6 87.7 87.3 67.9 50.6 76.5 43.7 72.3 18.8 61.8
-   RAIL-Primal (Xu et al., 2024) 77.9 40.4 85.6 83.3 68.3 62.2 76.6 45.8 80.4 41.7 66.2
-   CoDyRA (Lu et al., 2024) 80 39.2 92.5 85.2 69.2 73.7 79.6 46.2 78.6 44.1 68.8
-   MoRA 80.2 40.1 92.5 84.7 70.1 74 80.1 48.7 78.4 44.4 69.3
+   LwF [44] 49.0 27.4 69.7 83.0 65.7 42.2 63.5 33.1 68.5 17.5 52.0
+   WiSE-FT [86] 57.9 29.6 77.8 85.4 68.0 51.6 69.3 35.5 71.0 23.0 56.9
+   ZSCL [105] 74.4 36.4 86.7 88.7 68.9 50.0 75.1 40.1 72.5 43.7 63.6
+   MoE-Adapter [96] 74.4 38.6 87.7 87.3 67.9 50.6 76.5 43.7 72.3 18.8 61.8
+   RAIL-Primal [91] 77.9 40.4 85.6 83.3 68.3 62.2 76.6 45.8 80.4 41.7 66.2
+   CoDyRA [50] 80 39.2 92.5 85.2 69.2 73.7 79.6 46.2 78.6 44.1 68.8
+   MoRAM 80.2 40.1 92.5 84.7 70.1 74 80.1 48.7 78.4 44.4 69.3
 Last
-   LwF (Li & Hoiem, 2017) 29.6 17.5 63.0 83.8 67.7 44.9 79.3 44.8 84.6 39.0 55.4
-   WiSE-FT (Wortsman et al., 2022) 46.1 23.5 71.3 85.7 70.2 59.1 85.5 47.9 82.4 42.8 61.5
-   ZSCL (Zheng et al., 2023) 71.7 35.3 86.5 89.2 71.8 52.3 89.8 52.0 77.1 68.4 69.4
-   MoE-Adapter (Yu et al., 2024) 75.1 41.1 87.9 87.1 74.1 89.7 92.6 61.2 81.0 27.4 71.7
-   RAIL-Primal (Xu et al., 2024) 77.7 41.9 86.1 83.3 71.8 91.6 97.3 66.4 94.8 86.9 79.8
-   CoDyRA (Lu et al., 2024) 79 38.6 92.6 86.4 74.7 95.2 93 64.7 81.9 92.2 79.8
-   MoRA 79.3 38.9 93.1 85.4 74.9 96.4 94.1 69.9 82 92.9 80.7
+   LwF [44] 29.6 17.5 63.0 83.8 67.7 44.9 79.3 44.8 84.6 39.0 55.4
+   WiSE-FT [86] 46.1 23.5 71.3 85.7 70.2 59.1 85.5 47.9 82.4 42.8 61.5
+   ZSCL [105] 71.7 35.3 86.5 89.2 71.8 52.3 89.8 52.0 77.1 68.4 69.4
+   MoE-Adapter [96] 75.1 41.1 87.9 87.1 74.1 89.7 92.6 61.2 81.0 27.4 71.7
+   RAIL-Primal [91] 77.7 41.9 86.1 83.3 71.8 91.6 97.3 66.4 94.8 86.9 79.8
+   CoDyRA [50] 79 38.6 92.6 86.4 74.7 95.2 93 64.7 81.9 92.2 79.8
+   MoRAM 79.3 38.9 93.1 85.4 74.9 96.4 94.1 69.9 82 92.9 80.7
+A.9 More comparisons on X-TAIL
+To further validate MoRAM ’s robustness, we evaluated it under other continual-learning task orderings, i.e., X-TAIL (Order 2), as shown in Table 17.
+The results align with those in Table 1, confirming that MoRAM consistently achieves state-of-the-art performance.
 Table 18: Trainable parameters and averaged training GPU memory per task.
 Method Trainable Params. (Million) GPU Mem. (MiB)
-LWF (Li & Hoiem, 2017) 129.6 32172
-ZSCL (Zheng et al., 2023) 129.6 26290
-MoE-Adapters (Yu et al., 2024) 59.8 22358
-CoDyRA (Lu et al., 2024) 4.4 21770
-MoRA 4.4 21090
-A.7 Computation Cost
+LWF [44] 129.6 32172
+ZSCL [105] 129.6 26290
+MoE-Adapters [96] 59.8 22358
+CoDyRA [50] 4.4 21770
+MoRAM 4.4 21090
+A.10 Computation Cost
 Table 11 summarizes the per‐task trainable parameters of various continual‐learning methods.
-Standard LoRA (Hu et al., 2021), CoDyRA (Lu et al., 2024), and O-LoRA (Wang et al., 2023) each introduce $r(d_{\text{in}}+d_{\text{out}})$ new parameters per weight matrix.
-Mixture‐of‐Experts variants such as MoE-LoRA and MoE-Adapter (Yu et al., 2024) additionally train a router module to controll the usage of each LoRA experts, inducing $d_{\text{in}}$ additional paramters for each experts.
-LB-CL (Qiao & Mahdavi, 2024) introduce $r$ additional parameters, mimic the singular values of SVD.
-By contrast, MoRA requires only $rd_{\text{in}}+kd_{\text{out}}$ activated trainable parameters per task, where $k\leq r$ is the activation budget, and the trainable parameter is at most the same as a standard LoRA.
-Despite the small number of parameter activated and trained, MoRA achieves superior continual learning performance, and reaches comparable performance in general fine-tuning with only one‐third of the activated parameters of a standard LoRA.
+Standard LoRA [33], CoDyRA [50], and O-LoRA [80] each introduce $r(d_{\text{in}}+d_{\text{out}})$ new parameters per weight matrix.
+Mixture‐of‐Experts variants such as MoE-LoRA and MoE-Adapter [96] additionally train a router module to control the usage of each LoRA expert, inducing $d_{\text{in}}$ additional parameters for each expert.
+LB-CL [61] introduces $r$ additional parameters, mimic the singular values of SVD.
+By contrast, MoRAM requires only $rd_{\text{in}}+kd_{\text{out}}$ activated trainable parameters per task, where $k\leq r$ is the activation budget, and the trainable parameter is at most the same as a standard LoRA.
+Despite the small number of parameters activated and trained, MoRAM achieves superior continual learning performance, and reaches comparable performance in general fine-tuning with only one‐third of the activated parameters of a standard LoRA.
 Trainable parameters and training GPU memory.
 Beyond the estimated parameter counts in Table 11, in Table 18, we measured the actual trainable parameters for each continual‐learning task and GPU memory usage, under the same settings as Table 1 in the main paper.
-LWF (Li & Hoiem, 2017) and ZSCL (Zheng et al., 2023) perform full‐parameter fine‐tuning, consuming the most parameters and memory.
-MoE‐Adapters (Yu et al., 2024) maintains a router with 22 rank‐64 adapter experts (top-2 activated) and a DDAS domain predictor.
-CoDyRA (Lu et al., 2024) trains a single rank‐16 LoRA per task, reducing its footprint to 4.4 M parameters.
-MoRA introduces 16 rank-1 experts per task, with no additional router, for a total of 4.4 M trainable parameters and keeps a low gpu memory usage, thanks to our novel self‐activated sparse mixture of ranks design.
+LWF [44] and ZSCL [105] perform full‐parameter fine‐tuning, consuming the most parameters and memory.
+MoE‐Adapters [96] maintains a router with 22 rank‐64 adapter experts (top-2 activated) and a DDAS domain predictor.
+CoDyRA [50] trains a single rank‐16 LoRA per task, reducing its footprint to 4.4 M parameters.
+MoRAM introduces 16 rank-1 experts per task, with no additional router, for a total of 4.4 M trainable parameters and keeps a low GPU memory usage, thanks to our novel self‐activated sparse mixture of ranks design.
+Table 19: Analysis of post-pruning. After training of each task, we retain the subset required to capture the top 99% of the total activation.
+      Method Cars Aircraft OxfordPet Food SUN397 MNIST Flowers DTD Caltech101 EuroSAT Average
+Transfer
+     MoRAM 74.5 38.1 46.9 65.3 82.9 45.8 88.2 65.1 62.9 63.3
+     MoRAM w/ pruning 75.1 38.0 43.6 68.4 83.9 48.0 88.8 65.3 62.3 63.7
+Average
+     MoRAM 44.1 81.6 64.6 79.6 83.9 84.4 66.5 89.7 68.4 64.1 72.7
+     MoRAM w/ pruning 42.3 81.0 62.9 75.4 82.7 83.2 66.3 89.6 67.6 63.0 71.4
+Last
+     MoRAM 37.7 81.5 70.7 92.4 95.0 86.0 97.6 92.6 81.0 74.7 80.9
+     MoRAM w/ pruning 34.6 80.8 67.9 89.4 94.2 85.3 97.3 92.6 79.8 73.7 79.5
+Analysis of post-pruning.
+While this approach entails linear storage growth, our atomic structure offers a distinct advantage over conventional LoRA.
+Unlike standard matrices that require complex post-hoc decomposition (e.g., SVD) to compress, our rank-1 experts are independent and can be individually assessed.
+Consequently, the model inherently supports post-pruning, retaining only atoms that exceed a cumulative activation threshold, allowing for storage reduction with minimal performance degradation if required.
+We probe the effects of pruning in Table 19, where we collect the cumulative activation mass of all rank-1 memories during training and retain only the subset required to capture the top 99% of the total activation mass, which on average prunes approximately 30% of the parameter storage.
+We find that the degradation to performance is minimal, confirming that the model naturally learns a sparse representation where information is concentrated in a concise set of high-utility rank-1 memories.
+Figure 4: Averaged activation ratio of each rank-1 expert across three scenarios in Fig. 3
+A.11 Extended Visualizations of Aggregated Rank Activations
+To complement the qualitative examples in Fig.
+3, we include a statistical aggregation of rank-1 expert utilization over the entire test set for each task.
+This heatmap (Fig.
+4) provides a global view of the routing behavior and confirms that the patterns observed in Fig.
+3 are representative of the model’s overall dynamics.
+The heatmap (Fig.
+4) visualizes the activation ratio of each rank-1 expert (x-axis) across three scenarios (y-axis) in Fig.
+3:
+  1. 1.
+    Task 1 Data (after Task 1): Consistent with the qualitative results in Fig.
+3(a), we observe statistically dominant usage of Rank 0 (airplane semantics) and Rank 11 (background/sky).
+Rank 11 shows higher overall activation frequency as it captures common background tokens, which constitute a larger portion of image patches than the object itself.
+  2. 2.
+    Task 1 Data (after Task 2): Crucially, the activation pattern remains virtually unchanged after training on Task 2.
+The heatmap shows near-zero activation for the newly introduced Task 2 experts (Ranks 16-31).
+This statistically proves that our routing mechanism is stable: "old" data does not drift to "new" experts, effectively preventing catastrophic forgetting.
+  3. 3.
+    Task 2 Data (after Task 2): We observe a distinct, dual-mode behavior:
+    1. (a)
+      Knowledge Reuse: Old Rank 11 is reactivated, confirming that the model reuses the generic "blue sky" feature for the new task.
+    2. (b)
+      High Diversity: Unlike the concentrated pattern of Task 1 (Aircraft with homogeneous airplane images), Task 2 (Caltech101 with 101 diverse categories) utilizes a broad spectrum of new experts (e.g., Ranks 20, 24, 29, 30).
+This aligns with our design goal: fine-grained experts allow the model to dedicate different subspaces to the highly diverse semantics of the new task.
+We explicitly chose to visualize activations at specific representative layers rather than averaging across the entire depth of the model.
+In the modern deep models, different layers specialize in distinct feature types (e.g., low-level textures vs.
+high-level semantics).
+Averaging expert usage across all layers would smooth out these distinct signatures and obscure the specialized routing behavior we aim to demonstrate.
 […]
-(a) Rank Activations of MoRA on data from Task 1 after learning Task 1.
+(a) Rank Activations of MoRAM on data from Task 1 after learning Task 1.
 […]
-(b) Rank Activations of MoRA on data from Task 1 after learning Task 2.
+(b) Rank Activations of MoRAM on data from Task 1 after learning Task 2.
 […]
-(c) Rank Activations of MoRA on data from Task 1 after learning Task 10.
+(c) Rank Activations of MoRAM on data from Task 1 after learning Task 10.
 Figure 5: Extended view of Fig. 3 illustrating forgetting mitigation. Regions corresponding to object semantics are highlighted with orange bounding boxes. Zoom in for details.
 […]
-(a) Rank Activations of MoRA on data from Task 1 after learning Task 1.
+(a) Rank Activations of MoRAM on data from Task 1 after learning Task 1.
 […]
-(b) Rank Activations of MoRA on data from Task 9 after learning Task 1.
+(b) Rank Activations of MoRAM on data from Task 9 after learning Task 1.
 […]
-(c) Rank Activations of MoRA on data from Task 9 after learning Task 9.
+(c) Rank Activations of MoRAM on data from Task 9 after learning Task 9.
 Figure 6: Extended view of Fig. 3 illustrating knowledge reuse. Regions corresponding to generic input tokens (e.g. blue sky) are highlighted with blue bounding boxes. Zoom in for details.
-A.8 Extended Visualizations of Rank Activations
+A.12 Extended Visualizations of Rank Activations
 In Sec.4.2 (Fig.3), we illustrated rank activations during the learning of Task 1 and Task 2. Here, we extend these visualizations to additional tasks and scenarios in Fig.5 and Fig.6.
-MoRA retains task‐specific semantics without forgetting.
+MoRAM retains task‐specific semantics without forgetting.
 Fig.
@@
 Patches corresponding to the airplane object are outlined in orange.
-In all three snapshots, Rank at index 0 remains consistently and exclusively activated for those airplane patches, demonstrating that MoRA has effectively memorized the airplane‐specific knowledge into Rank 0.
+In all three snapshots, Rank at index 0 remains consistently and exclusively activated for those airplane patches, demonstrating that MoRAM has effectively memorized the airplane‐specific knowledge into Rank 0.
 Even after 10 subsequent tasks, this pattern remains unchanged, indicating that later updates do not overwrite or interfere with the learned airplane representations.
-In other words, MoRA effectively memorizes and preserves task‐relevant semantics, thereby mitigating catastrophic forgetting.
-MoRA encodes generic semantics that are reused across tasks.
+In other words, MoRAM effectively memorizes and preserves task‐relevant semantics, thereby mitigating catastrophic forgetting.
+MoRAM encodes generic semantics that are reused across tasks.
 Fig.
 6 examines an input image from Task 9 before and after learning Task 9.
-Panel (a) shows the activation map of data from Task 1 after learning Task 1: Rank 11 (outlined in blue) already responds strongly to sky‐background patches, demonstrating that MoRA has stored a generic “blue sky” concept in this rank.
-In panel (b), when we infer on the Task 9 image before training on Task 9, Rank 11 is again activated for the sky regions, confirming that MoRA reuses this shared knowledge for unseen data.
+Panel (a) shows the activation map of data from Task 1 after learning Task 1: Rank 11 (outlined in blue) already responds strongly to sky‐background patches, demonstrating that MoRAM has stored a generic “blue sky” concept in this rank.
+In panel (b), when we infer on the Task 9 image before training on Task 9, Rank 11 is again activated for the sky regions, confirming that MoRAM reuses this shared knowledge for unseen data.
 Finally, panel (c) shows the activation map after learning Task 9: Rank 11 remains dedicated to the sky background, while newly initialized ranks specialize in the “car” object semantics.
-This persistent reuse of Rank 11 across tasks illustrates MoRA’s ability to capture and retain common features as reusable memory slots, reducing redundancy and facilitating knowledge reuse.
-Figure 6: Statistical analyses on the number of ranks required to capture 99% of cumulative sum (indicated in red dashed line) of all rank activations.
+This persistent reuse of Rank 11 across tasks illustrates MoRAM ’s ability to capture and retain common features as reusable memory slots, reducing redundancy and facilitating knowledge reuse.
+Figure 7: Statistical analyses on the number of ranks required to capture 99% of the cumulative sum (indicated in red dashed line) of all rank activations.
 Activations were gathered from the model after training on all tasks, and results are shown for a representative selection of layers and positions within the pre-trained model.
-Figure 7: Statistical analyses on the number of ranks required to capture 99% of cumulative sum (indicated in red dashed line) of all rank activations.
+Figure 8: Statistical analyses on the number of ranks required to capture 99% of the cumulative sum (indicated in red dashed line) of all rank activations.
 Activations were gathered from the model after training on all tasks, and results are shown for a representative selection of layers and positions within the pre-trained model.
@@
 Figure 9: Required ranks to capture 99 % of cumulative activations, shown across different pre-trained model layers and projection locations.
-A.9 Statistical Analyses of Contributing Rank Activations
+A.13 Statistical Analyses of Contributing Rank Activations
 Fig.
@@
 Across all layers and positions, we find that fewer than 10% of the total ranks suffice to capture 99% of the activations.
-This highlights the extreme sparsity of MoRA’s self-activated mixture: most ranks remain dormant for any given input, while a compact set of highly relevant ranks drives the adaptation.
+This highlights the extreme sparsity of MoRAM ’s self-activated mixture: most ranks remain dormant for any given input, while a compact set of highly relevant ranks drives the adaptation.
 2.
@@
 For example, in Fig.
-6, the MLP’s output projection (c-proj) in Layer 1 of the vision encoder requires 16 ranks, whereas the same module in Layer 1 of the text encoder needs only 3 ranks.
+7, the MLP’s output projection (c-proj) in Layer 1 of the vision encoder requires 16 ranks, whereas the same module in Layer 1 of the text encoder needs only 3 ranks.
 To provide a broader view, Fig.
 9 shows the required rank counts for every module in the pre-trained model.
-We observe that most attention modules requires around 6–12 ranks, while the second MLP projection generally demands more ranks in early layers, peaking in the first few blocks, and then steadily declines in deeper layers.
-Coupling the rank activation budget with rank pruning, MoRA adapts the number of ranks needed to activate at each layer and module.
+We observe that most attention modules require around 6–12 ranks, while the second MLP projection generally demands more ranks in early layers, peaking in the first few blocks, and then steadily declines in deeper layers.
+Coupling the rank activation budget with rank pruning, MoRAM adapts the number of ranks needed to activate at each layer and module.
 This adaptive sparsity maximizes the efficient use of newly acquired knowledge during continual learning.

### v4

@@
 Relying on the key-value form of the rank-1 adapter, we propose a self-activation mechanism for routing the memory experts.
-We derive the mixing weights directly from the intrinsic activation of each atomic memory, leveraging the key vectors $\mathbf{A}_{i,:}$ for memory retrieval.
-Unlike standard MoE approaches that rely on additional routing networks, our self-activation routing performs routing as a content-addressable retrieval, reducing the possibility of forgetting caused by an additional router.
+We derive the mixing weights directly from the intrinsic activation of each atomic memory, leveraging the $\mathbf{A}_{i,:}$ of each rank-1 adapter component, instead of introducing a separate router for the pairs of $\mathbf{A}_{i,:}$ and $\mathbf{B}_{:,i}$ in a standard MoE-LoRA modeling.
+It is naturally the key vectors of the memory modeling for memory retrieval.
+Unlike standard MoE approaches that rely on an additional routing module, our self-activation routing performs routing as a content-addressable retrieval, reducing the possibility of forgetting caused by an additional router.
 Self-activated relevance scoring.
-By eliminating the external router, MoRAM avoids the optimization instability and forgetting associated with auxiliary networks.
+By eliminating the external router, MoRAM avoids the additional optimization instability and forgetting associated with auxiliary network parameters (i.e., the router).
 Instead, we derive mixing weights directly from the intrinsic alignment between the input and the memory keys.
@@
 This seamlessly integrates new knowledge while preserving the reusability of prior tasks.
-To enhance efficiency and enable sub-linear expansion rate, MoRAM supports low-utility atoms to be pruned after training with minimal performance degradation (more in Appendix A.10).
+To simply enhance efficiency and enable sub-linear expansion rate, MoRAM supports low-utility atoms to be pruned after training with minimal performance degradation (more in Appendix A.10), which can be further enhanced with more advanced dynamic expansion strategies.
 Training objectives.
@@
 Analysis of post-pruning.
-While this approach entails linear storage growth, our atomic structure offers a distinct advantage over conventional LoRA.
-Unlike standard matrices that require complex post-hoc decomposition (e.g., SVD) to compress, our rank-1 experts are independent and can be individually assessed.
-Consequently, the model inherently supports post-pruning, retaining only atoms that exceed a cumulative activation threshold, allowing for storage reduction with minimal performance degradation if required.
+In the default implementation setup, MoRAM increments memory units with a constant number for each task, which entails a linear growth rate for the memory size (i.e., number of experts or model size).
+To enhance the efficiency and enable a sub-linear expansion rate, a more advanced dynamic expansion strategy [78] can also be applied jointly with MoRAM.
+In simple terms, after incrementing $r$ rank-1 atomic experts and training them, low-utility components can be identified and pruned, achieving a sub-linear expansion rate.
 We probe the effects of pruning in Table 19, where we collect the cumulative activation mass of all rank-1 memories during training and retain only the subset required to capture the top 99% of the total activation mass, which on average prunes approximately 30% of the parameter storage.
 We find that the degradation to performance is minimal, confirming that the model naturally learns a sparse representation where information is concentrated in a concise set of high-utility rank-1 memories.
+The atomic design of MoRAM enables fine-grained modeling and flexible capacity expansion compared to conventional MoE–LoRA.
+In standard MoE–LoRA, experts are coarse-grained (i.e., LoRA adapters with higher rank), so adding or removing capacity is restricted to operate at a relatively large unit, or it requires post-hoc decomposition (e.g., SVD) or compression.
 Figure 4: Averaged activation ratio of each rank-1 expert across three scenarios in Fig. 3